### v1

# Quantum Blackwell’s Ordering and Differential Privacy

Abstract
We develop a framework for quantum differential privacy (QDP) based on quantum hypothesis testing and Blackwell’s ordering. This approach characterizes $(\varepsilon,\delta)$-QDP via hypothesis testing divergences and identifies the most informative quantum state pairs under privacy constraints. We apply this to analyze the stability of quantum learning algorithms, generalizing classical results to the case $\delta>0$. Additionally, we study privatized quantum parameter estimation, deriving tight bounds on the quantum Fisher information under QDP. Finally, we establish near-optimal contraction bounds for differentially private quantum channels with respect to the hockey-stick divergence.
I Introduction
A fundamental challenge in modern machine learning is the trade-off between privacy and information extraction. In this work, we explicitly treat both sides: privacy (ensuring that algorithmic outputs do not reveal significant information about the input data of the respondents) and the investigator’s goal to extract as much useful information as possible from data for accurate learning and estimation. With the rapid advancement of machine learning, a key concern is about ensuring the privacy of learning algorithms, meaning that their outputs should not reveal significant information about the input data. Differential privacy (DP) provides a rigorous mathematical framework to balance these opposing requirements. Accordingly, we structure our contributions in three steps: first step (privacy), second step (information extraction under privacy constraints), and third step, the quantum channel setup, where the situation is more complicated, and we mark the transition to each step explicitly in the text.
First step: privacy. This step develops the privacy side of the trade-off from the respondent’s perspective by studying the stability [1, 2] of learning algorithms. From the respondent’s viewpoint, privacy means that the inclusion or exclusion of their individual data should not materially affect the mechanism’s output, so that they can contribute data without fear of singled-out inference. An algorithm is considered stable if its output does not change drastically when a single respondent’s data is changed; this point-wise insensitivity is precisely the respondent-centric guarantee we seek. Differential privacy $($$\varepsilon,\delta$)-DP) [3, 4] formalizes this requirement as a strong, mathematically precise form of stability: by ensuring that the output states (distributions) corresponding to any two neighboring datasets are close, an $(\epsilon,\delta)$-DP algorithm guarantees that no respondent can be reliably distinguished from the output alone [5].
In the classical case, the stability of learning algorithms was studied in [6] only for the pure $\varepsilon$-DP ($\delta=0$) case. The main aim of this manuscript is to generalize the result of [6] to the quantum learning scenario for $(\varepsilon,\delta)$-DP. To achieve this, we build a framework for studying quantum differentially private mechanisms $($$\varepsilon,\delta$)-DP). This framework is inspired by [7] and is related to the problem of hypothesis testing (see for example, [8]). The main aim of differential privacy is to ensure that the outputs of a private mechanism corresponding to two neighboring inputs are hard to distinguish. To understand this, consider the following scenario. Suppose inside a database of datasets, we have two datasets $S$ and $T$ such that there is exactly one individual, say $A$ is present in set $S$ and not in $T$, i.e., $S\setminus(S\cap T)=\{A\}$ and suppose there is an oracle (randomized) which is attached to the database. From the perspective of an attacker, the only way it can break the privacy of this oracle is by performing a hypothesis test on the output of the oracle to detect whether the individual $A$ is a member of $S$ or not. From this discussion, it implies that the more informative the oracle is about the datasets, the less private it is.
A notion for comparing the “informativeness” between two mechanisms (oracles) was first studied by Blackwell [9]. In [10, Theorem 12.2.2], he showed that if a mechanism (oracle) is more informative as compared to the other mechanism, then there exists a Markov kernel (channel) which maps the more informative one to the less informative one. A quantum version of [10, Theorem 12.3.1] was proved by Buscemi in [11], wherein he defined a notion of ordering between quantum channels based on a certain definition of informativeness. Using this ordering, Buscemi showed that there exists a statistical morphism from the more informative quantum channel to the lesser one.
[…]
Using Lemma 1 (a quantum version of [10, Theorem 12.4.2]), we show that there exists a worst case $(\varepsilon,\delta)$-DP pair of quantum states which can be mapped to any other $(\varepsilon,\delta)$-DP pair of states by applying a CP-TP map. Thus, this version of the quantum Blackwell theorem (Lemma 1), along with this worst case $(\varepsilon,\delta)$-DP pair, provides us with a powerful machinery to analyze the stability of quantum differentially private $($$\varepsilon,\delta$)-DP) learning algorithms.
Second step: information extraction under privacy constraints. We now analyze, under the $(\varepsilon,\delta)$-DP constraints, how much useful information an investigator can still extract. That is, we investigate the statistical inference of quantum privatized parameter using our quantum version of Blackwell’s theorem (Lemma 1). This is a common task where we want to guess a hidden number or parameter from data that has been made private. How well the investigator can guess this number is limited by how much information the privacy method keeps. This limit is measured by a value called the Fisher information. In the quantum case, we use a related measure called the Symmetric Logarithmic Derivative (SLD) Fisher information to judge performance [12]. By using the idea of our “weakest” (most informative) private method, we can figure out the exact maximum SLD Fisher information the investigator can get while still respecting the $(\varepsilon,\delta)$-DP rule. This work improves on past studies that only looked at the simpler case where $\delta=0$ [13], and sets a hard limit on how accurate quantum estimations can be under the privacy condition.
Third step: We have two kinds of extensions. In the privacy of the quantum learning framework, we can relax the $(\varepsilon,\delta)$-DP condition by trusting the party, the data processor, who performs the learning algorithm. That is, motivated by the quantum learning framework of [14, 15], we propose a framework for a quantum learning algorithm which is “$1$-neighbor" $(\varepsilon,\delta)$-DP. For this proposed learning framework, we study its stability by obtaining an upper-bound on the Holevo information (between the training data and the quantum output of the algorithm). Further, we show that for $\delta=0$, our upper-bound recovers [6, Proposition 2] in the classical case.
Another interesting problem is to analyze quantum channels which are $(\varepsilon,\delta)$-DP. Its classical version has been extensively studied in the literature via the contraction of output divergence-measured with respect to the output distributions induced by $(\varepsilon,\delta)$-DP channels-in comparison to the input divergence, which is defined over the input distributions of these channels; see, for example, [16], [17], and the references therein. This contraction is studied using an integral representation which involves the hockey stick divergence [18]. We leverage our framework to first establish an almost tight upper-bound on the contraction coefficient of any $(\varepsilon,\delta)$-DP CPTP map with respect to the hockey stick divergence.
In the process of studying the contraction for divergences, we observe that in the classical case, there exists a pure $(\varepsilon+\log(1/1-\delta))$-DP pair of distributions which can be obtained by truncating the original $(\varepsilon,\delta)$-DP pair. Further, we show that this truncated pair is not far ($O(\delta)$, in $L_{1}$ distance) from the original pair. We use this truncated pair to obtain a meaningful upper-bound on the relative entropy (when it is well-defined) of the original $(\varepsilon,\delta)$-DP pair of distributions using the continuity of the relative entropy. This upper-bound recovers the known result for $\delta=0,$ [19, 7] as a special case.
Moreover, these contraction-based analysis serve to connect the channel-level action with both privacy guarantees and investigator-agnostic limits on information extraction, by bounding how much input distinctions can be suppressed at the output, we obtain direct, worst-case control of output distinguishability (useful for hypothesis-testing-based privacy statements) and of output information quantities (useful for bounding mutual information, Fisher-information-based estimation limits).
TABLE I: Relationship between results obtained in this work and related studies
Results Classical Domain Quantum Domain
$\varepsilon$-DP $(\varepsilon,\delta)$-DP $\varepsilon$-DP $(\varepsilon,\delta)$-DP
Blackwell’s Informativeness
Theorem
[10, Theorem 12.4.1 and 12.4.2]
(Proposition 1 in this manuscript.)
Lemma 1 in this manuscript.
Existence of Weakest
DP Mechanism
[19, Theorem 5]
[19, Theorem 18]
[13, Theorem 2]
Lemma 4
of this manuscript.
Data Processing based
upper-bounds using Weakest
DP Mechanism
[7, Lemma D.8]
Corollary 3
of this manuscript.
Corollary 2
of this manuscript.
Privatized Parameter
Estimation
[19, Theorem 18]
[19, Theorem 18]
[13, Section IIA]
Theorem 3
of this manuscript.
Stability Upper-bounds
of Private Learning
Algorithms
[6, Proposition 2]
Proposition 2
of this manuscript.
Equations (i)),
(ii)) and (iii)) of
this manuscript.
Theorem 5
of this manuscript.
Theorem 5
of this manuscript.
Privatized Contraction
Coefficient of Trace Distance
[19, Corollary 11]
[20, Theorem 22]
[20, Theorem 55]
Contraction Coefficient
of Hockey-stick Divergence
[17, Theorem 1]
Corollary 9
of this manuscript.
[20, Theorem 1]
Lemma 7
of this manuscript.
Contraction Coefficient based
Upper-bounds on Relative entropy
of LDP channels
Theorem 6
of this manuscript.
[20, Proposition 3]
I-A Organization of this Manuscript and our Contributions
The rest of this manuscript is organized as follows:
  • In Section II, we list the essential notations, definitions, and facts which will be used throughout this manuscript.
  • In Section III, we develop a quantum generalization of Blackwell’s theorem, which establishes an ordering of informativeness for pairs of quantum states. In Lemma 1, we show that one pair is less informative than the other if and only if there exists a completely positive map connecting them. This forms the foundation for comparing privacy mechanisms in the quantum setting.
  • In Section V, we prove Lemma 3, which shows that a pair of quantum states is $(\varepsilon,\delta)$-differentially private if and only if $D^{\alpha}_{H}{(\rho\|\sigma)}\leq\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1],$ where $D^{\alpha}_{H}{(\rho\|\sigma)}$ is the hypothesis testing divergence between $\rho$ and $\sigma.$ Using this in Lemma 4, we identify an explicit “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states that dominates all other.
  • In Section VI, we apply our framework to quantum parameter estimation under privacy constraints. In particular, in Theorem 3, we derive the exact maximum SLD Fisher information achievable by any $(\varepsilon,\delta)$-DP mechanism.
  • In Section VII, we study the stability of $(\varepsilon,\delta)$-DP learning algorithms. We first discuss a framework for differentially private learning algorithms. Then, using this framework, in Theorem 5, we study their stability by deriving upper-bound on the Holevo information.
  • In Section VIII, we analyze the contraction of divergences under differentially private channels. In Lemma 7, we obtain nearly tight upper and lower bounds on the contraction coefficient of $(\varepsilon,\delta)$-LDP channels with respect to the quantum hockey-stick divergence. We then use this to derive a novel upper-bound on the relative entropy for classical $(\varepsilon,\delta)$-LDP channels in Theorem 6 via its integral representation [18].
Table I above summarizes the relation between the results obtained in our manuscript and the existing results.
II Notations, Definitions and Facts
We use $\mathcal{H}$ to denote a finite-dimensional Hilbert space and we denote its dimension with $\left|\mathcal{H}\right|$, $\mathcal{D}(\mathcal{H})$ to represent the set of all state density matrices acting on $\mathcal{H}$. For any quantum state $\rho\in\mathcal{D}(\mathcal{H})$, we define $\mbox{supp}(\rho):=\mbox{Span}\{|i\rangle:\lambda_{i}>0\},$ where $\left\{\lambda_{i}\right\}$ represent non-zero eigenvalues of $\rho$. For any finite and non-empty set $\mathcal{X}$, we denote $\mathcal{P}(\mathcal{X})$ as the set of all probability distributions over $\mathcal{X}$. Similarly, for any distribution $P\in\mathcal{P}(\mathcal{X})$, $\mbox{supp}(P):=\{x\in\mathcal{X}:P(x)>0\}.$ For any sequence $x^{n}:(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, a permutation $\pi:\mathcal{X}^{n}\to\mathcal{X}^{n},$ is a map such that $\pi\left(x^{n}\right):=(x_{\pi^{-1}(1)},\cdots,x_{\pi^{-1}(n)}).$ Let $S_{n}$ be the permutation group containing all permutations of length $n$ and it is also known as $n$-th symmetric group.
Definition 1.
(Type and Set of all types) A vector of integers ${\bf f}:=\left(f_{1},\cdots,f_{d}\right)$ (where $d>0$) is called a type of size $n$ and length $d$ if $\forall i\in[d],f_{i}\geq 0$ and $\sum_{i}f_{i}=n.$ $T^{n}_{d}$ is denoted as the set of all distinct types of size $n$ and length $d$.
Definition 2 (Frequency type classes).
For a non-empty finite set $\mathcal{X}$ and an integer $n>0$, given a type function ${\bf f}:=(f_{1},\cdots,f_{\left|\mathcal{X}\right|}])$ ( such that $\sum_{i\in\mathcal{X}}f_{i}=n$), we define a set $T_{{\bf f}}\subset\mathcal{X}^{n}$ to be the frequency type class or type set corresponding to $\bf f$} as follows,
$T_{\bf f}:=\left\{(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}:\frac{\left\lvert\left\{k:x_{k}=i\right\}\right\rvert}{n}=\bar{f}(i),\forall i\in\mathcal{X}\right\},$
where $\bar{f}(i):=\frac{f_{i}}{n}$ for all $i\in\mathcal{X}$.
Definition 3 ([6]).
A pair of $n$-length (where $n\in\mathbb{N}$ ) sequences $x^{n}:=(x_{1},\cdots,x_{n}),\tilde{x}^{n}:=(\tilde{x}_{1},\cdots,\tilde{x}_{n})\in\mathcal{X}^{n}$ (where $\left|\mathcal{X}\right|>0$) is said to be $k$-neighbors, represented by $x^{n}\overset{k}{\sim}\tilde{x}^{n}$ if,
$k=\frac{1}{2}\sum_{a\in\mathcal{X}}\left|f(a|x^{n})-f(a|\tilde{x}^{n})\right|,$ (1)
where, $f(a|x^{n})$ denotes the number of appearance of the alphabet $a\in\mathcal{X}$ in the sequence $x^{n}$.
Further, for any sequence $x^{n}$ and its some permutation $\tilde{x}^{n}$, we denote $x^{n}\overset{0}{\sim}\tilde{x}^{n}$.
Definition 4 (Minimun type-II error).
Given a pair of probability distributions $P_{1},P_{2}$ over a finite set $\mathcal{X}$, the minimum type-II error of a fixed order $\alpha\in[0,1]$ is defined as follows,
$\beta(\alpha,P_{1},P_{2}):=\min_{\begin{subarray}{c}\phi:\mathcal{X}\to[0,1]\\ \sum_{x\in\mathcal{X}}P_{1}(x)\phi(x)\leq\alpha\end{subarray}}1-\sum_{x\in\mathcal{X}}P_{2}(x)\phi(x).$ (2)
Definition 5 (Classical hypothesis testing divergence).
For any pair of probability distributions $P,Q$ over a finite set $\mathcal{X}$, and for any $\alpha\in[0,1]$, the classical hypothesis testing divergence of order $\alpha$ is defined as
$D^{\alpha}_{H}(P\|Q):=\max_{\begin{subarray}{c}\phi:\mathcal{X}\to[0,1]\\ \sum_{x\in\mathcal{X}}P(x)\phi(x)\leq\alpha\end{subarray}}-\ln\left(1-\sum_{x\in\mathcal{X}}Q(x)\phi(x)\right).$ (3)
Further, note that for any $\alpha\in[0,1]$, $D^{\alpha}_{H}(P\|Q)=-\ln\beta(\alpha,P,Q)$ where $\beta(\alpha,P,Q)$ is as defined in Definition 4.
Definition 6 (Classical hockey-stick divergence[21]).
Let $P$ and $Q$ be probability distributions on a finite (or measurable) space $\mathcal{X}$, and let $\gamma\geq 1$. Then, the hockey-stick divergence between $P$ and $Q$ is defined as
$E_{\gamma}(P\|Q):=\sup_{A\subseteq\mathcal{X}}\Big(P(A)-\gamma\,Q(A)\Big)=\sum_{x\in\mathcal{X}}\big[P(x)-\gamma Q(x)\big]_{+},$
where $t]_{+}:=\max\{t,0\}$ denotes the positive part of $t$.
Definition 7 (Quantum Instrument [22]).
A quantum instrument consists of a collection $\{\mathcal{E}_{j}\}$ of completely positive, trace non-increasing maps such that the sum map $\sum_{j}\mathcal{E}_{j}$ is trace preserving. Let $\{|j\rangle\}_{j}$ be an orthonormal basis for a Hilbert space $\mathcal{H}_{J}$. The action of a quantum instrument on a density operator $\rho\in\mathcal{D}(\mathcal{H})$ is the following quantum channel, which features a quantum and classical output:
$\rho\mapsto\sum_{j}\mathcal{E}_{j}(\rho)\otimes|j\rangle\langle j|_{J}.$ (4)
Definition 8 (Smooth Max-Relative Entropy).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\varepsilon\in[0,1)$, the smooth max-relative entropy of $\rho$ with respect to $\sigma$ is defined as
$D_{\max}^{\varepsilon}(\rho\|\sigma):=\min_{\tilde{\rho}\in\mathcal{B}^{\varepsilon}(\rho)}D_{\max}(\tilde{\rho}\|\sigma),$ (5)
where $D_{\max}(\tilde{\rho}\|\sigma):=\inf\{\lambda\in\mathbb{R}:\tilde{\rho}\leq 2^{\lambda}\sigma\}$ and $\mathcal{B}^{\varepsilon}(\rho):=\left\{\tilde{\rho}\in\mathcal{D}(\mathcal{H}_{A}):\frac{1}{2}\|\tilde{\rho}-\rho\|_{1}\leq\varepsilon\right\}$.
Definition 9.
A function $\mathbb{D}:\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})\to\mathbb{R}^{+}$ (where $\mathcal{X}$ is an arbitrary set of finite cardinality) is called a divergence if for any classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$, (where $\mathcal{Y}$ is another arbitrary set of finite cardinality), the following holds,
$\mathbb{D}(P_{1}\|P_{2})\geq\mathbb{D}(P^{\mathcal{K}}_{1}\|P^{\mathcal{K}}_{2}),\forall P_{1},P_{2}\in\mathcal{P}(\mathcal{X}),$ (6)
where for each $i=1,2$, $P^{\mathcal{K}}_{i}(y):=\mathbb{E}_{X\sim P_{i}}\left[\mathcal{K}(y\mid X)\right]$.
Definition 10.
[…]
Definition 11.
Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, then, we define the quantum Divergence between $\rho$ and $\sigma$ as follows
$D(\rho\|\sigma):=\begin{cases}\mathrm{Tr}[\rho(\log\rho-\log\sigma)],&\text{ if }\rho\ll\sigma,\\ +\infty,&\text{ else}.\end{cases}$
Definition 12 (Petz Quantum Rényi Divergence [23]).
Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1]\cup(1,+\infty)$. Then, Petz Quantum Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
$D_{\alpha}(\rho\|\sigma):=\begin{cases}\frac{1}{\alpha-1}\log\mathrm{Tr}\left[\rho^{\alpha}\sigma^{1-\alpha}\right],&{\mbox{if }}(\alpha<1\cap\rho\not\perp\sigma)\cup(\rho\ll\sigma),\\ +\infty,&\mbox{else}.\end{cases}$
Definition 13 (Quantum hockey stick divergence [24]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we define the quantum hockey stick divergence $E_{\gamma}(\rho\|\sigma)$ of order $\gamma\geq 0$ as follows,
$E_{\gamma}(\rho\|\sigma):=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}\end{subarray}}\mathrm{Tr}[\Lambda(\rho-\gamma\sigma)].$ (8)
Fact 1.
For any type ${\bf f}=(f_{1},\cdots,f_{d})$, the set $T_{\bf f}$ (see Definition 2) is a permutation invariant set, i.e., for every sequence $x^{n}:(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, under any permutation $\pi\in S_{n},$ $\pi\left(x^{n}\right)\in T_{\bf f}$ where $\pi_{n}\left(x^{n}\right):=(x_{\pi^{-1}(1)},\cdots,x_{\pi^{-1}(n)}).$
Fact 2.
[25] $T^{n}_{d}$ (See Definition 1) satisfies the following upper-bound :
$\left|\Lambda^{n}_{d}\right|\leq\left|T^{n}_{d}\right|\leq(n+1)^{d-1}.$ (9)
Fact 3.
For any two pair $(P,Q)$ and $(P^{\prime},Q^{\prime})$ probability distributions, the following holds for any $\gamma\geq 1$,
$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\leq E_{\gamma}(P\|Q)+{\left\lVert P^{\prime}-P\right\rVert}_{1}+\gamma{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$ (10)
Proof.
From Definition 13, we have
$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $=\max_{\begin{subarray}{c}0\leq\Lambda(x)\leq 1,\\ x\in\mathcal{X}\end{subarray}}\sum_{x\in\mathcal{X}}\Lambda(x)(P^{\prime}(x)-\gamma Q^{\prime}(x))$
$\leq\max_{\begin{subarray}{c}0\leq\Lambda(x)\leq 1,\\ x\in\mathcal{X}\end{subarray}}\sum_{x\in\mathcal{X}}\Lambda(x)(P(x)-\gamma Q(x))+\sum_{x\in\mathcal{X}}|P^{\prime}(x)-P(x)|+\gamma\sum_{x\in\mathcal{X}}|Q^{\prime}(x)-Q(x)|$
$\leq E_{\gamma}(P\|Q)+{\left\lVert P^{\prime}-P\right\rVert}_{1}+\gamma{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$
This proves Fact 3.
Fact 4 (Hölder’s inequality).
For any two vectors $x=(x_{1},\dots,x_{n})$ and $y=(y_{1},\dots,y_{n})$ in $\mathbb{R}^{n}$ and for any real numbers $p,q\in[1,\infty)$ such that $\frac{1}{p}+\frac{1}{q}=1$, we have,
$\sum_{i=1}^{n}\left|x_{i}y_{i}\right|\leq\left(\sum_{i=1}^{n}\left|x_{i}\right|^{p}\right)^{1/p}\left(\sum_{i=1}^{n}\left|y_{i}\right|^{q}\right)^{1/q}.$ (11)
Further, if $p=1$ and $q=\infty$, then the right-hand side of (11) can be interpreted as follows,
$\left(\sum_{i=1}^{n}\left|x_{i}\right|^{p}\right)^{1/p}\left(\sum_{i=1}^{n}\left|y_{i}\right|^{q}\right)^{1/q}=\left(\sum_{i=1}^{n}\left|x_{i}\right|\right)\max_{i\in[n]}\left|y_{i}\right|.$ (12)
Fact 5 ([18, Eq. $428$]).
Let $P$ and $Q$ be probability distributions on a finite (or measurable) space $\mathcal{X}$. Then, the classical relative entropy (Kullback-Leibler divergence) admits the following integral representation in terms of the classical hockey-stick divergence:
$D(P\|Q)=\int_{0}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(P\|Q)+\frac{1}{\gamma^{2}}E_{\gamma}(Q\|P)\right)d\gamma,$ (13)
where $E_{\gamma}(P\|Q)$ is the classical hockey-stick divergence as defined in Definition 6.
Fact 6 ([26]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the quantum relative entropy $D(\rho\|\sigma)$ has the following integral representation in terms of quantum hockey stick divergence,
$D(\rho\|\sigma)=\int_{0}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(\rho\|\sigma)+\frac{1}{\gamma^{2}}E_{\gamma}(\sigma\|\rho)\right)d\gamma.$ (14)
Fact 7.
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, we have the following equality,
$\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]=\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (15)
Proof.
[…]
$\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]$ $=\delta\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$
$\overset{a}{<}\beta^{\prime},$
where $a$ follows from the fact that $\delta<1$ and $\beta^{\prime}=\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$. Thus, for any POVM $\{\tilde{\Lambda},\mathbb{I}-\tilde{\Lambda}\}$, such that $\mathrm{Tr}[\tilde{\Lambda}\rho]<\alpha$, there exists a POVM $\{\hat{\Lambda},\mathbb{I}-\hat{\Lambda}\}$ such that $\mathrm{Tr}[\tilde{\Lambda}\rho]=\alpha$ and $\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]<\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$. Therefore, the optimum $\Lambda$ in the LHS of (15) satisfies $\mathrm{Tr}[{\Lambda}\rho]=\alpha$. This proves Fact 7.
Fact 8 ([24]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the quantum hockey stick divergence $E_{\gamma}(\rho\|\sigma)$ of order $\gamma\geq 0$ (see Definition 13) has the following equivalent form,
$E_{\gamma}(\rho\|\sigma)=\mathrm{Tr}|\Lambda(\rho-\gamma\sigma)|_{+},$ (16)
where $|\cdot|_{+}$ denotes the positive part of the operator, i.e., $|O|_{+}:=\sum_{i}\max\{0,\lambda_{i}\}|\psi_{i}\rangle\langle\psi_{i}|$, where $\lambda_{i}$ and $|\psi_{i}\rangle$ are the eigenvalues and eigenvectors of $O$, respectively.
Fact 9.
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$ the following equality holds,
$\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\geq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]=\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (17)
Proof.
Proof of Fact 9 follows directly from the proof of Fact 7.
Fact 10 (Reverse quantum data processing inequality).
Consider $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$. Then, if $\forall\gamma\geq 0$, $E_{\gamma}(\rho_{1}\|\rho_{2})\geq E_{\gamma}(\sigma_{1},\sigma_{2})$, then there exists a completely positive map $\mathcal{T}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ such that $\sigma_{i}=\mathcal{T}(\rho_{i})$, for each $i\in\{1,2\}$.
Proof.
The proof of Fact 10 follows directly from Theorem 55 and Theorem $6$ of [27].
Fact 11 (Data-processing inequality of quantum relative entropy).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, $D(\rho\|\sigma)$ the following holds,
$D(\rho\|\sigma)\geq D(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (18)
where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
Fact 12 (Data-processing inequality for quantum hypothesis testing divergence [28]).
Consider any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and any CP-TP map $\mathcal{N}:\mathcal{H}_{A}\to\mathcal{H}_{B}$. Then for any $\alpha\in[0,1]$ the following satisfies,
$\mathcal{D}^{\alpha}_{H}(\rho\|\sigma)\geq\mathcal{D}^{\alpha}_{H}(\left(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\right).$ (19)
Fact 13 (Data-processing inequality of Petz quantum Rényi divergence [23]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ and $\forall\alpha\in[0,1)\cup(1,2]$, $D_{\alpha}(\rho\|\sigma)$ satisfies the following,
$D_{\alpha}(\rho\|\sigma)\geq D_{\alpha}(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (20)
where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
Fact 14 (Data-processing inequality of quantum Hockey stick divergence [24, Lemma $4$]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ and $\forall k\geq 0$, $E_{k}(\rho\|\sigma)$ satisfies the following,
$E_{k}(\rho\|\sigma)\geq E_{k}(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (21)
where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
Fact 15 (Monotonicity of SLD Fisher Information [29]).
Let $\rho_{\theta}$ be a differentiable family of quantum states and let $\mathcal{E}$ be a completely positive trace-preserving (CP-TP) map. Then, the SLD Fisher information is monotonic under $\mathcal{E}$:
$J_{\theta}(\rho_{\theta})\geq J_{\theta}(\mathcal{E}(\rho_{\theta})).$ (22)
Fact 16 (Quantum Reversed Pinsker inequality[30, theorem 2]).
For quantum states $\rho$ and $\sigma$, the following inequality holds:
$D(\rho\,\|\,\sigma)\leq\frac{2}{\lambda_{\min}(\sigma)}E_{1}(\rho\,\|\,\sigma)^{2}.$ (23)
where $\lambda_{\min}(\sigma)$ is the smallest non-zero eigenvalue of $\sigma$.
Fact 17.
Consider $\rho,\rho^{\prime},\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, the following inequality holds:
$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$ (24)
Proof.
By the definition of the max-relative entropy, we have
$\rho^{\prime}\leq e^{D_{\max}(\rho^{\prime}\|\sigma)}\,\sigma.$
Equivalently,
$\sigma\geq e^{-D_{\max}(\rho^{\prime}\|\sigma)}\,\rho^{\prime}.$
Since the logarithm is operator monotone, this implies
$\log\sigma\succeq\log\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
Multiplying both sides by $\rho$ and taking the trace (which reverses the inequality), we obtain
$\mathrm{Tr}[\rho\log\sigma]\leq-\mathrm{Tr}[\rho\log\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
Adding $\mathrm{Tr}[\rho\log\rho]$ to both sides gives
$\mathrm{Tr}[\rho(\log\rho-\log\sigma)]\leq\mathrm{Tr}[\rho(\log\rho-\log\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
which can be written as
$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
This completes the proof of Fact 17.
Fact 18.
Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, for any $\varepsilon\geq 0,\delta\in[0,1)$, the following statements are equivalent,
(1) $\rho$ and $\sigma$ satisfies,
$\mathrm{Tr}[\Lambda\rho]\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,\mbox{~for any }0\preceq\Lambda\preceq\mathbb{I}.$ (25)
(2) $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$, where $E_{e^{\varepsilon}}(\cdot\|\cdot)$ is the quantum hockey stick divergence (see Definition 13) of order $e^{\varepsilon}$.
Proof.
We first prove that (1) implies (2). From (25), we have that for any $0\preceq\Lambda\preceq\mathbb{I}$,
$\mathrm{Tr}[\Lambda\rho]-e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]\leq\delta,$
which can be rewritten as
$\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
Taking the maximum over all $0\preceq\Lambda\preceq\mathbb{I}$ on the left-hand side, we get
$\max_{0\preceq\Lambda\preceq\mathbb{I}}\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
By the definition of the quantum hockey stick divergence (Definition 13), the left-hand side is exactly $E_{e^{\varepsilon}}(\rho\|\sigma)$. Thus, $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$.
Next, we prove that (2) implies (1). Assume $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$. By definition, this means
$\max_{0\preceq\Lambda^{\prime}\preceq\mathbb{I}}\mathrm{Tr}[\Lambda^{\prime}(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
This implies that for any specific choice of $0\preceq\Lambda\preceq\mathbb{I}$, we must have
$\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\max_{0\preceq\Lambda^{\prime}\preceq\mathbb{I}}\mathrm{Tr}[\Lambda^{\prime}(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
Rearranging the terms, we get
$\mathrm{Tr}[\Lambda\rho]\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,$
which is the condition in (25). This completes the proof.
Fact 19 ([31, Lemma $6.9$]).
If $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$, for a pair $(\rho,\sigma)\in\mathcal{D}(\mathcal{H}_{A})$, then,
$D_{\max}(\tilde{\rho}\|\sigma)\leq\varepsilon-\log\left({1-\delta}\right),$
where $\varepsilon\geq 0,\delta\in[0,1)$ and $\tilde{\rho}:=\frac{G\rho G^{\dagger}}{\mathrm{Tr}\left[G\rho G^{\dagger}\right]}$, has the property that,
$\frac{1}{2}{\left\lVert\rho-\tilde{\rho}\right\rVert}_{1}\leq\sqrt{\delta(2-\delta)},$
where $G:=\left(e^{\varepsilon}\sigma\right)^{\frac{1}{2}}\left(e^{\varepsilon}\sigma+\left|\rho-e^{\varepsilon}\sigma\right|_{+}\right)^{-\frac{1}{2}}.$ Further, if $\operatorname{supp}(\rho)=\operatorname{supp}(\sigma),$ then, $\operatorname{supp}(\tilde{\rho})=\operatorname{supp}(\rho)=\operatorname{supp}(\sigma).$
III Blackwell’s Dominance of Quantum Mechanisms
Blackwell [9] in classical statistics provided a framework to compare two statistical experiments in terms of their informativeness. An experiment is considered more informatively dominating if it allows for better decision-making based on the observed data. This idea of dominance can be viewed from the point of view of hypothesis testing. Consider $\mathcal{E}_{1}:=\left\{P_{1},P_{2}\right\}$ and $\mathcal{E}_{2}:=\left\{Q_{1},Q_{2}\right\}$ be two classical statistical experiments, where $P_{1},P_{2}$ and $Q_{1},Q_{2}$ are probability distributions over two finite sets $\mathcal{X}$ and $\mathcal{Y}$. Consider the two following Hypothesis tests:
$H(\mathcal{E}_{1})$ $:=\begin{cases}H^{\mathcal{E}_{1}}_{0}:\mbox{A random variable }X\sim P_{1},\\ H^{\mathcal{E}_{1}}_{1}:\mbox{A random variable }X\sim P_{2}.\end{cases}$
$H(\mathcal{E}_{2})$ $:=\begin{cases}H^{\mathcal{E}_{2}}_{0}:\mbox{A random variable }Y\sim Q_{1},\\ H^{\mathcal{E}_{2}}_{1}:\mbox{A random variable }Y\sim Q_{2}.\end{cases}$
We now say that $\mathcal{E}_{2}$ is dominated by $\mathcal{E}_{1}$ (written as $\mathcal{E}_{2}\preceq_{B}\mathcal{E}_{1}$) if for any test (decision rule) $\phi_{\mathcal{E}_{2}}:\mathcal{Y}\to[0,1]$ for $H(\mathcal{E}_{2})$, there exists a test $\phi_{\mathcal{E}_{1}}:\mathcal{X}\to[0,1]$ for $H(\mathcal{E}_{1})$ such that $\phi_{\mathcal{E}_{1}}$ performs at least as well as $\phi_{\mathcal{E}_{2}}$. We formalize this notion in the following proposition.
Proposition 1 (Blackwell’s Theorem of Informative Dominance[10, Theorems $12.4.1$ & $12.4.2$]).
Given two pairs of probability distributions $\left(P_{1},P_{2}\right)$ and $\left(Q_{1},Q_{2}\right)$ over two finite sets $\mathcal{X}$ and $\mathcal{Y}$, respectively, the following statements are equivalent,
  1. 1.
    $\left\{Q_{1},Q_{2}\right\}\preceq_{B}\left\{P_{1},P_{2}\right\}$.
  2. 2.
    $\beta(\alpha,Q_{1},Q_{2})\geq\beta(\alpha,P_{1},P_{2})$ for all $\alpha\in[0,1]$, where $\beta(\alpha,P_{1},P_{2})$ and $\beta(\alpha,Q_{1},Q_{2})$ is the minimum type-II error of order $\alpha$ (see Definition 4) of the hypothesis tests $H(\left\{P_{1},P_{2}\right\})$ and $H(\left\{Q_{1},Q_{2}\right\})$ respectively.
  3. 3.
    $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}\leq D^{\alpha}_{H}{(P_{1}\|P_{2})}$ for all $\alpha\in[0,1]$, where $D^{\alpha}_{H}{(P_{1}\|P_{2})}$ and $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ are the classical hypothesis testing divergence of order $\alpha$ (see Definition 5) between the pair of distributions $\left\{P_{1},P_{2}\right\}$ and $\left\{Q_{1},Q_{2}\right\}$ respectively.
  4. 4.
    There exists a stochastic map $\mathcal{T}:\mathcal{X}\to\mathcal{Y}$ such that $Q_{i}=\mathcal{T}(P_{i})$ for each $i\in\{1,2\}$.
Observe that Proposition 1 above provides a partial order $\preceq_{B}$ on experiments, i.e., we say an experiment $\mathcal{E}_{2}$ is dominated by (or less informative than) another experiment $\mathcal{E}_{1}$ if $\mathcal{E}_{2}\preceq_{B}\mathcal{E}_{1}$.
Similarly, in binary asymmetric quantum hypothesis testing, given any pair of quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we perform a hypothesis test between $\rho$, then, the null hypothesis, and $\sigma$, the alternative hypothesis. The test is biased towards the null hypothesis, i.e., if the given state is originally $\rho$, then the test should reject the null hypothesis with a small probability, while accepting the null hypothesis with a small probability if the given state is $\sigma$. A quantum hypothesis test is typically performed using a binary POVM $\{\Lambda,\mathbb{I}-\Lambda\}$, where $\Lambda$ is the rejection operator and $\mathbb{I}-\Lambda$ is the acceptance operator. The hypothesis test can incur two types of errors, namely, the type-I error (false negative) and the type-II error (false positive). The type I error is the probability of rejecting the null hypothesis when it is true, while the type II error is the probability of accepting the null hypothesis when it is false. Under the binary POVM $\{\Lambda,\mathbb{I}-\Lambda\}$, the type-I error is given by $\mathrm{Tr}[\Lambda\rho]$ and the type-II error is given by $\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]$. To characterize the trade-off between the type-I and type-II errors, in asymmetric quantum hypothesis testing, we fix the type-I error to a value $\alpha\in[0,1]$ and maximize the negative logarithm of type-II error over all possible POVMs $\{\Lambda,\mathbb{I}-\Lambda\}$, which gives us the following definition of quantum hypothesis testing divergence.
Definition 14 (Quantum hypothesis testing divergence [28]).
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, the quantum hypothesis testing divergence $D^{\alpha}_{H}{(\rho\|\sigma)}$ of order $\alpha$ is defined as follows,
$D^{\alpha}_{H}{(\rho\|\sigma)}:=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}-\ln\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (26)
For any pair of quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we denote its fixed point as the level of type-I error $\alpha\in[0,1]$ for which the type-II error is equal to the type-I error, i.e., $e^{-D^{\alpha}_{H}{(\rho\|\sigma)}}=\alpha$. This is the point where both hypotheses are equally likely. Similar to classical Blackwell’s ordering, we define quantum Blackwell’s order as follows,
Definition 15 (Quantum Blackwell’s order).
Given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$ over two finite-dimensional Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, we say that $\left\{\sigma_{1},\sigma_{2}\right\}$ is dominated by (or less informative than) $\left\{\rho_{1},\rho_{2}\right\}$ (denoted as $\left\{\sigma_{1},\sigma_{2}\right\}\preceq_{B_{Q}}\left\{\rho_{1},\rho_{2}\right\}$) if for any $\alpha\in[0,1]$, the following holds,
$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})}.$ (27)
Lemma 1 (Quantum Blackwell’s Theorem).
Given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$ over two finite-dimensional Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, the following statements are equivalent,
  1. 1.
    $\left\{\sigma_{1},\sigma_{2}\right\}\preceq_{B_{Q}}\left\{\rho_{1},\rho_{2}\right\}$.
  2. 2.
    There exists a completely positive map $\mathcal{T}:\mathcal{D}(\mathcal{H}_{A})\to\mathcal{D}(\mathcal{H}_{B})$, such that $\sigma_{1}=\mathcal{T}(\rho_{1})$ and $\sigma_{2}=\mathcal{T}(\rho_{2})$.
Proof.
Before proceeding with the proof of Lemma 1, we first prove the following lemma, which will be useful in proving Lemma 1.
Lemma 2.
Consider $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$. If
$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})},\forall\alpha\in[0,1],$ (28)
then,
$E_{\gamma}(\rho_{1}\|\rho_{2})\geq E_{\gamma}(\sigma_{1}\|\sigma_{2}),\forall\gamma\geq 0,$
where $E_{\gamma}(\cdot\|\cdot)$ is defined in Definition 13 for any $\gamma\geq 0$.
Proof.
See Appendix -A for the proof.
It is easy to follow that $(2)\implies(1)$ as a consequence of Definition 15 and Fact 12. We now proceed to prove $(1)\implies(2)$. From Definition 15 $(1)$ implies (28). Further, from Lemma 2 and Fact 10, it follows that, if for the pairs $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$, (28) holds, then there exists a completely positive map $\mathcal{T}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ such that $\sigma_{i}=\mathcal{T}(\rho_{i})$, for each $i\in\{1,2\}$. This completes the proof of Lemma 1.
The concept of Blackwell dominance for pairs of quantum states, as presented in Definition 15, is related to the framework for the statistical comparison of quantum channels developed by Buscemi [11]. In that work, an informational ordering is defined between two quantum channels, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, based on the guessing probability of any ensemble of states passed through them. Specifically, channel $\mathcal{N}_{1}$ is considered more informative than $\mathcal{N}_{2}$ if, for every ensemble of input states, the guessing probability after passing through $\mathcal{N}_{1}$ is at least as high as that after passing through $\mathcal{N}_{2}$. This ordering is shown to be equivalent to the existence of a completely positive map that transforms the output of $\mathcal{N}_{1}$ into that of $\mathcal{N}_{2}$, mirroring the condition in Lemma 1. However, it remains an open question whether this ordering can be fully characterized by the quantum hypothesis testing divergence, as in Definition 15. Addressing this question would deepen our understanding of the relationship between quantum channel comparison and hypothesis testing.
IV Characteristic Region of $(\varepsilon,\delta)$-Differentially Private Quantum Mechanisms
In this section, we introduce the concept of the characteristic region of a quantum mechanism that satisfies $(\varepsilon,\delta)$-quantum differential privacy (QDP) as defined in [4, Definition $2$]. The characteristic region provides a geometric representation of the trade-offs between the privacy parameters $\varepsilon$ and $\delta$ for a given private quantum mechanism. To illustrate this concept, we consider the following definition of a pair of quantum states to be $(\varepsilon,\delta)$-DP.
Definition 16.
A pair $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ is defined to be $(\varepsilon,\delta)$-DP (differentially private) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, if every POVM measurement $0\preceq\Lambda\preceq\mathbb{I}$, the following holds,
$\begin{split}\mathrm{Tr}[\Lambda\rho]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,\\ \mathrm{Tr}[\Lambda\sigma]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\rho]+\delta.\end{split}$ (29)
Further, we denote $\mathfrak{D}_{(\varepsilon,\delta)}$ to be the collection of all pairs of quantum states that satisfy (29). Moreover, for $\delta=0$, we denote $(\rho,\sigma)$ to be pure $\varepsilon$-DP or just $\varepsilon$-DP i.e. $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$.
Consider a hypothesis test between two quantum states $\rho$ and $\sigma$ over a finite dimensional Hilbert space $H$ as follows,
$H$ $:=\begin{cases}H_{0}:\mbox{Given quantum state is}~\rho,\\ H_{1}:\mbox{Given quantum state is}~\sigma.\end{cases}$
Under the choice of a rejection POVM $0\preceq\Lambda\preceq\mathbb{I}$ for $H_{0}$, the type-I error is given by $\alpha_{\Lambda}:=\mathrm{Tr}\left[\Lambda\rho\right]$ and the type-II error is given by $\beta_{\Lambda}:=1-\mathrm{Tr}\left[\Lambda\sigma\right]$. Now, for any pair $(\rho,\sigma)$ of quantum states, the characteristic region is defined as follows,
Definition 17.
For any pair of quantum states $(\rho,\sigma)$, the characteristic region $\mathcal{R}(\rho,\sigma)$ is defined as follows,
$\mathcal{R}(\rho,\sigma):=\left\{(\alpha_{\Lambda},\beta_{\Lambda}):0\preceq\Lambda\preceq\mathbb{I}\right\}.$ (31)
Note that from Definition 15 it follows that for any two pairs of quantum states $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, we have $\{\sigma_{1},\sigma_{2}\}\preceq_{B_{Q}}\{\rho_{1},\rho_{2}\}$ if and only if $\mathcal{R}(\sigma_{1},\sigma_{2})\subseteq\mathcal{R}(\rho_{1},\rho_{2})$. Further, if $(\rho,\sigma)$ is $(\varepsilon,\delta)$-QDP, then for any rejection POVM $\Lambda$, from (29) the following constraints hold,
$\beta_{\Lambda}$ $\geq e^{-\varepsilon}(1-\delta-\alpha_{S}),$ (32)
$\beta_{\Lambda}$ $\geq 1-\delta-e^{\varepsilon}\alpha_{S},$ (33)
$\beta_{\Lambda}$ $\leq 1-e^{-\varepsilon}(\alpha_{S}-\delta),$ (34)
$\beta_{\Lambda}$ $\leq e^{\varepsilon}(1-\alpha_{S})+\delta.$ (35)
$\alpha\longrightarrow$
(Type-I error)
$\beta\longrightarrow$
(Type-II error)
[…]
$\left(\frac{1-\delta}{1+e^{\varepsilon}},\frac{1-\delta}{1+e^{\varepsilon}}\right)$
(Worst fixed-point)
$\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}\right)$
(Best fixed-point)
[…]
Figure 1: Graphical representation of $\mathcal{R}(\varepsilon,\delta)$ : characteristic region of $(\varepsilon,\delta)$-QDP (in shaded region), where we define the boundaries (a), (b), (c) and (d) to be $\beta=1-\delta-e^{\varepsilon}\alpha$, $\beta=e^{-\varepsilon}(1-\delta-\alpha)$, $\beta=1-e^{-\varepsilon}(\alpha-\delta)$ and $\beta=e^{\varepsilon}(1-\alpha)+\delta$ respectively and $\mathcal{R}(\varepsilon,\delta)$ has two fixed points $(\frac{1-\delta}{1+e^{\varepsilon}},\frac{1-\delta}{1+e^{\varepsilon}})$ and $(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}})$ as its extremal points, where the former and the latter points are known to be the worst and the best fixed points respectively from the perspective of privacy.
1n Figure 1, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ of $(\varepsilon,\delta)$-QDP.
Therefore, under the constraints mentioned in eqs.˜32, 33, 34 and 35, we define the characteristic (or operating) region of quantum $(\varepsilon,\delta)$-QDP as follows,
Definition 18 (Characteristic region of $(\varepsilon,\delta)$-QDP).
for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, we define the characteristic region of $(\varepsilon,\delta)$-DP as follows,
$\mathcal{R}(\varepsilon,\delta):=\left\{(\alpha,\beta)\in[0,1]^{2}:~~\begin{split}\beta&\geq e^{-\varepsilon}(1-\delta-\alpha),\\ \beta&\geq 1-\delta-e^{\varepsilon}\alpha,\\ \beta&\leq 1-e^{-\varepsilon}(\alpha-\delta),\\ \beta&\leq e^{\varepsilon}(1-\alpha)+\delta.\end{split}\right\}.$ (36)
In Figure 1 above, we illustrate a graphical representation of the characteristic region of $(\varepsilon,\delta)$-DP. In [19], the authors studied a characteristic region for classical $(\varepsilon,\delta)$-DP. However, they only considered the region below $\alpha+\beta=1$. This is because, in Figure 1, the characteristic region $\mathcal{R}(\varepsilon,\delta)$ is symmetric over the straight line $\alpha+\beta=1$. Hirche et al. in [32] plotted $\mathcal{R}(\varepsilon,\delta)$ for certain values of $\varepsilon,\delta$.
From eqs.˜32, 33, 34 and 35, it is trivial to see that for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$), we have $\mathcal{R}(\rho,\sigma)\subseteq\mathcal{R}(\varepsilon,\delta)$. Moreover, the characteristic region $\mathcal{R}(\varepsilon,\delta)$ is a convex set. This is because for any two positive operators $0\preceq\Lambda_{1},\Lambda_{2}\preceq\mathbb{I}$, for any $\theta\in[0,1]$, the operator $\Lambda=\theta\Lambda_{1}+(1-\theta)\Lambda_{2}$ is also a positive operator satisfying $0\preceq\Lambda\preceq\mathbb{I}$. Therefore, it can be verified that $\alpha_{\Lambda}=\theta\alpha_{\Lambda_{1}}+(1-\theta)\alpha_{\Lambda_{2}}$ and $\beta_{\Lambda}=\theta\beta_{\Lambda_{1}}+(1-\theta)\beta_{\Lambda_{2}}$, where $(\alpha_{\Lambda_{1}},\beta_{\Lambda_{1}})$ and $(\alpha_{\Lambda_{2}},\beta_{\Lambda_{2}})$ belongs to $\mathcal{R}(\varepsilon,\delta)$. Thus, as $(\alpha_{\Lambda},\beta_{\Lambda})$ also belongs to $\mathcal{R}(\varepsilon,\delta)$, this implies that $\mathcal{R}(\varepsilon,\delta)$ is convex. Further, $\mathcal{R}(\varepsilon,\delta)$ is also a closed set. We will prove this by constructing a pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\in\mathfrak{D_{(\varepsilon,\delta)}}$, which achieves all the corner points of $\mathcal{R}(\varepsilon,\delta)$, as mentioned in Figure 1 above.
Further observe that, in Figure 1, $\mathcal{R}(\varepsilon,\delta)$ has two fixed points (points where $\alpha=\beta$) at $(\frac{1-\delta}{e^{\varepsilon}+1},\frac{1-\delta}{e^{\varepsilon}+1})$ and $(\frac{e^{\varepsilon}+\delta}{e^{\varepsilon}+1},\frac{e^{\varepsilon}+\delta}{e^{\varepsilon}+1})$. The former is the worst fixed point and the latter is the best fixed point of $\mathcal{R}(\varepsilon,\delta)$ from the perspective of privacy. This is because the worst fixed point represents the scenario where the adversary has the most amount of information about the underlying quantum state, i.e., it’s least private, while the best fixed point represents the scenario where the attacker has the least information about the underlying quantum state, i.e., it’s most private.
The characteristic region $\mathcal{R}(\varepsilon,\delta)$ also has six more corner/extremal points at $(0,1-\delta)$, $(1-\delta,0)$, $(\delta,1)$, $(1,\delta)$, $(1,\delta)$ and $(\delta,1)$ along with the two fixed points.
In the discussion below, we will give a constructive proof for the existence of a pair of quantum states $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\in\mathfrak{D}_{(\varepsilon,\delta)}$ such that $\mathcal{R}(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})=\mathcal{R}(\varepsilon,\delta)$ for any given $\varepsilon\geq 0$ and $\delta\in[0,1]$ i.e. we show that for $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$, under certain choices of measurements, the pair of type-I ($\alpha$) and type-II errors ($\beta$) achieves the extremal/corner points of $\mathcal{R}(\varepsilon,\delta)$.
Consider that $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ are two quantum states over a finite-dimensional Hilbert space $\mathcal{H}$ such that for an one-dimensional projector $|v\rangle\langle v|$ we have,
$\mathrm{Tr}[|v\rangle\langle v|\rho_{(\varepsilon,\delta)}]$ $\triangleq\delta,$ (37)
$\mathrm{Tr}[|v\rangle\langle v|\sigma_{(\varepsilon,\delta)}]$ $\triangleq 0.$ (38)
Then, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the corner point $(\delta,1)$ of $\mathcal{R}(\varepsilon,\delta)$. Similarly, for the one-dimensional projector $\mathbb{I}-|v\rangle\langle v|$, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the corner point $(1-\delta,0)$ of $\mathcal{R}(\varepsilon,\delta)$.
Further, consider another one-dimensional projector $|x\rangle\langle x|$ which is perpendicular to $|v\rangle\langle v|$, such that,
$\mathrm{Tr}[|x\rangle\langle x|\rho_{(\varepsilon,\delta)}]\triangleq 0,$ (39)
$\mathrm{Tr}[|x\rangle\langle x|\sigma_{(\varepsilon,\delta)}]\triangleq\delta.$ (40)
Then, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the corner point $(0,1-\delta)$ of $\mathcal{R}(\varepsilon,\delta)$. Similarly, for the one-dimensional projector $\mathbb{I}-|x\rangle\langle x|$, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the corner point $(1,\delta)$ of $\mathcal{R}(\varepsilon,\delta)$.
Now, we notice that the cardinality of the support of the state pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ must be at least $3$ since we need two more corner points to be achieved.
Thus, we consider another one-dimensional projector $\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|$ which is perpendicular to both $|v\rangle\langle v|$ and $|x\rangle\langle x|$. Therefore, from the above discussions, we have,
$\mathrm{Tr}[\left(\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|\right)\rho_{(\varepsilon,\delta)}]$ $=1-\delta,$ (41)
$\mathrm{Tr}[\left(\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|\right)\sigma_{(\varepsilon,\delta)}]$ $=1-\delta.$ (42)
Thus, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the point $(1-\delta,\delta)$ of $\mathcal{R}(\varepsilon,\delta)$. However, this is an interior point of $\mathcal{R}(\varepsilon,\delta)$ and not a corner point. Thus, it can be easily verified that combining the above one-dimensional projectors, we can never achieve the two fixed points of $\mathcal{R}(\varepsilon,\delta)$. Therefore, we consider the support of the state pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ to be at least $4$.
We now consider another one-dimensional projector $|y\rangle\langle y|$ which is perpendicular to both $|v\rangle\langle v|$ and $|x\rangle\langle x|$, and we consider the two-dimensional projector $(|v\rangle\langle v|+|y\rangle\langle y|)$. For this projector, we have two choices: we can either choose it to achieve the worst fixed point, i.e., $(\frac{1-\delta}{1+e^{\varepsilon}},\frac{1-\delta}{1+e^{\varepsilon}})$ or the best fixed point, i.e., $(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}})$. We will see that the former choice will lead us to some construction of the pair of quantum states, which achieves all the corner points of $\mathcal{R}(\varepsilon,\delta)$ but fails to satisfy $(\varepsilon,\delta)$-DP condition. See Remark 1 below for more details. Thus, we choose the two-dimensional projector $(|v\rangle\langle v|+|y\rangle\langle y|)$ to achieve the best fixed point. Towards this, we assume that
$\mathrm{Tr}[(|v\rangle\langle v|+|y\rangle\langle y|)\rho_{(\varepsilon,\delta)}]$ $\triangleq\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},$ (43)
$\mathrm{Tr}[\left(\mathbb{I}-(|v\rangle\langle v|+|y\rangle\langle y|)\right)\sigma_{(\varepsilon,\delta)}]$ $\triangleq\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}.$ (44)
Thus, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the best fixed point $(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}})$ of $\mathcal{R}(\varepsilon,\delta)$ with the help of the two-dimensional projector $|v\rangle\langle v|+|y\rangle\langle y|$.
Now eqs.˜37, 38, 43 and 44 gives us,
$\mathrm{Tr}[|y\rangle\langle y|\rho_{(\varepsilon,\delta)}]$ $=\frac{e^{\varepsilon}(1-\delta)}{1+e^{\varepsilon}},$ (45)
$\mathrm{Tr}[|y\rangle\langle y|\sigma_{(\varepsilon,\delta)}]$ $=\frac{1-\delta}{1+e^{\varepsilon}}.$ (46)
Finally, consider the one-dimensional projector $|z\rangle\langle z|=\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|-|y\rangle\langle y|$ which is perpendicular to $|v\rangle\langle v|$, $|x\rangle\langle x|$ and $|y\rangle\langle y|$. Then, from eqs.˜37, 38, 39, 40, 45 and 46 we have,
$\mathrm{Tr}[|z\rangle\langle z|\rho_{(\varepsilon,\delta)}]$ $=\frac{1-\delta}{1+e^{\varepsilon}},$ (47)
$\mathrm{Tr}[|z\rangle\langle z|\sigma_{(\varepsilon,\delta)}]$ $=\frac{e^{\varepsilon}(1-\delta)}{1+e^{\varepsilon}}.$ (48)
Thus, one can verify that the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the point $(\frac{1-\delta}{1+e^{\varepsilon}},\frac{1-\delta}{1+e^{\varepsilon}})$ of $\mathcal{R}(\varepsilon,\delta)$ with the help of the two-dimensional projector $|x\rangle\langle x|+|z\rangle\langle z|$.
Therefore, we have constructed a pair of quantum states $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ which can be written as follows,
$\rho_{(\varepsilon,\delta)}$ $=\delta|v\rangle\langle v|+\frac{e^{\varepsilon}(1-\delta)}{1+e^{\varepsilon}}|y\rangle\langle y|+\frac{1-\delta}{1+e^{\varepsilon}}|z\rangle\langle z|+0|x\rangle\langle x|,$ (49)
$\sigma_{(\varepsilon,\delta)}$ $=0|v\rangle\langle v|+\frac{1-\delta}{1+e^{\varepsilon}}|y\rangle\langle y|+\frac{e^{\varepsilon}(1-\delta)}{1+e^{\varepsilon}}|z\rangle\langle z|+\delta|x\rangle\langle x|.$ (50)
Further, one can verify that the pair of quantum states $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\in\mathfrak{D}_{(\varepsilon,\delta)}$ where $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ are defined in (49) and (50) and satisfies all the corner points of $\mathcal{R}(\varepsilon,\delta)$. Hence, we have $\mathcal{R}(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})=\mathcal{R}(\varepsilon,\delta)$. Thus, from the existence of $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ and the convexity, we observe that $\mathcal{R}(\varepsilon,\delta)$ is a closed convex set.
To simplify the definition of $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$, we consider the Hilbert space $\mathcal{H}$ to be four-dimensional with the orthonormal basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Thus, we can rewrite $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ as follows,
$\rho_{(\varepsilon,\delta)}$ $=\delta|00\rangle\langle 00|+(1-\delta)\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{1}{1+e^{\varepsilon}}|10\rangle\langle 10|\right),$ (51)
$\sigma_{(\varepsilon,\delta)}$ $=(1-\delta)\left(\frac{1}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|10\rangle\langle 10|\right)+\delta|11\rangle\langle 11|.$ (52)
Remark 1.
  • It is important to note that in (43) and (44), if we had chosen the RHS to be $(\frac{1-\delta}{1+e^{\varepsilon}}$ and $\frac{1-\delta}{1+e^{\varepsilon}}$ we would get another pair of quantum states $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ which also achieves the extremal points of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ for any given $\varepsilon\geq 0$ and $\delta\in[0,1]$. This pair is given as follows,
    $\rho^{\prime}_{(\varepsilon,\delta)}$ $=\delta|00\rangle\langle 00|+\frac{1-2\delta-e^{\varepsilon}\delta}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}|10\rangle\langle 10|,$ (53)
    $\sigma^{\prime}_{(\varepsilon,\delta)}$ $=\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{1-2\delta-e^{\varepsilon}\delta}{1+e^{\varepsilon}}|10\rangle\langle 10|+\delta|11\rangle\langle 11|.$ (54)
    However, the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ also achieves two points $\left(\frac{1-\delta}{1+e^{\varepsilon}}-\delta,\frac{1-\delta}{1+e^{\varepsilon}}-\delta\right)$ and $\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta,\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta\right)$, which are outside the characteristic region $\mathcal{R}(\varepsilon,\delta)$. From this, we can conclude that the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})\not\in\mathfrak{D}_{(\varepsilon,\delta)}$.
  • Further, one can verify that the above pair satisfies $(\log(\frac{e^{\varepsilon}}{1-\delta(2+e^{\varepsilon})}),\delta)$-DP.
In Figure 2 below, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ along with the extremal points achieved by the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$.
$\alpha\longrightarrow$
(Type-I error)
$\beta\longrightarrow$
(Type-II error)
[…]
$\left(\frac{1-\delta}{1+e^{\varepsilon}}-\delta,\frac{1-\delta}{1+e^{\varepsilon}}-\delta\right)$
(Point outside $\mathcal{R}(\varepsilon,\delta)$)
$\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta,\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta\right)$
(Point outside $\mathcal{R}(\varepsilon,\delta)$)
Figure 2: Graphical representation of $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ (in the whole shaded region) which is strictly larger than $\mathcal{R}(\varepsilon,\delta)$ (in inner shaded region), where the two extremal points of $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ outside $\mathcal{R}(\varepsilon,\delta)$ are $\left(\frac{1-\delta}{1+e^{\varepsilon}}-\delta,\frac{1-\delta}{1+e^{\varepsilon}}-\delta\right)$ and $\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta,\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta\right)$.
V Differential Privacy : A Quantum Hypothesis Testing Perspective
V-A Main characterization
In this section, we consider differential privacy from a quantum hypothesis testing perspective. The main goal of quantum differential privacy is to ensure that a quantum mechanism produces output states that are nearly indistinguishable when applied to neighboring quantum inputs. Formally, for two neighboring quantum inputs producing output states $\rho,\sigma$, the distinguishability between $\rho$ and $\sigma$ must be limited by the privacy parameters. Intuitively, this means that even if an investigator (adversary) knows the mechanism and observes its output, the investigator should not be able to reliably infer which quantum input (respondent) generated it. More formally, with high probability, no hypothesis test—regardless of the investigator’s strategy—can reliably infer the respondent’s individual contribution from the output.
Further, it is important to observe that given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, if for any $\alpha\in[0,1]$ the hypothesis testing divergence satisfies if $D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})}$, then $(\sigma_{1},\sigma_{2})$ is harder to distinguish than $(\rho_{1},\rho_{2})$ for all possible values of the Type-I error parameter $\alpha$. Intuitively, this indicates that the states $\rho_{1}$ and $\rho_{2}$ are more distinguishable—or equivalently, more well-separated—than the states $\sigma_{1}$ and $\sigma_{2}$. This observation naturally leads to the following result.
Theorem 1.
Consider two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, for which the following holds,
$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})},\forall\alpha\in[0,1].$ (55)
[…]
Proof.
[…]
Motivated by the close relation between privacy and hypothesis testing in the lemma below, we give an equivalent condition for a pair of quantum states to be $(\varepsilon,\delta)$-DP.
Lemma 3.
A pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 16) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1)$ if and only if the following holds,
$D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1].$ (56)
where $f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1]$.
Proof.
See Appendix -B for the proof.
Note that the above version of quantum $(\varepsilon,\delta)$-DP implies that if for each $\alpha\in[0,1]$ $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}(\alpha)$ for some pair of quantum states $\rho^{\prime},\sigma^{\prime}$ over any arbitrary finite dimension Hilbert space, then for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 16) is at least as hard as distinguishing between the pair $(\rho^{\prime},\sigma^{\prime})$. The above intuition provides a notion of the weakest (most informative) quantum $(\varepsilon,\delta)$-DP pairs of quantum states, which are the least “hard to differentiate”. We formalize this notion in the definition below.
Definition 19 (Weakest quantum $(\varepsilon,\delta)$-DP).
For $\varepsilon\geq 0$ and $\delta\in[0,1),$ a pair of quantum states $(\rho,\sigma)$ is defined to be the weakest (most informative) quantum $(\varepsilon,\delta)$-DP if
$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1].$ (57)
where $f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1]$.
[…]
Theorem 2.
[…]
where $(\rho^{\prime},\sigma^{\prime})$ is some pair of quantum states which is the weakest (most informative) $(\varepsilon,\delta)$-DP (see Definition 19).
Proof.
If a pair $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then, from Lemma 3 and the fact that $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}$ for every $\alpha\in[0,1]$, the following holds,
$D^{\alpha}_{H}(\rho\|\sigma)\leq D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime}),\forall\alpha\in[0,1].$ (59)
[…]
In the lemma below, we show the existence of a pair of quantum states that is the weakest (most informative) quantum $(\varepsilon,\delta)$-DP (see Definition 19) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$ in the Blackwell sense.
Lemma 4.
Consider the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in (51) and (52), for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$. Then, we have $D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=-\log\max\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\}$ for any $\alpha\in[0,1]$.
Proof.
See Appendix -C for the proof.
V-B Blackwell Dominance under Quantum Differential Privacy
From Lemmas 3 and 4 and Definition 15, we note that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, gets dominated by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$. Formally, it implies the following.
Corollary 1.
If any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then the following holds,
$\left\{(\rho,\sigma)\right\}\preceq_{B_{Q}}\left\{(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\right\}.$ (60)
Remark 2.
  • Note that the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in Lemma 4 satisfies weakest (most informative) $(\varepsilon,\delta)$-DP condition for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$ and therefore, they have a fixed-point at $\alpha=\frac{1-\delta}{e^{\varepsilon}+1}$.
  • For $\delta=0$ and any $\varepsilon\geq 0$, we define two quantum states $\rho_{\varepsilon}:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|$ and $\sigma_{\varepsilon}:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|$. Observe that, it can be verified that $\rho_{\varepsilon}$ and $\sigma_{\varepsilon}$ are weakest (most informative) $(\varepsilon,0)$-DP states.
[…]
Corollary 2.
[…]
The reference [19, Theorem 18] showed Corollary 2 in the classical case.
Corollary 3.
[…]
where $\rho_{(\varepsilon)}:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|$ and $\sigma_{(\varepsilon)}:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|$.
Now using Corollary 2, for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, we can derive upper-bounds in terms of quantum hockey-stick divergence. However, it is important to note that for $\delta>0$, $\operatorname{supp}(\rho_{(\varepsilon,\delta)})\not\subseteq\operatorname{supp}(\sigma_{(\varepsilon,\delta)})$. Therefore, the divergence that requires support inclusion is not well-defined for the pair of quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$. Although for $\delta=0$, using Corollary 3, since $\operatorname{supp}(\rho_{\varepsilon})=\operatorname{supp}(\sigma_{\varepsilon})$, for any $\varepsilon,0$-DP quantum state pairs, we can derive upper-bounds on quantum relative entropy, quantum Rényi divergence in the corollary below.
Corollary 4.
If a pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$ for some fixed $\varepsilon\geq 0$, then, we have the following upper-bounds,
  1. (i)
    $D(\rho\|\sigma)\leq\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)$,
  2. (ii)
    $D_{\alpha}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\left[\log\left(e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}\right)-\log\left(1+e^{\varepsilon}\right)\right]$,  for all $\alpha\in[0,1]\cup(1,2)$ ,
  3. (iii)
    ${\left\lVert\rho-\sigma\right\rVert}_{1}\leq\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)$.
Proof.
(i) From Fact 11 it follows that $D(\cdot\|\cdot)$ is a valid divergence and thus, we can write the following,
$D(\rho^{\prime}\|\sigma^{\prime})$ $=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\varepsilon-\frac{1}{1+e^{\varepsilon}}\varepsilon$
$=\varepsilon\left(\frac{e^{\frac{\varepsilon}{2}}-e^{-\frac{\varepsilon}{2}}}{e^{\frac{\varepsilon}{2}}+e^{-\frac{\varepsilon}{2}}}\right)$
$=\varepsilon\tanh\left(\frac{\varepsilon}{2}\right).$ (63)
Therefore, (61) of Corollary 2 completes the proof of 1) of Corollary 4.
(ii) From Fact 13 it follows that for $\alpha\in[0,1]\cup(1,2]$, $D_{\alpha}(\cdot\|\cdot)$ is a valid divergence and thus, we can write the following,
$D_{\alpha}(\rho^{\prime}\|\sigma^{\prime})$ $=\frac{1}{\alpha-1}\log\left(\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\right)^{\alpha}\left(\frac{1}{1+e^{\varepsilon}}\right)^{1-\alpha}+\left(\frac{1}{1+e^{\varepsilon}}\right)^{\alpha}\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\right)^{1-\alpha}\right)$
$=\frac{1}{\alpha-1}\log\left(\frac{e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}}{1+e^{\varepsilon}}\right)$
$=\frac{1}{\alpha-1}\left[\log\left(e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}\right)-\log\left(1+e^{\varepsilon}\right)\right].$ (64)
Therefore, (61) of Corollary 2 completes the proof of 2) of Corollary 4.
(iii) As ${\left\lVert\mathcal{N}(\rho)-\mathcal{N}(\sigma)\right\rVert}_{1}=2E_{1}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))$, from Fact 14 it follows that ${\left\lVert\cdot-\cdot\right\rVert}_{1}$ is a valid divergence and thus, we can write the following,
${\left\lVert\rho^{\prime}-\sigma^{\prime}\right\rVert}_{1}$ $=\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|+\left|\frac{1-e^{\varepsilon}}{1+e^{\varepsilon}}\right|$
$\overset{a}{=}2\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|$
$=2\tanh\left(\frac{\varepsilon}{2}\right).$ (65)
Therefore, (61) of Corollary 2 completes the proof of 3) of Corollary 4.
VI Quantum Privatized Parameter Inference
In the context of randomized response, a fundamental challenge is balancing the privacy of respondents with the statistical utility of the information gathered by an investigator. Let’s consider a scenario where binary information, represented by $\{0,1\}$, is encoded into two quantum states, $\rho$ and $\sigma$. The privacy of this encoding is quantified by the $(\varepsilon,\delta)$-DP constraint on the pair $(\rho,\sigma)$.
We assume that the behavior of the respondents is considered as random sampling with replacement from the investigator’s perspective, and the investigator is interested only in the ratio between 0 and $1$ so that the underlying statistical information $\theta$ follows a binomial distribution $p_{\theta}$, where $p_{\theta}(0)=\theta$ and $p_{\theta}(1)=1-\theta$. That is, the key task is the inference of the parameter $\theta$ by the investigator.
With the quantum encoding, this translates to a parametrized state $\rho_{\theta}:=\theta\rho+(1-\theta)\sigma$, where $\theta\in[0,1]$. If the investigator collects $n$ responses, the composite state is $\rho_{\theta}^{\otimes n}$.
More generally, in quantum parameter estimation, the goal is to estimate an unknown parameter $\theta$ encoded in a family of quantum states $\{\rho_{\theta}\}$, where the encoding is constrained by privacy requirements. In the privatized setting, the states $\rho$ and $\sigma$ must satisfy the $(\varepsilon,\delta)$-DP condition, which restricts the distinguishability of the states and thus limits the amount of information that can be extracted about $\theta$.
The estimation performance is fundamentally limited by the quantum Fisher information, which quantifies the sensitivity of the state $\rho_{\theta}$ to changes in $\theta$. In the quantum setting, the relevant quantity is the Symmetric Logarithmic Derivative (SLD) Fisher information, defined as
$J_{\theta}:=\mathrm{Tr}\left[L_{\theta}^{2}\rho_{\theta}\right],$
where $L_{\theta}$ is the SLD operator satisfying
$\frac{1}{2}(L_{\theta}\rho_{\theta}+\rho_{\theta}L_{\theta})=\frac{d\rho_{\theta}}{d\theta}.$
The quantum Cramér-Rao bound states that the mean squared error (MSE) of any unbiased estimator $\hat{\theta}$ is lower bounded by $1/J_{\theta}$.
In the privatized scenario, the investigator is restricted to the pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$, and thus the achievable Fisher information is maximized over all such admissible pairs. This leads to a constrained optimization problem,
$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}J_{\theta},$
where the maximum is taken over all pairs of quantum states satisfying the privacy constraint. The optimal value quantifies the fundamental tradeoff between privacy and estimation accuracy in quantum settings.
Furthermore, the structure of the optimal $(\varepsilon,\delta)$-DP pair, as characterized in Lemma 4, allows for explicit computation of the Fisher information and the corresponding Cramér-Rao bound. This provides a precise benchmark for the best possible estimation performance under quantum differential privacy constraints, and generalizes classical results to the quantum regime.
The investigator’s goal is to estimate the parameter $\theta$. The quality of this estimation is typically measured by the mean square error (MSE). For a single-parameter model, the MSE is lower-bounded by the Cramér-Rao bound, which is the inverse of the Fisher information. In the quantum setting, we use the Symmetric Logarithmic Derivative (SLD) Fisher information. The SLD, denoted $L_{\theta}$, is defined by the equation:
$\frac{1}{2}(L_{\theta}\rho_{\theta}+\rho_{\theta}L_{\theta})=\frac{d\rho_{\theta}}{d\theta}=\rho-\sigma.$ (66)
The SLD Fisher information, $J_{\theta}$, is then given by:
$J_{\theta}:=\mathrm{Tr}(L_{\theta}^{2}\rho_{\theta}).$ (67)
The MSE of any unbiased estimator for $\theta$ is lower-bounded by $1/J_{\theta}$. This bound is asymptotically achievable, for instance, via a two-step estimation process. Consequently, to optimize the estimation, it is natural to maximize the SLD Fisher information $J_{\theta}$ subject to the $(\varepsilon,\delta)$-DP privacy constraint on the states $\rho$ and $\sigma$.
Theorem 3.
The maximum SLD Fisher information achievable under the $(\varepsilon,\delta)$-DP constraint is given by:
$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}J_{\theta}=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$ (68)
This maximum is uniformly attained by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
Proof.
Let $\rho_{(\varepsilon,\delta),\theta}:=\theta\rho_{(\varepsilon,\delta)}+(1-\theta)\sigma_{(\varepsilon,\delta)}$. We first compute the SLD Fisher information for this specific state, which we denote by $J_{(\varepsilon,\delta),\theta}$. The state $\rho_{(\varepsilon,\delta),\theta}$ is diagonal in the standard basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Its eigenvalues are:
$p_{00}$ $=\theta\delta,$
$p_{01}$ $=\theta\frac{(1-\delta)e^{\varepsilon}}{1+e^{\varepsilon}}+(1-\theta)\frac{1-\delta}{1+e^{\varepsilon}}=\frac{1-\delta}{1+e^{\varepsilon}}(\theta e^{\varepsilon}+1-\theta),$
$p_{10}$ $=\theta\frac{1-\delta}{1+e^{\varepsilon}}+(1-\theta)\frac{(1-\delta)e^{\varepsilon}}{1+e^{\varepsilon}}=\frac{1-\delta}{1+e^{\varepsilon}}(\theta+(1-\theta)e^{\varepsilon}),$
$p_{11}$ $=(1-\theta)\delta.$
The derivative $\frac{d\rho_{(\varepsilon,\delta),\theta}}{d\theta}=\rho_{(\varepsilon,\delta)}-\sigma_{(\varepsilon,\delta)}$ is also diagonal, with eigenvalues:
$d_{00}$ $=\delta,$
$d_{01}$ $=\frac{(1-\delta)(e^{\varepsilon}-1)}{1+e^{\varepsilon}},$
$d_{10}$ $=-\frac{(1-\delta)(e^{\varepsilon}-1)}{1+e^{\varepsilon}},$
$d_{11}$ $=-\delta.$
For diagonal states, the SLD Fisher information is given by the sum of the classical Fisher informations for the eigenvalues: $J_{\theta}=\sum_{i}\frac{(\frac{dp_{i}}{d\theta})^{2}}{p_{i}}$. In this case, $\frac{dp_{i}}{d\theta}=d_{i}$. Thus, we have,
$J_{(\varepsilon,\delta),\theta}$ $=\frac{d_{00}^{2}}{p_{00}}+\frac{d_{01}^{2}}{p_{01}}+\frac{d_{10}^{2}}{p_{10}}+\frac{d_{11}^{2}}{p_{11}}$
$=\frac{\delta^{2}}{\theta\delta}+\frac{\left(\frac{(1-\delta)(e^{\varepsilon}-1)}{1+e^{\varepsilon}}\right)^{2}}{\frac{1-\delta}{1+e^{\varepsilon}}(\theta e^{\varepsilon}+1-\theta)}+\frac{\left(-\frac{(1-\delta)(e^{\varepsilon}-1)}{1+e^{\varepsilon}}\right)^{2}}{\frac{1-\delta}{1+e^{\varepsilon}}(\theta+(1-\theta)e^{\varepsilon})}+\frac{(-\delta)^{2}}{(1-\theta)\delta}$
$=\frac{\delta}{\theta}+\frac{\delta}{1-\theta}+\frac{(1-\delta)(e^{\varepsilon}-1)^{2}}{1+e^{\varepsilon}}\left(\frac{1}{\theta e^{\varepsilon}+1-\theta}+\frac{1}{\theta+(1-\theta)e^{\varepsilon}}\right)$ (69)
$=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(e^{\varepsilon}-1)^{2}}{1+e^{\varepsilon}}\left(\frac{(\theta+(1-\theta)e^{\varepsilon})+(\theta e^{\varepsilon}+1-\theta)}{(\theta e^{\varepsilon}+1-\theta)(\theta+(1-\theta)e^{\varepsilon})}\right)$
$=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(e^{\varepsilon}-1)^{2}}{1+e^{\varepsilon}}\left(\frac{1+e^{\varepsilon}}{\theta^{2}e^{\varepsilon}+\theta(1-\theta)+\theta(1-\theta)(e^{\varepsilon})^{2}+(1-\theta)^{2}e^{\varepsilon}}\right)$
$=\frac{\delta}{\theta(1-\theta)}+(1-\delta)(e^{\varepsilon}-1)^{2}\left(\frac{1}{\theta^{2}e^{\varepsilon}+\theta(1-\theta)(1+(e^{\varepsilon})^{2})+(1-\theta)^{2}e^{\varepsilon}}\right)$
$=\frac{\delta}{\theta(1-\theta)}+(1-\delta)(e^{\varepsilon}-1)^{2}\left(\frac{1}{e^{\varepsilon}(\theta^{2}+(1-\theta)^{2})+\theta(1-\theta)(1+(e^{\varepsilon})^{2})}\right)$
$=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$
Now, consider any pair of states $(\rho,\sigma)$ that satisfies the $(\varepsilon,\delta)$-DP condition. From Corollary 1, there exists a CP-TP map $\Gamma$ such that $\Gamma(\rho_{(\varepsilon,\delta)})=\rho$ and $\Gamma(\sigma_{(\varepsilon,\delta)})=\sigma$. From Fact 15, the SLD Fisher information preserves monotonicity under CP-TP maps. Therefore, since $J_{\theta}$ is the SLD Fisher information for the state $\rho_{\theta}=\theta\rho+(1-\theta)\sigma$, Corollary 2 implies
$J_{\theta}\leq J_{(\varepsilon,\delta),\theta}.$ (70)
Since the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ itself satisfies the $(\varepsilon,\delta)$-DP condition, it is a valid choice for the maximization. Thus, the equality in (70) is achievable, and the maximum value is $J_{(\varepsilon,\delta),\theta}$.
As an alternative application, consider the scenario where the investigator aims to distinguish between two original information sources, $\rho_{\theta_{0}}$ and $\rho_{\theta_{1}}$. The performance in this hypothesis testing task is characterized by quantities like the hypothesis testing divergence $D_{H}^{\alpha}(\rho_{\theta_{0}}^{\otimes n}\|\rho_{\theta_{1}}^{\otimes n})$.
Also, Corollary 2 implies the following theorem.
Theorem 4.
The maximum hypothesis testing divergence under the $(\varepsilon,\delta)$-DP constraint is given by:
$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}D_{H}^{\alpha}(\rho_{\theta_{0}}^{\otimes n}\|\rho_{\theta_{1}}^{\otimes n})=D_{H}^{\alpha}(\rho_{(\varepsilon,\delta),\theta_{0}}^{\otimes n}\|\rho_{(\varepsilon,\delta),\theta_{1}}^{\otimes n}),$ (71)
This maximum is uniformly attained by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
The asymptotic behavior of the hypothesis testing problem for the optimal states is characterized by the quantum Rényi divergences between $\rho_{(\varepsilon,\delta),\theta_{0}}$ and $\rho_{(\varepsilon,\delta),\theta_{1}}$. The specific expressions for $D(\rho_{(\varepsilon,\delta),\theta_{0}}\|\rho_{(\varepsilon,\delta),\theta_{1}})$ and $D_{\alpha}(\rho_{(\varepsilon,\delta),\theta_{0}}\|\rho_{(\varepsilon,\delta),\theta_{1}})$ can be computed from their definitions.
VII Stability of Learning Algorithms
In the previous setting, the investigator receives information directly from the respondents. However, when there exists a trusted third party, the data processor, who applies the learning algorithm to the training data sent by the respondents between the respondents and the investigator, it is sufficient to discuss the output of the learning algorithm for the privacy. Under this assumption, we can relax our condition for privacy.
To discuss this case, we use the framework developed earlier to study a fundamental problem in learning theory: the privacy of training data. Specifically, we ask how much information the output of a learning algorithm reveals about its training data. Clearly, if the output were completely independent of the training data, it would reveal no information. However, such an algorithm would be useless for learning because it could not leverage the data to make predictions. Thus, there is an inherent trade-off between preserving privacy and maintaining the utility of the learning algorithm.
To address this trade-off, a private learning algorithm should ensure that its output does not change significantly when a single training example is modified. This property is known as stability, which was first introduced in [1]. Formally, stability refers to the insensitivity of the algorithm’s output to small changes in the training dataset, such as replacing one data point with a neighboring one. The authors in [1] studied this in terms of $L_{1}$ distance. In the classical setting, this notion is well established, and differentially private learning algorithms are known to provide strong stability guarantees [33].
In [6], the authors studied the stability of $\varepsilon$-DP learning algorithms and obtained an upper-bound on the mutual information between the training data at the respondent’s end and the algorithm’s output at the investigator’s end.
The main aim of the subsequent subsections is to generalize the [6, Proposition 2] in the quantum setting for $\delta\neq 0$. Towards this, we first discuss a differentially private quantum learning framework in the subsection below.
VII-A Framework for $1$-neighbor $(\varepsilon,\delta)$-DP quantum learning algorithms
Motivated by the learning frameworks of [14] and [15], we will assume that the input to the quantum learning algorithm is a classical $n$-length ($n\in\mathbb{N}$) data $s:=(z_{1},\cdots,z_{n})\in\mathcal{S}:=\mathcal{Z}^{n}$ (where $\left|\mathcal{Z}\right|>0$) and a $\rho_{s}\in\mathcal{D}((\mathbb{C}^{d})^{\otimes n})$. For each $i\in[n]$, $z_{i}:=(x_{i},y_{i})$, where $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ for some non-empty sets $\mathcal{X}$ and $\mathcal{Y}$ and $y_{i}=f(x_{i})$ for some concept function $f:\mathcal{X}\to\mathcal{Y}$, which is unknown to the the learning algorithm.
In a typical case, each data $z_{i}$ for $i\in[n]$ has the form $z_{i}:=(x_{i},y_{i})$, where $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ for some non-empty sets $\mathcal{X}$ and $\mathcal{Y}$ and $y_{i}=f(x_{i})$ for some concept function $f:\mathcal{X}\to\mathcal{Y}$, which is unknown to the the data processor. In this quantum learning framework, the algorithm’s objective is to learn a concept function $f$ from a classical sequence $s$ and a quantum state $\rho_{s}$. Thus, the manuscripts [14] and [15] formulated the input to a learning algorithm to be a classical-quantum state $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$. The quantum mechanics formalism restricts the algorithm to only use measurements and CP-TP maps. Therefore, the joint state between the training data $S$ and the output of any learning algorithm in this framework can be modeled by a classical-quantum state,
$\sigma^{SB}:=\sum_{s\in\mathcal{S}}P_{S}(s)|s\rangle\langle s|\otimes\sigma^{B}_{s},$ (72)
where,
$\sigma^{B}_{s}:=\sum_{w\in\mathcal{W}}(\mathcal{N}_{s,w}(\rho_{s}))^{B^{\prime}}\otimes|w\rangle\langle w|^{W},$ (73)
and $w$ is a hypothesis in the hypothesis class $\mathcal{W}$. Further, the map $\mathcal{N}_{s}:\rho_{s}\to\sum_{w\in\mathcal{W}}\mathcal{N}_{s,w}(\rho_{s})\otimes|w\rangle\langle w|$ is a quantum instrument, i.e., for every $(s,w)\in\mathcal{S}\times\mathcal{W},$ $\mathcal{N}_{s,w}$ is a completely positive trace non-increasing map. In the discussions below for every $s\in\mathcal{S},$ we will use following the notation
$\mathcal{N}_{s}(\rho_{s}):=\sum_{w\in\mathcal{W}}\mathcal{N}_{s,w}(\rho_{s})\otimes|w\rangle\langle w|.$
From the learning point of view, it is quite natural to expect that the order in which the training data is fed to the algorithm should not change the algorithm’s output. Therefore, we will assume that if $s$ and $s^{\prime}$ have the same type as $T_{s},$ then both $\rho_{s}$ and $\rho_{s^{\prime}}$ get mapped to the same state at the output of the learning algorithm. Formally, upon getting an input $\{s,\rho_{s}\}$, the algorithm maps, $\rho_{s}\to V_{\pi_{s}}\rho_{s}V^{\dagger}_{\pi_{s}}$, where $\pi_{s}:s\to T_{s}$ permutes $s$ to $T_{s}$ which is the type representative of $s$. Thus, for every $s$ in a type $T_{s},$ its associated quantum output is $\sum_{w\in\mathcal{W}}\mathcal{N}_{T_{s},w}(\rho_{T_{s}})\otimes|w\rangle\langle w|.$
This property is also desirable from the viewpoint of the privacy because this property disables the investigator to identify which respondent has the respective data $z$. Therefore, we consider that this property is a part of the conditions for the privacy as follows.
Definition 20.
An algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$ is said to be a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm if and only if the following properties are true.
  1. 1.
    For every $s\in\mathcal{S}$ the quantum instrument $\mathcal{N}_{s}$ should be 11-neighbor $(\varepsilon,\delta)$-DP, i.e., for every $s\overset{1}{\sim}s^{\prime}$ (see Definition 3), the following holds,
    $\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]+\delta,$ (74)
    $\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]+\delta,$ (75)
    for every $0\preceq\Lambda\preceq\mathbb{I}.$ The property desired in (74) and (75) is to ensure privacy during the learning process.
  2. 2.
    It should be support consistent. Formally, for each $s\overset{1}{\sim}s^{\prime}\in\mathcal{S}$, the following holds,
    $\operatorname{supp}(\mathcal{N}_{s}(\rho_{s}))=\operatorname{supp}(\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})).$ (76)
    The property mentioned in (76) is restrictive and may not be true for many learning algorithms. However, many of the information-theoretic quantities which can be used to analyze the stability require this condition. Therefore, in light of this requirement, we make the above assumption. Further, for $\delta=0$, this condition is inherently there. Otherwise, for no finite $\varepsilon\geq 0$ the pair of quantum states will satisfy the $\varepsilon$-DP criterion.
A property which follows from the Definition 20 is that if a learning algorithm $\mathcal{A}$ is $1$-neighbor $(\varepsilon,\delta)$-DP, then it is also a differentially private algorithm with respect to data points which are $k$-neighbors. However, the privacy parameters will degrade with $k.$ We formalize this in the corollary below.
Corollary 5.
If a learning algorithm $\mathcal{A}$ satisfies Definition 20, then, any pair of $k$-neighbor (where $k>1$) inputs $s\overset{k}{\sim}s^{\prime}$ and for every $0\preceq\Lambda\preceq\mathbb{I}$, the following holds,
$\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]+g_{k}(\delta),$ (77)
$\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]+g_{k}(\delta),$ (78)
where,
$g_{k}(\varepsilon,\delta):=\frac{e^{k\varepsilon}-1}{e^{\varepsilon}-1}\delta,$ (79)
and is assumed to be less than $1.$
Proof.
See Appendix -F for the proof.
Remark 3.
It is important to note that in this section, our focus is not on the training process of the learning algorithm itself. Instead, we are primarily concerned with the privacy preservation between the input training data and the output of the quantum learning algorithm. We analyze how the $(\varepsilon,\delta)$-DP property of the algorithm limits the information an investigator can infer about the training data from the algorithm’s output.
VII-B Stability of a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent quantum learning algorithm
From the point of view of privacy-preserving learning, the output of a learning algorithm should reveal minimal information about the training dataset it was trained on. In this scenario, the trusted party, the data processor, has access to the learning algorithm and the input state $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$, and the investigator, who has access to the output subsystem $B$ and may be interested in learning about $S$ from $B$. The investigator attempts to infer as much information as possible about the respondent’s original training data $S$ from the algorithm’s output $B$. In contrast, the data processor’s goal is to ensure that the learning process discloses little to no information about the input dataset. Figure 3 below illustrates this framework.
[…]
Figure 3: Privacy based learning framework.
In the context of the above adversarial setting between the respondents and the investigator, motivated from [34], we focus on the mutual information $I\left[S;B\right]$, which is calculated with respect to $\sum_{s\in\mathcal{S}}P_{S}(s)|s\rangle\langle s|\otimes{\mathcal{N}_{s}(\rho_{s})}^{B}$. This quantity satisfies the following chain rule.
$I[S;B]=I[S;B^{\prime}W]=I[S;B^{\prime}\mid W]+I[S;W].$ (80)
The respondents impose the requirement that the output should remain essentially unchanged under the modification of a single training example. This property, called stability, is formalized in the following definition.
Definition 21.
(Stability) A quantum learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s}$ is $\gamma$-stable, if
$\max_{Q_{S}}I\left[S;B\right]\leq\gamma.$
The above definition provides a quantitative upper-bound on the maximum classical information that an adversary or investigator can extract from $B$ about $S$. Consequently, a small upper-bound implies that the algorithm’s output is not strongly dependent on any single training data point, indicating that the algorithm is information-theoretically stable. Therefore, the respondent will always aim to minimize $\gamma$ by using algorithms that are not sensitive to small changes in the input training sequence $s$. Towards this, we have the following definition.
Now, using the framework presented in Section VII-A, we analyze the stability of the quantum private learning algorithm by deriving an upper-bound on the Holevo information in the theorem below, under the assumption that
$g_{n(\left|\mathcal{Z}\right|-1)}(\varepsilon,\delta)=\frac{e^{n(\left|\mathcal{Z}\right|-1)\varepsilon}-1}{e^{\varepsilon}-1}\delta<1.$ (81)
We now have the following theorem.
Theorem 5.
For $\varepsilon\in\left[\frac{1}{n},1)\right$., consider a learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 20 and (81). Then, the following holds,
$I[S;B]_{\sigma}\leq(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (82)
where, $n$ is the length of the training data and for some constant $m\in(0,1],$ $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\log\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ and has a property that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$.
Proof.
See Appendix -D for the proof.
[…]
Corollary 6.
For $\varepsilon\in\left.[0,\frac{1}{n}\right)$, consider a learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 20 and (81). Then, the following holds,
$I[S;B]_{\sigma}\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$ (83)
Corollary 7.
[…]
$I[S;B]_{\sigma}\leq(|\mathcal{Z}|-1)\log\left(n+1\right).$ (84)
The proof of Theorem 5 above relies on the lemma below, which can be thought of as a quantum analog of [6, Lemma $2$].
Lemma 5.
Consider $\rho$ and $\sigma$ be two quantum states over Hilbert space $\mathcal{H}_{A}$ such that $\rho\ll\sigma$ and $\sigma$ is a finite mixture of probability distributions such that $\sigma=\sum_{b=1}^{m}P(b)\sigma_{b}$, where $\sum_{b=1}^{m}P(b)=1$, and $\rho\ll\sigma_{b}$ for all $b\in[m]$. Then, the following holds,
$D(\rho\|\sigma)\leq\min_{b\in[m]}\left\{D\left(\rho\|\sigma_{b}\right)-\log P(b)\right\}.$ (85)
Further, a tighter bound in comparison to (85) is as follows,
$D(\rho\|\sigma)\leq-\log\left(\sum_{b=1}^{m}P(b)\exp\left(-D\left(\rho\|\sigma_{b}\right)\right)\right).$
Proof.
See Appendix -E for the proof.
Theorem 5 establishes a crucial quantitative link between differential privacy and stability by providing an explicit upper-bound on the Holevo information $I[S;B]_{\sigma}$ between the training dataset $S$ and the algorithm output $B$.
Furthermore, the upper-bound given in Theorem 5 is uniform and explicitly depends on dataset size $n$, alphabet size $|\mathcal{Z}|$, and privacy parameters $(\varepsilon,\delta)$. This makes clear how stability scales with these variables. In particular, the theorem translates the $(\varepsilon,\delta)$-privacy guarantees into a provable stability bound, thereby linking differential privacy directly to stability-based generalization controls measured by mutual information. This establishes a precise and quantitative connection between privacy and algorithmic stability.
Theorem 5 allows us to obtain a classical version which can be considered as $(\varepsilon,\delta)$-DP generalization of Proposition 2. Before discussing this result, we first briefly discuss a classical learning framework, which is $1$-neighbor $(\varepsilon,\delta)$-DP.
Remark 4.
The upper-bound obtained in Theorem 5 is independent of $P_{S}$ and thus, Theorem 5 implies that if a quantum learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$ satisfies Definition 20, then $\mathcal{A}$ is $\left((|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+\right$. $\left.h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right)$-stable (see Definition 21). A similar observation also follows for Corollaries 6 and 7.
VII-C $1$-neighbor $(\varepsilon,\delta)$-DP support consistent classical learning algorithms and its stability bound
In the classical setting a learning algorithm takes input $s\in\mathcal{S}$ and produces a $w\in\mathcal{W}$ according to some condition distribution $p_{W\mid s}.$ For more details on $s$ and $w$ see subsection VII-A Further, in the context of learning, it is reasonable to assume that the condition distribution is independent of the order in which the training samples are fed to it (i.e., it depends only on the type of $s$). We say that a learning algorithm is $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm if for every $s\overset{1}{\sim}s^{\prime}$ (see Definition 3), we have
$p_{W\mid s}(\mathcal{J})$ $\leq e^{\varepsilon}p_{W\mid s^{\prime}}(\mathcal{J})+\delta,$
$p_{W\mid s^{\prime}}(\mathcal{J})$ $\leq e^{\varepsilon}p_{W\mid s}(\mathcal{J})+\delta,$ (86)
for $\forall\mathcal{J}\subseteq\mathcal{W}.$
When the learning algorithm $\mathcal{A}=\left\{p_{W\mid s}\right\}_{s\in\mathcal{S}}$ satisfies the above condition for every $s\overset{1}{\sim}s^{\prime},$ $\operatorname{supp}(p_{W\mid s})=\operatorname{supp}(p_{W\mid s^{\prime}})$, Theorem 5 implies the following,
Corollary 8.
Any classical learning algorithm $\mathcal{A}=\left\{p_{W\mid s}\right\}_{s\in\mathcal{S}}$ which satisfies the properties mentioned in (86), satisfies the following upper-bounds,
  1. i)
    for $\varepsilon>1$,
    $I[S;W]\leq(|\mathcal{Z}|-1)\log\left(n+1\right).$
  2. ii)
    for $\frac{1}{n}\leq\varepsilon\leq 1$,
    $I[S;W]\leq(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
  3. iii)
    for $\varepsilon<\frac{1}{n}$,
    $I[S;W]\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
Further, for $\delta=0$, Corollary 8 above, recovers [6, Proposition 22] mentioned below.
Proposition 2 ([6, Proposition 22]).
For $\varepsilon<1$ and $\delta=0$ any classical learning algorithm $\mathcal{A}=\left\{p_{W\mid s}\right\}_{s\in\mathcal{S}}$ which satisfies the properties mentioned in (86), satisfies the following upper-bound,
$I[S;W]\leq(|\mathcal{Z}|-1)\log\big(1+\varepsilon en\big),$
where $n$ is the length of the training data.
VII-D Comparison between Theorem 5 and [20, Proposition 10]
In [20, Proposition 10], the authors claimed to have derived a stability upper-bound on Holevo information for quantum $(\varepsilon,\delta)$-LDP quantum channels. However, it is doubtful and misleading for two reasons.
The first reason is that the applicability of the upper-bound in [20, Proposition 10] to $(\varepsilon,\delta)$-LDP quantum channels is questionable. Although the proposition claims that the bound holds for general $(\varepsilon,\delta)$-LDP channels, the derived inequality is valid only in the special case $\delta=0$. That is, it is evident that the bound can be violated when $\delta$ becomes sufficiently large.
The second and more important reason is that the stability of a learning algorithm is closely related to how sensitive the output of a learning algorithm is with respect to minor changes in the input training data. Even though the authors in [20, Proposition 10] claim to study the stability, the results obtained by them nowhere capture the relation between the stability and the sensitivity. Formally, they have not considered the algorithms which are $(\varepsilon,\delta)$ differentially private quantum channel only for neighboring (defined appropriately) input quantum states.
In contrast, Theorem 5 takes care of all the issues discussed above.
VII-E Comparison between Theorem 5 and [14, Appendix C.7]
In [14, Appendix C.7], the authors obtain an upper-bound on the Holevo information assuming their learning framework under pure $\varepsilon$-LDP constraints. However, they don’t study this from the point of view of the stability of a quantum learning algorithm.
The bound on the Holevo information that they obtain (under pure $\varepsilon$-LDP constraints) is conceptually wrong. This is because of the reasons mentioned below.
  1. 1.
    In their context of the learning algorithm, the $\varepsilon$-LDP channels (or measurements) act only on the training part of the data available. In particular, to obtain [14, Eq. (C.7.1)], the authors assume that $\Lambda^{\mathcal{A}}_{s,w}$, for any $0\preceq M\preceq\mathbb{I}^{\mbox{hyp}}$ and $\rho^{\mbox{train}}_{1},\rho^{\mbox{train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mbox{train}})$, $\Lambda^{\mathcal{A}}_{s,w}$ satisifies the following,
    $\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{1})\right]\leq e^{\varepsilon}\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{2})\right].$
    The authors further claim that even this local action would maintain the $\varepsilon$-LDP globally. However, for [14, Eq. (C.7.1)] to be true, the following must hold,
    $\mathrm{Tr}\left[O\left(\mathbb{I}^{\mbox{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\right)(\rho^{\mbox{test;train}}_{1})\right]\leq e^{\varepsilon}\mathrm{Tr}\left[O\left(\mathbb{I}^{\mbox{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\right)(\rho^{\mbox{test;train}}_{2})\right],$ (90)
    for every $0\preceq O\preceq\mathbb{I}^{\mbox{test;hyp}}$. From [4, Theorem $4$], (90) does not hold because $\mathbb{I}^{\mbox{test}}$ is not a differentially private channel for some fixed $\varepsilon\geq 0$.
    Therefore, the validity of [14, Eq. (C.7.1)] is questionable under the given assumptions and may require further scrutiny.
  2. 2.
    Further, in their framework, they consider the training part of their learning algorithm to consist of certain POVM measurements (which depend on the classical training data) that act on the quantum part of the training data. However, while studying the Holevo information under $\varepsilon$-LDP constraints, they assume that the POVMs are independent of the training data. In particular, on the page $59$ of [14] the authors mention the following,
    “Next, we turn our attention to the classical MI term in our generalization bounds. Here, we assume that the data processor $\mathcal{A}$ uses an overall $\varepsilon$-LDP POVM. As the POVM $\left\{|s\rangle\langle s|\otimes E_{s}^{\mathcal{A}}(w)\right\}_{s,w}$ is not LDP even if every $\left\{E_{s}^{\mathcal{A}}(w)\right\}_{w}$ is, we make the simplifying assumption that the data processor uses an $s$-independent $\varepsilon$-LDP POVM $\left\{E^{\mathcal{A}}(w)\right\}_{w}$.”
    The above over-simplified assumption is as good as saying that the learning algorithm doesn’t depend on the training data. In that case, the algorithm is trivially private. However, it will incur severe errors.
Theorem 5 does not have any such issues.
VIII Upper-Bounds on Relative Entropy between the outputs of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
VIII-A Formulation of $(\varepsilon,\delta)$-LDP classical channels
[…]
In this section, we will obtain a meaningful bound on the relative entropy $D(\cdot\|\cdot)$ between the output distributions induced by an $(\varepsilon,\delta)$-LDP channel (defined below) for the case when at least the support of one of the output distributions is a subset of the support of the other distribution. We will accomplish this with the help of the integral representation of the relative entropy mentioned in Fact 5. This integral representation uses the contraction coefficient of $(\varepsilon,\delta)$-LDP channels for the hockey stick divergence.
Towards this, we first formulate the channel-based setup of $(\varepsilon,\delta)$-LDP in both classical and quantum settings. In the classical case, a channel from the system $\mathcal{X}$ to the system $\mathcal{Y}$ is given as a transition matrix $P_{Y|X}$. Assume that a respondent generates private data in $\mathcal{X}$ and converts it to a distribution on $\mathcal{Y}$ via the channel $P_{Y|X}$ and an investigator can access only the system $\mathcal{Y}$. In this case, a $(\varepsilon,\delta)$-LDP channel is formulated as a generalization of $(\varepsilon,\delta)$-LDP in the following way.
Definition 22.
A classical channel $P_{Y|X}$ is defined to be $(\varepsilon,\delta)$-LDP (locally differentially private) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, if any pair of $x\neq x^{\prime}\in\mathcal{X}$ satisfies
$P_{Y|X}(S|x)\leq e^{\varepsilon}P_{Y|X}(S|x^{\prime})+\delta,$ (91)
for every subset $S\subseteq\mathcal{Y}$. Further, for $\delta=0$, we denote $\mathcal{K}$ to be a pure $\varepsilon$-LDP or just $\varepsilon$-LDP channel.
Now, we write the channel by using the map $\mathcal{K}$ from a distribution on $\mathcal{X}$ to a distribution on $\mathcal{Y}$ as $\mathcal{K}(P)(y):=\sum_{x\in\mathcal{X}}P_{Y|X}(y|x)P(x)$. This definition is rewritten as follows.
Lemma 6.
A classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP, if and only if any pair of distributions $P_{X},Q_{X}\in\mathcal{P}(\mathcal{X})$ satisfies the relation
$\mathcal{K}(P)(S)\leq e^{\varepsilon}\mathcal{K}(Q)(S)+\delta,$ (92)
for every subset $S\subseteq\mathcal{Y}$.
Proof.
When the condition (92) holds, the classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP by considering the case when $P$ and $Q$ are delta distributions.
Assume that the classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP. Given $x\in\mathcal{X}$, (91) implies $P_{Y|X}(S|x)\leq e^{\varepsilon}P_{Y|X}(S|x^{\prime})+\delta$ for any $x^{\prime}\in\mathcal{X}$. Thus, $P_{Y|X}(S|x)=\sum_{x^{\prime}\in\mathcal{X}}Q(x^{\prime})P_{Y|X}(S|x)\leq\sum_{x^{\prime}\in\mathcal{X}}Q(x^{\prime})(e^{\varepsilon}P_{Y|X}(S|x^{\prime})+\delta)=e^{\varepsilon}\mathcal{K}(Q)(S)+\delta$. Then, we have $\mathcal{K}(P)(S)=\sum_{x\in\mathcal{X}}P(x)P_{Y|X}(S|x)\leq=\sum_{x\in\mathcal{X}}P(x)(e^{\varepsilon}\mathcal{K}(Q)(S)+\delta)=e^{\varepsilon}\mathcal{K}(Q)(S)+\delta$, which shows (92).
We note here that the above definition of $(\varepsilon,\delta)$-LDP channel is relaxed version of the standard definition of $(\varepsilon,\delta)$-LDP channel [16], where the privacy condition is required to hold for every pair of input symbols $x,x^{\prime}\in\mathcal{X}$. However, in the upcoming discussion, we will see that the above definition is sufficient to obtain meaningful upper-bounds on the relative entropy (when it is well defined) between the output distributions induced by $(\varepsilon,\delta)$-LDP channels. Now, as a quantum generalization of (92), we have the following definition.
Definition 23 ([35]).
A CP-TP map $\mathcal{N}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ is defined to be quantum $(\varepsilon,\delta)$-LDP (locally differentially private) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, if for all pairs $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and every POVM measurement $0\preceq\Lambda\preceq\mathbb{I}$, the following holds,
$\mathrm{Tr}[\Lambda\mathcal{N}(\rho)]\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\sigma)]+\delta.$ (93)
Further, for $\delta=0$, we denote $\mathcal{N}$ to be a pure quantum $\varepsilon$-LDP or just quantum $\varepsilon$-LDP CP-TP map.
Remark 5.
Observe that Definition 23 appears to be very strict in the sense that for a CP-TP map to be $(\varepsilon,\delta)$-DP, it should behave differentially private for every pair of quantum states. In [20], the authors give an example of one such map in terms of a measurement channel composed by a depolarizing channel.
In the subsections below, we define the contraction coefficient of classical and quantum $(\varepsilon,\delta)$-LDP channels with respect to general divergences and obtain almost matching upper and lower bounds on the contraction coefficient of the hockey stick divergence for $(\varepsilon,\delta)$-LDP mechanism.
VIII-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
For any classical divergence $\mathbb{D}$, from Definition 9, we know that for any pair of probability distributions $P,Q\in\mathcal{P}(\mathcal{X})$ (where $\mathcal{X}$ is some finite set) and a classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ (where $\mathcal{Y}$ is another finite set), $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\mathbb{D}(P_{1}\|P_{2})$, where for each $i=1,2$, $\mathcal{K}(P_{i})$ is the marginal output distribution of $\mathcal{K}$ corresponding to $P_{i}$. However, this inequality is not always strict. Thus, we are interested in how much the channel $\mathcal{K}$ shrinks the KL divergence for the privacy of the respondent. To clarify this, we study the largest constant $\eta^{(c),\mathcal{K}}_{\mathbb{D}}$ satisfying the condition $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\eta^{(c),\mathcal{K}}_{\mathbb{D}}\mathbb{D}(P_{1}\|P_{2})$ for all pairs of distributions $P_{1},P_{2}\in\mathcal{P}(\mathcal{X})$. The quantity $\eta^{(c),\mathcal{K}}_{\mathbb{D}}$ is called the contraction coefficient of the channel $\mathcal{K}$ with respect to the divergence $\mathbb{D}$.
In particular, if we consider the set of all classical $(\varepsilon,\delta)$-LDP channels (see Definition 6) for some $\varepsilon\geq 0$ and $\delta\in[0,1)$, denoted by $\mathfrak{Q}_{\varepsilon,\delta}$, then it is interesting to study the worst contraction coefficient of any channel $\mathcal{K}$ with respect to the classical hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ (see Definition 6). This quantity is defined as follows,
$\eta^{(c),\varepsilon,\delta}_{E_{\gamma}}:=\sup_{\mathcal{K}\in\mathfrak{Q}_{\varepsilon,\delta}}\eta^{(c),\mathcal{K}}_{E_{\gamma}}\triangleq\sup_{\begin{subarray}{c}\mathcal{K}\in\mathfrak{Q}_{\varepsilon,\delta}\\ P_{1},P_{2}\in\mathcal{P}(\mathcal{X}):\\ E_{\gamma}(P_{1}\|P_{2})\neq 0\end{subarray}}\frac{E_{\gamma}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))}{E_{\gamma}(P_{1}\|P_{2})},$ (94)
where $\gamma\geq 1$ is some fixed constant.
[…]
$\eta^{\mathcal{N}}_{E_{\gamma}}:=\sup_{\begin{subarray}{c}\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A}):\\ E_{\gamma}(\rho\|\sigma)\neq 0\end{subarray}}\frac{E_{\gamma}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))}{E_{\gamma}(\rho,\sigma)}.$ (95)
Hirche et al. in [32] have studied the following bound on $\eta^{\mathcal{N}}_{E_{\gamma}}$ for any CP-TP some fixed $\gamma\geq 1$.
Proposition 3 ([32, Lemma II.4]).
For any $\gamma\geq 1$ and a CP-TP map $\mathcal{N}:\mathcal{D}(\mathcal{H}_{A})\to\mathcal{D}(\mathcal{H}_{B})$, the contraction coefficient $\eta^{\mathcal{N}}_{E_{\gamma}}$ with respect to the quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ (see Definition 13) satisfies the following,
$1-\gamma\left(1-\eta^{\mathcal{N}}_{E_{1}}\right)\leq\eta^{\mathcal{N}}_{E_{\gamma}}\leq\eta^{\mathcal{N}}_{E_{1}},$ (96)
where $\eta^{\mathcal{N}}_{E_{1}}$ is the contraction coefficient of quantum hockey stick divergence of order 1, i.e., trace distance with respect to $\mathcal{N}$.
From the context of Private CP-TP maps, the contraction coefficient of quantum hockey stick divergence under $(\varepsilon,\delta)$-LDP quantum mechanisms is defined as follows,
Definition 24 ([20]).
For any $\gamma\geq 1$, the contraction coefficient of $(\varepsilon,\delta)$-LDP quantum CP-TP maps(see Definition 23) with respect to $E_{\gamma}$ is defined as follows,
$\eta^{\varepsilon,\delta}_{E_{\gamma}}:=\sup_{\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta}}\eta^{\mathcal{N}}_{E_{\gamma}}\triangleq\sup_{\begin{subarray}{c}\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta}\\ \rho,\sigma\in\mathcal{D}(\mathcal{H}_{A}):\\ E_{\gamma}(\rho\|\sigma)\neq 0\end{subarray}}\frac{E_{\gamma}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))}{E_{\gamma}(\rho,\sigma)},$ (97)
where $\mathfrak{P}_{\varepsilon,\delta}$ is the set of all quantum $(\varepsilon,\delta)$-LDP CP-TP maps (see Definition 23) for some $\varepsilon\geq 0$ and $\delta\in[0,1)$.
Further, in [20], the authors studied the contraction coefficient for the trace distance $\eta^{\varepsilon,\delta}_{E_{1}}$ under quantum $(\varepsilon,\delta)$-LDP CP-TP maps where $\varepsilon\geq 0$ and $\delta\in[0,1]$. This is given in the proposition below.
Proposition 4 ([20, Theorem 5]).
For any $\varepsilon\geq 0$ and $\delta\in[0,1)$, the contraction coefficient $\eta^{\varepsilon,\delta}_{E_{1}}$ with respect to the trace distance satisfies
$\eta^{\varepsilon,\delta}_{E_{1}}=\frac{(e^{\varepsilon}-1+2\delta)}{e^{\varepsilon}+1}.$ (98)
In the following lemma, we now generalize the above result for the contraction coefficient of quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ for any $\gamma\geq 1$ under $(\varepsilon,\delta)$-LDP quantum mechanisms for any $\varepsilon\geq 0$ and $\delta\in[0,1]$.
Lemma 7.
For any $\gamma\geq 1,\varepsilon\geq 0$ and $\delta\in[0,1)$, the contraction coefficient $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ with respect to the quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ (see Definition 13) satisfies the following,
$\eta^{\varepsilon,\delta}_{E_{1}}+\frac{(2-\delta)(1-\gamma)}{e^{\varepsilon}+1}\leq\eta_{E_{\gamma}}^{\varepsilon,\delta}\leq\begin{cases}\eta^{\varepsilon,\delta}_{E_{1}}+\frac{(1-\delta)(1-\gamma)}{e^{\varepsilon}+1},&~\mbox{if}~\gamma\in[1,e^{\varepsilon}],\\ \delta,&~\mbox{if}~\gamma>e^{\varepsilon}.\end{cases}$ (99)
Proof.
(1) Lower-bound: Consider the following series of inequalities
$\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{a}{\geq}1-\gamma(1-\eta^{\varepsilon,\delta}_{E_{1}})$
$\overset{b}{=}1-2\gamma\frac{(1-\delta)}{e^{\varepsilon}+1}$
$=\frac{e^{\varepsilon}+1-2\gamma+2\gamma\delta}{e^{\varepsilon}+1}$
$\overset{c}{\geq}\frac{e^{\varepsilon}+\delta-\gamma(1-\delta)}{e^{\varepsilon}+1}+\frac{(1-\gamma)}{e^{\varepsilon}+1}$
$=\eta^{\varepsilon,\delta}_{E_{1}}+\frac{(1-\gamma)(2-\delta)}{e^{\varepsilon}+1},$
where $a$ follows from Proposition 3, $b$ follows from Proposition 4, $c$ follows because $\gamma\geq 1.$
(2) Upper-bound:
  1. (i)
    For the case when $\gamma\in[1,e^{\varepsilon}]$. Consider the following series of inequalities,
    $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{a}{=}\sup_{\begin{subarray}{c}\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta},\\ |\psi\rangle\perp|\phi\rangle\end{subarray}}E_{\gamma}(\mathcal{N}(|\psi\rangle\langle\psi|)\|\mathcal{N}(|\phi\rangle\langle\phi|))$
    $\overset{b}{\leq}E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$
    $\overset{c}{=}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},$ (100)
    $=\frac{e^{\varepsilon}-1+2\delta}{e^{\varepsilon}+1}+\frac{(1-\delta)(1-\gamma)}{e^{\varepsilon}+1}$
    $=\eta^{\varepsilon,\delta}_{E_{1}}+\frac{(1-\delta)(1-\gamma)}{e^{\varepsilon}+1},$ (101)
    where in $a$ $|\psi\rangle,|\phi\rangle\in\mathcal{H}_{A}$ are two orthogonal quantum states and the inequality follows from [32, Theorem II.2], $b$ follows because $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜51 and 53 and $c$ follows from Claim 1 below.
    Claim 1.
    For any $\gamma\geq 1,\varepsilon\geq 0$ and $\delta\in[0,1)$, the quantum hockey stick divergence $E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$ satisfies the following,
    $E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=\begin{cases}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},&~\mbox{if}~\gamma\in[1,e^{\varepsilon}],\\ \delta,&~\mbox{if}~\gamma>e^{\varepsilon}.\end{cases}$ (102)
    where $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜51 and 53.
    Proof.
    See Appendix -G for the proof.
  2. (ii)
    For the case when $\gamma>e^{\varepsilon}$. The RHS of (100), will be replaced by $\delta$ using Claim 1. This completes the proof of Lemma 7.
Remark 6.
The difference between the upper and lower bound in Lemma 7 is $\frac{\gamma-1}{e^{\varepsilon}+1}\leq\tanh(\frac{\varepsilon}{2}).$ For small values of $\varepsilon,\tanh(\frac{\varepsilon}{2})\leq\mathcal{O}(\varepsilon).$ Thus, making these bounds almost tight in the small $\varepsilon$ regime. Also, it trivially follows that for $\gamma=1$ both the upper and lower bound coincide and are equal to $\eta^{\varepsilon,\delta}_{E_{1}}$.
Further, the upper-bound obtained in Lemma 7 is tighter in comparison to the upper-bound obtained in Proposition 3 by a subtractive factor of $\frac{(1-\delta)(\gamma-1)}{e^{\varepsilon}+1}.$
In a classical scenario, it is trivial to observe that the privatized contraction coefficient $\eta^{(c),\varepsilon,\delta}_{E_{\gamma}}$ with respect to classical hockey stick divergence has the same upper and lower bounds as $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ mentioned in Lemma 7 i.e. we have the following,
Corollary 9.
For any $\varepsilon\geq 0$ and $\delta\in[0,1)$, the contraction coefficient $\eta_{E_{\gamma}}^{(c),\varepsilon,\delta}$ with respect to the classical hockey stick divergence (see Definition 6) satisfies the following,
$\eta_{E_{\gamma}}^{(c),\varepsilon,\delta}\leq\begin{cases}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},&~\mbox{if}~\gamma\in[1,e^{\varepsilon}],\\ \delta,&~\mbox{if}~\gamma>e^{\varepsilon}.\end{cases}$ (103)
In the quantum setting, Nuradha et al in [20, Proposition 3] obtain an upper-bound on quantum relative entropy under $\varepsilon$-LDP CP-TP maps via an integral representation of quantum relative entropy (Fact 6) in terms of quantum hockey stick divergence. Moreover, the upper-bound obtained in [20, Proposition 3] is tighter than the upper-bound obtained in $(i)$ of Corollary 4.
However, for $\delta>0$, it is not clear if one can obtain a similar upper-bound on quantum relative entropy under $(\varepsilon,\delta)$-LDP CP-TP maps via an integral representation of quantum relative entropy in terms of quantum hockey stick divergence. This is because, unlike the $\varepsilon$-LDP channels, for $(\varepsilon,\delta)$-LDP channels (with $\delta>0$), the contraction coefficient of quantum hockey stick divergence never becomes 0 for any $\gamma\geq 1$ (see (99)). This makes the integration mentioned in Fact 6 un-integrable.
To resolve this issue, in the subsection below, we come up with a technique which we call truncation. This technique allows us to distill pure DP from non-pure DP.
VIII-C upper-bound on the relative entropy of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
In this subsection, we obtain a tight upper-bound on the relative entropy of the output distributions of any $(\varepsilon,\delta)$-LDP classical channel (Markov kernel) via the integral representation of relative entropy in terms of hockey-stick divergence. To obtain this, we define a two-sided truncation of a $(\varepsilon,\delta)$-DP pair of distributions $(P,Q)$ in the definition below.
Definition 25 (Truncated Pair of Distributions).
[…]
$\tilde{P}(x):=\frac{P^{\prime}(x)}{\sum_{x^{\prime}\in\mathcal{X}}P^{\prime}(x)},$ (104)
$\tilde{Q}(x):=\frac{Q^{\prime}(x)}{\sum_{x^{\prime}\in\mathcal{X}}Q^{\prime}(x)},$ (105)
where, $P^{\prime}(x):=\min\left\{P(x),e^{\varepsilon}Q(x)\right\}$ and $Q^{\prime}(x):=\min\left\{Q(x),e^{\varepsilon}P(x)\right\}$.
[…]
Lemma 8.
[…]
  1. i)
    $\|\tilde{P}-P\|_{1}\leq 2\delta$ and $\|\tilde{Q}-Q\|_{1}\leq 2\delta.$
  2. ii)
[…]
  3. iii)
    if $\operatorname{supp}(P)\subseteq\operatorname{supp}(Q)$, then $\left|D(P\|Q)-D(\tilde{P}\|\tilde{Q})\right|\leq 2\delta\left(\varepsilon+\log\frac{1}{1-\delta}+\frac{2}{m}\right)$.
where $m=\min_{x\in\operatorname{supp}(\tilde{P})}\left\{\min\{\tilde{P}(x),P(x)\},\min\{\tilde{Q}(x),Q(x)\}\right\}.$
Proof.
See Appendix -H for the proof.
Using Lemma 8 and the integral representation of relative entropy in terms of hockey-stick divergence (Fact 6), we now obtain an upper-bound on the relative entropy of the output distributions of any $(\varepsilon,\delta)$-LDP classical channel (Markov kernel) in Theorem 6 below.
Theorem 6.
Let $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ be a $(\varepsilon,\delta)$-LDP classical channel (Markov kernel). Further, for any pair of probability distributions $P_{X}$ and $Q_{X}$ over $\mathcal{X}$, let $\mathcal{K}(P_{X})$ and $\mathcal{K}(Q_{X})$ be their respective output distribution with respect to $\mathcal{K}$ such that $\operatorname{supp}(\mathcal{K}(P_{X}))\subseteq\operatorname{supp}(\mathcal{K}(Q_{X}))$. Then,
$D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ $\leq\frac{1}{2}{\left\lVert P_{X}-Q_{X}\right\rVert}_{1}\left(\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)+\delta\left(\frac{2\varepsilon}{e^{\varepsilon}+1}+\frac{e^{\varepsilon}-1}{e^{\varepsilon}}+\log\frac{1}{1-\delta}+\frac{\delta}{e^{\varepsilon}}\right)\right)$
$\quad+\delta\left(\frac{e^{\varepsilon}}{1-\delta}+2\log\frac{e^{\varepsilon}}{1-\delta}-\frac{1-\delta}{e^{\varepsilon}}+2\left(\varepsilon+\log\frac{1}{(1-\delta)}+\frac{2}{m}\right)\right),$
where $m=\min_{y\in\operatorname{supp}(\mathcal{K}(P_{X}))}\left\{\min\{\tilde{P}_{Y}(y),\mathcal{K}(P_{X})(y)\},\min\{\tilde{Q}_{Y}(y),\mathcal{K}(Q_{X})(y)\}\right\}$ and $(\tilde{P}_{Y},\tilde{Q}_{Y})$ is the truncated pair with respect to $(\mathcal{K}(P_{X}),\mathcal{K}(Q_{X}))$.
Proof.
See Appendix -I for the proof.
Alternatively, we can obtain a different upper-bound on $D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ by directly applying the continuity bound for relative entropy from (ii) of Lemma 8 and substituting $\varepsilon\leftarrow\left(\varepsilon+\log\frac{1}{1-\delta}\right)$ in Corollary 4.
Theorem 7.
Let $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ be a $(\varepsilon,\delta)$-LDP classical channel (Markov kernel). Further, for any pair of probability distributions $P_{X}$ and $Q_{X}$ over $\mathcal{X}$, let $\mathcal{K}(P_{X})$ and $\mathcal{K}(Q_{X})$ be their respective output distribution with respect to $\mathcal{K}$ such that $\operatorname{supp}(\mathcal{K}(P_{X}))\subseteq\operatorname{supp}(\mathcal{K}(Q_{X}))$. Then,
$D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ $\leq\varepsilon^{\prime}\tanh\left(\frac{\varepsilon^{\prime}}{2}\right)+2\delta\left(\varepsilon^{\prime}+\frac{2}{m}\right),$
where $\varepsilon^{\prime}=\left(\varepsilon+\log\frac{1}{1-\delta}\right)$ and $m=\min_{y\in\operatorname{supp}(\mathcal{K}(P_{X}))}\left\{\min\{\tilde{P}_{Y}(y),\mathcal{K}(P_{X})(y)\},\min\{\tilde{Q}_{Y}(y),\mathcal{K}(Q_{X})(y)\}\right\}$ and $(\tilde{P}_{Y},\tilde{Q}_{Y})$ is the truncated pair with respect to $(\mathcal{K}(P_{X}),\mathcal{K}(Q_{X}))$.
Remark 7.
It is important to note that the bound obtained in Theorem 6 is tighter than the bound in Theorem 7. This is because the bound in Theorem 7 is obtained by first approximating the $(\varepsilon,\delta)$-DP pair with an $(\varepsilon^{\prime},0)$-DP pair, where $\varepsilon^{\prime}=\varepsilon+\log\frac{1}{1-\delta}$, and then applying the known bounds for pure DP. This approximation introduces looseness, particularly in the leading term, which becomes $\mathcal{O}((\varepsilon+\delta)\tanh(\varepsilon+\delta))$. In contrast, Theorem 6 uses a more direct approach via the integral representation of the KL divergence, resulting in a leading term of $\mathcal{O}(\varepsilon\tanh(\varepsilon)){\left\lVert P_{X}-Q_{X}\right\rVert}_{1}$. For small $\varepsilon$, the bound in Theorem 6 is therefore significantly tighter.
IX Conclusion
In this work, inspired by [7], we develop a framework for studying quantum differential privacy from the perspective of hypothesis testing and Blackwell’s ordering [9]. We provided a characterization of quantum $(\varepsilon,\delta)$-differential privacy in terms of quantum hypothesis testing divergences, and identified the most informative (in the Blackwell sense) pair of quantum states. We use this framework for studying the stability of differentially private quantum learning algorithms. Our stability result also generalizes (in the sense that $\delta>0$) the existing bound in the classical settings.
We also study the problem of quantum privatized parameter estimation, where the trade-off between privacy and statistical utility is characterized via the quantum Fisher information. We derived explicit expressions for the maximal Fisher information achievable under quantum $(\varepsilon,\delta)$-DP constraints, thereby quantifying the fundamental limits of parameter estimation in the presence of quantum privacy mechanisms.
Further, we derive near-optimal bounds on the contraction coefficient of $(\varepsilon,\delta)$-DP CP-TP maps with respect to the hockey stick divergence. This allows us to prove bounds on the relative entropy between the output pair induced by any $(\varepsilon,\delta)$-DP classical channels.
Acknowledgments
The work of NAW was supported in part by MTR/2022/000814. The work of MH was supported in part by the National Natural Science Foundation of China under Grant 62171212 and in part by the General Research and Development Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund under Grant GRDP2025-022.

### v3

@@
-Quantum Blackwell’s Ordering and Differential Privacy
+Quantum Information Ordering and Differential Privacy
 Abstract
-We develop a framework for quantum differential privacy (QDP) based on quantum hypothesis testing and Blackwell’s ordering.
-This approach characterizes $(\varepsilon,\delta)$-QDP via hypothesis testing divergences and identifies the most informative quantum state pairs under privacy constraints.
-We apply this to analyze the stability of quantum learning algorithms, generalizing classical results to the case $\delta>0$.
-Additionally, we study privatized quantum parameter estimation, deriving tight bounds on the quantum Fisher information under QDP.
+We study quantum differential privacy (QDP) by defining a notion of the order of informativeness between pairs of quantum states.
+In particular, we show that if the hypothesis testing divergence of one pair dominates over that of the other pair, then this dominance holds for every $f$-divergence.
+This approach completely characterizes $(\varepsilon,\delta)$-QDP mechanisms by identifying the most informative $(\varepsilon,\delta)$-DP quantum state pairs.
+We apply this to study precise limits for privatized hypothesis testing and privatized quantum parameter estimation, including tight upper-bounds on the quantum Fisher information under QDP.
 Finally, we establish near-optimal contraction bounds for differentially private quantum channels with respect to the hockey-stick divergence.
 I Introduction
-A fundamental challenge in modern machine learning is the trade-off between privacy and information extraction.
-In this work, we explicitly treat both sides: privacy (ensuring that algorithmic outputs do not reveal significant information about the input data of the respondents) and the investigator’s goal to extract as much useful information as possible from data for accurate learning and estimation.
-With the rapid advancement of machine learning, a key concern is about ensuring the privacy of learning algorithms, meaning that their outputs should not reveal significant information about the input data.
-Differential privacy (DP) provides a rigorous mathematical framework to balance these opposing requirements.
-Accordingly, we structure our contributions in three steps: first step (privacy), second step (information extraction under privacy constraints), and third step, the quantum channel setup, where the situation is more complicated, and we mark the transition to each step explicitly in the text.
+The trade-off between extracting useful insights by an investigator and preserving a respondent’s privacy lies at the heart of modern quantum machine learning.
+While classical Differential Privacy (DP) [13, 14] provides a rigorous framework to limit adversarial inference via hypothesis testing, extending these guarantees to the quantum regime presents unique challenges.
+Recent advances in Quantum Differential Privacy (QDP) [36] have made non-trivial attempts [12, 18, 1, 25, 26] to understand differentially private mechanisms in the quantum setting.
+In this manuscript, we advance the understanding of quantum differential privacy by grounding QDP in the geometry of asymmetric hypothesis testing.
+We establish a universal information ordering for differentially private mechanisms, proving that the entire privacy regime can be analyzed by a single “most-informative” pair of quantum states.
+Leveraging this, we derive exact, mechanism-agnostic limits for different statistical tasks: the optimal error exponent in hypothesis testing, the maximum achievable Fisher information for parameter estimation, and the fundamental contraction of distinguishability under private quantum channels.
+Framework for Differential Privacy.— In the quantum setting, differential privacy constrains the output which the quantum algorithms produce.
+Formally, consider a scenario where a respondent prepares a quantum system in state $\tau\in\{\rho,\sigma\}$.
+The pair $(\rho,\sigma)$ satisfies $(\varepsilon,\delta)$-DP if, for any general quantum measurement (POVM) performed by an investigator, the probability distributions of the measurement outcomes are statistically close.
+Thus, as defined in [36] for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, $(\rho,\sigma)$ is $(\varepsilon,\delta)$-DP if they satisfy the following.
+Definition 1.
+For every POVM measurement $0\preceq\Lambda\preceq\mathbb{I}$, the following holds:
+$\begin{split}\mathrm{Tr}[\Lambda\rho]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,\\ \mathrm{Tr}[\Lambda\sigma]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\rho]+\delta.\end{split}$ (1)
+Further, we denote $\mathfrak{D}_{(\varepsilon,\delta)}$ to be the collection of all pairs of quantum states that satisfy (1).
+Moreover, for $\delta=0$, we denote $(\rho,\sigma)$ to be pure $\varepsilon$-DP or just $\varepsilon$-DP i.e.
+$(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$.
+For the case when $\delta=0$, any value of $\varepsilon$ satisfying Eq.
+(1) must be at least $\max\{D_{\max}(\rho\|\sigma),\,D_{\max}(\sigma\|\rho)\},$ where $D_{\max}$ denotes the max-relative entropy.
+As a consequence, the required $\varepsilon$ is typically very large, which severely weakens privacy as discussed in the section below.
+To mitigate this issue, one introduces a positive additive constant $\delta$, allowing the permissible value of $\varepsilon$ to be significantly reduced.
+Further in Eq.
+(1), if we consider only binary POVMs then it motivates us to analyze QDP from the point of view of binary quantum hypothesis testing.
+The Operational Scenario: The Respondent and the Investigator.— This manuscript is based on the strategic interaction between two distinct parties: the Respondent (the data owner) and the Investigator (the adversary).
+The Respondent serves as the gatekeeper of a sensitive database containing potential datasets $S$ and $T$, which differ solely by the presence of a single individual $A$, such that $S\setminus(S\cap T)=\{A\}$.
+To make data accessible while maintaining confidentiality, the Respondent instantiates a statistical oracle (the mechanism).
+This oracle operates as a conditional distribution $\mathbb{P}^{(D)}(O\mid I)$—or in the quantum setting, a quantum channel—generating an output $O$ for a specific query $I$ based on the underlying dataset $D\in\{S,T\}$.
+The Respondent’s primary objective is to configure this oracle to satisfy strict differential privacy constraints, effectively masking the specific identity of the dataset.
+The Investigator operates as the Adversary in this setting.
+They interact with the Respondent’s oracle by issuing queries and processing the resulting outputs.
+Their goal is to execute a binary hypothesis test to breach privacy: specifically, to distinguish whether the underlying dataset is $S$ or $T$, thereby revealing the presence of individual $A$.
+The central thesis of this work is that the “informativeness” of the Respondent’s oracle dictates the success of the Investigator: the more useful the oracle is for legitimate statistical utility, the more vulnerable it is to this adversarial attack.
+We structure our contributions in three steps corresponding to this dynamic: the privacy limit (comparing the Respondent’s oracles), the adversarial extraction limit (the Investigator’s power), and the channel analysis (the properties of the Respondent’s defense mechanism).
 First step: Privacy.
-This step develops the privacy side of the trade-off from the respondent’s perspective by studying the stability [1, 2] of learning algorithms.
-From the respondent’s viewpoint, privacy means that the inclusion or exclusion of their individual data should not materially affect the mechanism’s output, so that they can contribute data without fear of singled-out inference.
-An algorithm is considered stable if its output does not change drastically when a single respondent’s data is changed; this point-wise insensitivity is precisely the respondent-centric guarantee we seek.
-Differential privacy $($$\varepsilon,\delta$)-DP) [3, 4] formalizes this requirement as a strong, mathematically precise form of stability: by ensuring that the output states (distributions) corresponding to any two neighboring datasets are close, an $(\epsilon,\delta)$-DP algorithm guarantees that no respondent can be reliably distinguished from the output alone [5].
-In the classical case, the stability of learning algorithms was studied in [6] only for the pure $\varepsilon$-DP ($\delta=0$) case.
-The main aim of this manuscript is to generalize the result of [6] to the quantum learning scenario for $(\varepsilon,\delta)$-DP.
-To achieve this, we build a framework for studying quantum differentially private mechanisms $($$\varepsilon,\delta$)-DP).
-This framework is inspired by [7] and is related to the problem of hypothesis testing (see for example, [8]).
-The main aim of differential privacy is to ensure that the outputs of a private mechanism corresponding to two neighboring inputs are hard to distinguish.
-To understand this, consider the following scenario.
-Suppose inside a database of datasets, we have two datasets $S$ and $T$ such that there is exactly one individual, say $A$ is present in set $S$ and not in $T$, i.e., $S\setminus(S\cap T)=\{A\}$ and suppose there is an oracle (randomized) which is attached to the database.
-From the perspective of an attacker, the only way it can break the privacy of this oracle is by performing a hypothesis test on the output of the oracle to detect whether the individual $A$ is a member of $S$ or not.
-From this discussion, it implies that the more informative the oracle is about the datasets, the less private it is.
-A notion for comparing the “informativeness” between two mechanisms (oracles) was first studied by Blackwell [9].
-In [10, Theorem 12.2.2], he showed that if a mechanism (oracle) is more informative as compared to the other mechanism, then there exists a Markov kernel (channel) which maps the more informative one to the less informative one.
-A quantum version of [10, Theorem 12.3.1] was proved by Buscemi in [11], wherein he defined a notion of ordering between quantum channels based on a certain definition of informativeness.
-Using this ordering, Buscemi showed that there exists a statistical morphism from the more informative quantum channel to the lesser one.
-[…]
-Using Lemma 1 (a quantum version of [10, Theorem 12.4.2]), we show that there exists a worst case $(\varepsilon,\delta)$-DP pair of quantum states which can be mapped to any other $(\varepsilon,\delta)$-DP pair of states by applying a CP-TP map.
-Thus, this version of the quantum Blackwell theorem (Lemma 1), along with this worst case $(\varepsilon,\delta)$-DP pair, provides us with a powerful machinery to analyze the stability of quantum differentially private $($$\varepsilon,\delta$)-DP) learning algorithms.
+The primary objective of differential privacy is to ensure that the outputs of the oracle corresponding to two neighboring inputs are difficult for the Investigator to distinguish.
+This implies a direct inverse relationship: the more informative the oracle is about the datasets, the less private it is.
+To formalize this, we require a rigorous method to compare the “informativeness” of different oracles.
+We consider a pair of oracles modeled as $\mathbb{P}^{(D)}(O\mid I)$ and $\mathbb{Q}^{(D)}(O\mid I)$.
+We say that $\mathbb{P}^{(D)}(O\mid I)$ is more informative than $\mathbb{Q}^{(D)}(O\mid I)$ (represented as the order $\mathbb{P}^{(D)}(O\mid I)\succeq\mathbb{Q}^{(D)}(O\mid I)$ or simply $\mathbb{P}^{(D)}(O\mid I)$ dominates $\mathbb{Q}^{(D)}(O\mid I)$) if the pair $\{\mathbb{P}^{(S)}(O\mid I),\mathbb{P}^{(T)}(O\mid I)\}$ is statistically easier to distinguish as compared to the pair $\{\mathbb{Q}^{(S)}(O\mid I),\mathbb{Q}^{(T)}(O\mid I)\}$.
+In $1951$, Blackwell [4, 3] formalized this notion of comparing “informativeness” for classical statistical experiments.
+He showed that $\mathbb{P}^{(D)}(O\mid I)\succeq\mathbb{Q}^{(D)}(O\mid I)$ if and only if there exists a Markov kernel (stochastic map) $\mathcal{M}$ such that $\mathbb{Q}^{(S)}(O\mid I)=\mathcal{M}(\mathbb{P}^{(S)}(O\mid I))$ and $\mathbb{Q}^{(T)}(O\mid I)=\mathcal{M}(\mathbb{P}^{(T)}(O\mid I))$.
+Further, using the data-processing inequality (monotonicity property) we have $D_{f}(\mathbb{P}^{(S)}(O\mid I)\|\mathbb{P}^{(T)}(O\mid I))\geq D_{f}(\mathbb{Q}^{(S)}(O\mid I)\|\mathbb{Q}^{(T)}(O\mid I))$, where $D_{f}$ is the $f$-divergence [29].
+For the case when these oracles are differentially private (with the same privacy parameters), it is obvious that the above monotonicity would help us obtain security bounds of the less informative in terms of the one that is more informative.
+Likewise, in the quantum domain, we consider two quantum oracles whose outputs are $\{\rho^{S},\rho^{T}\}$ and $\{\sigma^{S},\sigma^{T}\}$.
+We say that $\{\rho^{S},\rho^{T}\}$ is more informative than $\{\sigma^{S},\sigma^{T}\}$ if the former is statistically easier for the Investigator to distinguish.
+However, a significant theoretical hurdle arises here: unlike the classical case, it follows from [15] and [23] that the statistical ordering $\{\rho^{S},\rho^{T}\}\succeq\{\sigma^{S},\sigma^{T}\}$ does not imply the existence of a completely positive trace-preserving (CP-TP) map—or even a positive trace-preserving map—that transforms the former into the latter.
+As a consequence, we cannot rely on the standard data-processing inequalities of quantum $f$-divergences [19] to transfer bounds.
+We overcome this in Theorem 2 by proving that an explicit channel (a CP-TP map) construction is unnecessary.
+We show that if the quantum hockey stick divergence [31] between the oracle outputs $\{\rho^{S},\rho^{T}\}$ dominates $\{\sigma^{S},\sigma^{T}\}$ for every parameter, then the inequality $D_{f}(\rho^{S}\|\rho^{T})\geq D_{f}(\sigma^{S}\|\sigma^{T})$ holds universally.
+This result allows us to define a “worst-case” (most informative) pair of states that characterizes the upper bound on the Investigator’s adversarial capability.
 Second step: Information extraction under privacy constraints.
-We now analyze, under the $(\varepsilon,\delta)$-DP constraints, how much useful information an investigator can still extract.
-That is, we investigate the statistical inference of quantum privatized parameter using our quantum version of Blackwell’s theorem (Lemma 1).
-This is a common task where we want to guess a hidden number or parameter from data that has been made private.
-How well the investigator can guess this number is limited by how much information the privacy method keeps.
-This limit is measured by a value called the Fisher information.
-In the quantum case, we use a related measure called the Symmetric Logarithmic Derivative (SLD) Fisher information to judge performance [12].
-By using the idea of our “weakest” (most informative) private method, we can figure out the exact maximum SLD Fisher information the investigator can get while still respecting the $(\varepsilon,\delta)$-DP rule.
-This work improves on past studies that only looked at the simpler case where $\delta=0$ [13], and sets a hard limit on how accurate quantum estimations can be under the privacy condition.
-Third step: We have two kinds of extensions.
-In the privacy of the quantum learning framework, we can relax the $(\varepsilon,\delta)$-DP condition by trusting the party, the data processor, who performs the learning algorithm.
-That is, motivated by the quantum learning framework of [14, 15], we propose a framework for a quantum learning algorithm which is “$1$-neighbor" $(\varepsilon,\delta)$-DP.
-For this proposed learning framework, we study its stability by obtaining an upper-bound on the Holevo information (between the training data and the quantum output of the algorithm).
-Further, we show that for $\delta=0$, our upper-bound recovers [6, Proposition 2] in the classical case.
-Another interesting problem is to analyze quantum channels which are $(\varepsilon,\delta)$-DP.
-Its classical version has been extensively studied in the literature via the contraction of output divergence-measured with respect to the output distributions induced by $(\varepsilon,\delta)$-DP channels-in comparison to the input divergence, which is defined over the input distributions of these channels; see, for example, [16], [17], and the references therein.
-This contraction is studied using an integral representation which involves the hockey stick divergence [18].
-We leverage our framework to first establish an almost tight upper-bound on the contraction coefficient of any $(\varepsilon,\delta)$-DP CPTP map with respect to the hockey stick divergence.
-In the process of studying the contraction for divergences, we observe that in the classical case, there exists a pure $(\varepsilon+\log(1/1-\delta))$-DP pair of distributions which can be obtained by truncating the original $(\varepsilon,\delta)$-DP pair.
-Further, we show that this truncated pair is not far ($O(\delta)$, in $L_{1}$ distance) from the original pair.
-We use this truncated pair to obtain a meaningful upper-bound on the relative entropy (when it is well-defined) of the original $(\varepsilon,\delta)$-DP pair of distributions using the continuity of the relative entropy.
-This upper-bound recovers the known result for $\delta=0,$ [19, 7] as a special case.
-Moreover, these contraction-based analysis serve to connect the channel-level action with both privacy guarantees and investigator-agnostic limits on information extraction, by bounding how much input distinctions can be suppressed at the output, we obtain direct, worst-case control of output distinguishability (useful for hypothesis-testing-based privacy statements) and of output information quantities (useful for bounding mutual information, Fisher-information-based estimation limits).
+We now analyze, under the $(\varepsilon,\delta)$-DP constraints, the fundamental limits of the Investigator’s ability to extract information from Respondent’s data.
+Here, we treat the Investigator as an adversary whose goal is to infer the hidden parameter or state.
+We investigate this statistical inference using our quantum $f$-divergence dominance result (Theorem 2).
+The Investigator’s success in guessing the hidden parameter is strictly limited by the privacy constraints imposed by the mechanism.
+We examine this limit through two adversarial viewpoints: binary hypothesis testing and parameter estimation.
+In the former scenario, we begin with the one-shot setting, where the Investigator attempts to distinguish between two nearby sources $\rho_{\theta_{0}}$ and $\rho_{\theta_{1}}$ (parametrized by two private parameters $\theta_{1}$ and $\theta_{2}$) after their encoding has been privatized through an $(\varepsilon,\delta)$-DP mechanism.
+This scenario exposes the operational limit of privacy breaches.
+We show that, among all $(\varepsilon,\delta)$-DP mechanisms, the Investigator’s optimal performance—quantified by the hypothesis testing divergence—is maximized by the “weakest” (most informative) privatized mechanism permitted by the definition.
+This optimality carries over to KL, hockey stick Rényi (see Definition 12), and measured Rényi divergences.
+We then extend the analysis to the asymptotic regime and establish that this same privatized mechanism determines the optimal error exponent.
+This yields a unified characterization: the privacy requirement compresses the geometry of the state space, imposing a fundamental ceiling on the Investigator’s probability of successfully distinguishing hypotheses.
+In the regime of parameter estimation, the limit of the Investigator’s ability to pinpoint the private parameter is measured by the Fisher information.
+In the quantum case, we use the Symmetric Logarithmic Derivative (SLD) Fisher information [17].
+By leveraging the idea of our “weakest” (most informative) private method, we determine the exact maximum SLD Fisher information the Investigator can obtain while the mechanism respects the $(\varepsilon,\delta)$-DP rule.
+This result improves on past studies that only looked at the simpler case where $\delta=0$ [34], and sets a hard physical limit on how accurately an adversarial Investigator can estimate sensitive parameters under the privacy condition.
+Third step: The Quantum Channel.
+Finally, we analyze the mechanism itself by studying quantum channels which are $(\varepsilon,\delta)$-DP.
+From the adversarial perspective, the channel acts as a defense mechanism that suppresses the distinguishability of inputs.
+We quantify this suppression via the contraction of divergence—comparing the divergence of the output distributions induced by $(\varepsilon,\delta)$-DP channels to the divergence of the input states; see, for example, [2], [35], and the references therein.
+This contraction is studied using an integral representation involving the hockey stick divergence [29].
+We leverage our framework to first establish an almost tight upper bound on the contraction coefficient of any $(\varepsilon,\delta)$-DP CP-TP map with respect to the hockey stick divergence.
+A crucial technical contribution in this step addresses the difficulty of the $\delta>0$ regime.
+The additive $\delta$ term typically complicates the analysis of entropic quantities.
+We observe that, similar to the classical case, there exists a pure $(\varepsilon+\log(1/1-\delta))$-DP pair of distributions that can be obtained by truncating the original $(\varepsilon,\delta)$-DP pair.
+We formally extend this to the quantum domain, showing that this truncated pair is close ($O(\delta)$ in $L_{1}$ distance) to the original pair.
+We use this truncated pair to obtain a meaningful upper bound on the relative entropy of the original $(\varepsilon,\delta)$-DP pair using the continuity of relative entropy.
+This upper bound recovers the known result for $\delta=0$ [22, 11] as a special case.
+Moreover, these contraction-based analyses serve to connect the channel-level action with both privacy guarantees and the limits on adversarial extraction.
+By bounding how much input distinctions are suppressed at the output, we obtain direct, worst-case control of output distinguishability (useful for bounding the success of hypothesis-testing attacks) and of output information quantities (limiting the precision of Fisher-information-based estimation).
 TABLE I: Relationship between results obtained in this work and related studies
@@
 $\varepsilon$-DP $(\varepsilon,\delta)$-DP $\varepsilon$-DP $(\varepsilon,\delta)$-DP
-Blackwell’s Informativeness
-Theorem
-[10, Theorem 12.4.1 and 12.4.2]
-(Proposition 1 in this manuscript.)
-Lemma 1 in this manuscript.
-Existence of Weakest
-DP Mechanism
-[19, Theorem 5]
-[19, Theorem 18]
-[13, Theorem 2]
-Lemma 4
-of this manuscript.
-Data Processing based
-upper-bounds using Weakest
-DP Mechanism
-[7, Lemma D.8]
-Corollary 3
-of this manuscript.
-Corollary 2
-of this manuscript.
-Privatized Parameter
-Estimation
-[19, Theorem 18]
-[19, Theorem 18]
-[13, Section IIA]
-Theorem 3
-of this manuscript.
-Stability Upper-bounds
-of Private Learning
-Algorithms
-[6, Proposition 2]
-Proposition 2
-of this manuscript.
-Equations (i)),
-(ii)) and (iii)) of
-this manuscript.
-Theorem 5
-of this manuscript.
-Theorem 5
-of this manuscript.
-Privatized Contraction
-Coefficient of Trace Distance
-[19, Corollary 11]
-[20, Theorem 22]
-[20, Theorem 55]
-Contraction Coefficient
-of Hockey-stick Divergence
-[17, Theorem 1]
-Corollary 9
-of this manuscript.
-[20, Theorem 1]
-Lemma 7
-of this manuscript.
-Contraction Coefficient based
-Upper-bounds on Relative entropy
-of LDP channels
-Theorem 6
-of this manuscript.
-[20, Proposition 3]
+Existence of Weakest DP Mechanism [22, Theorem 5] [22, Theorem 18] [34, Theorem 2] Lemma 3 of this manuscript.
+Data Processing based upper-bounds using Weakest DP Mechanism [11, Lemma D.8] Corollary 2 of this manuscript. Theorem 2 of this manuscript.
+Privatized Hypothesis Testing [22, Theorem 18] [22, Theorem 18] [34, Section IIB] Theorem 3 and Corollary 4 of this manuscript.
+Privatized Parameter Estimation [22, Theorem 18] [22, Theorem 18] [34, Section IIA] Theorem 4 of this manuscript.
+Privatized Contraction Coefficient of Trace Distance [22, Corollary 11] [26, Theorem $2$] [26, Theorem $5$]
+Contraction Coefficient of Hockey-stick Divergence [35, Theorem 1] Corollary 5 of this manuscript. [26, Theorem 1] Lemma 5 of this manuscript.
+Contraction Coefficient based Upper-bounds on Relative entropy of LDP channels Theorem 5 of this manuscript. [26, Proposition 3]
 I-A Organization of this Manuscript and our Contributions
@@
   • In Section II, we list the essential notations, definitions, and facts which will be used throughout this manuscript.
-  • In Section III, we develop a quantum generalization of Blackwell’s theorem, which establishes an ordering of informativeness for pairs of quantum states.
-In Lemma 1, we show that one pair is less informative than the other if and only if there exists a completely positive map connecting them.
-This forms the foundation for comparing privacy mechanisms in the quantum setting.
-  • In Section V, we prove Lemma 3, which shows that a pair of quantum states is $(\varepsilon,\delta)$-differentially private if and only if $D^{\alpha}_{H}{(\rho\|\sigma)}\leq\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1],$ where $D^{\alpha}_{H}{(\rho\|\sigma)}$ is the hypothesis testing divergence between $\rho$ and $\sigma.$ Using this in Lemma 4, we identify an explicit “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states that dominates all other.
-  • In Section VI, we apply our framework to quantum parameter estimation under privacy constraints.
-In particular, in Theorem 3, we derive the exact maximum SLD Fisher information achievable by any $(\varepsilon,\delta)$-DP mechanism.
-  • In Section VII, we study the stability of $(\varepsilon,\delta)$-DP learning algorithms.
-We first discuss a framework for differentially private learning algorithms.
-Then, using this framework, in Theorem 5, we study their stability by deriving upper-bound on the Holevo information.
-  • In Section VIII, we analyze the contraction of divergences under differentially private channels.
+  • In Section III, we provide a geometric characterization of the privacy region of $(\varepsilon,\delta)$-QDP and provide a construction of the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states.
+  • In Section IV, we first define a notion of order of informativeness between two pairs of quantum states via quantum hypothesis testing divergence.
+Using this ordering, in Theorem 1, we give an $f$-divergence dominance of a more informative pair of quantum states over a lesser informative pair of quantum states.
+Further, in Lemma 2, we show that a pair of quantum states is $(\varepsilon,\delta)$-differentially private if and only if $D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\right\}$ for any $\alpha\in[0,1],$ where $D^{\alpha}_{H}{(\rho\|\sigma)}$ is the hypothesis testing divergence between $\rho$ and $\sigma.$ Using this in Lemma 3, we identify an explicit “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states that dominates all others.
+Finally, in Theorem 2, we show that the $f$-divergence of any pair of $(\varepsilon,\delta)$-DP quantum states can be upper-bounded by that of the “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states.
+  • In Section V, we apply our framework to the regime of the quantum hypothesis testing and the quantum parameter estimation under privacy constraints.
+In particular, in Theorem 3, we show that among all $(\varepsilon,\delta)$-DP mechanisms, the hypothesis testing divergence is maximized by the weakest (most informative) privatized mechanism.
+Further, in Theorem 4, we derive the exact maximum SLD Fisher information achievable by any $(\varepsilon,\delta)$-DP mechanism.
+  • In Section VI, we analyze the contraction of divergences under differentially private channels.
 In Lemma 5, we obtain nearly tight upper and lower bounds on the contraction coefficient of $(\varepsilon,\delta)$-LDP channels with respect to the quantum hockey-stick divergence.
@@
 II Notations, Definitions and Facts
-We use $\mathcal{H}$ to denote a finite-dimensional Hilbert space and we denote its dimension with $\left|\mathcal{H}\right|$, $\mathcal{D}(\mathcal{H})$ to represent the set of all state density matrices acting on $\mathcal{H}$.
-For any quantum state $\rho\in\mathcal{D}(\mathcal{H})$, we define $\mbox{supp}(\rho):=\mbox{Span}\{|i\rangle:\lambda_{i}>0\},$ where $\left\{\lambda_{i}\right\}$ represent non-zero eigenvalues of $\rho$.
+We use $\mathcal{H}$ to denote a finite-dimensional Hilbert space and we denote its dimension with $\left|\mathcal{H}\right|$, $\mathcal{D}(\mathcal{H})$ to represent the set of all density matrices acting on $\mathcal{H}$.
+For any quantum state $\rho\in\mathcal{D}(\mathcal{H})$, we define $\mbox{supp}(\rho):=\mbox{Span}\{|i\rangle\mid\lambda_{i}>0\},$ where $\left\{\lambda_{i}\right\}$ represent non-zero eigenvalues of $\rho$.
 For any finite and non-empty set $\mathcal{X}$, we denote $\mathcal{P}(\mathcal{X})$ as the set of all probability distributions over $\mathcal{X}$.
-Similarly, for any distribution $P\in\mathcal{P}(\mathcal{X})$, $\mbox{supp}(P):=\{x\in\mathcal{X}:P(x)>0\}.$ For any sequence $x^{n}:(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, a permutation $\pi:\mathcal{X}^{n}\to\mathcal{X}^{n},$ is a map such that $\pi\left(x^{n}\right):=(x_{\pi^{-1}(1)},\cdots,x_{\pi^{-1}(n)}).$ Let $S_{n}$ be the permutation group containing all permutations of length $n$ and it is also known as $n$-th symmetric group.
-Definition 1.
-(Type and Set of all types) A vector of integers ${\bf f}:=\left(f_{1},\cdots,f_{d}\right)$ (where $d>0$) is called a type of size $n$ and length $d$ if $\forall i\in[d],f_{i}\geq 0$ and $\sum_{i}f_{i}=n.$ $T^{n}_{d}$ is denoted as the set of all distinct types of size $n$ and length $d$.
-Definition 2 (Frequency type classes).
-For a non-empty finite set $\mathcal{X}$ and an integer $n>0$, given a type function ${\bf f}:=(f_{1},\cdots,f_{\left|\mathcal{X}\right|}])$ ( such that $\sum_{i\in\mathcal{X}}f_{i}=n$), we define a set $T_{{\bf f}}\subset\mathcal{X}^{n}$ to be the frequency type class or type set corresponding to $\bf f$} as follows,
-$T_{\bf f}:=\left\{(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}:\frac{\left\lvert\left\{k:x_{k}=i\right\}\right\rvert}{n}=\bar{f}(i),\forall i\in\mathcal{X}\right\},$
-where $\bar{f}(i):=\frac{f_{i}}{n}$ for all $i\in\mathcal{X}$.
-Definition 3 ([6]).
-A pair of $n$-length (where $n\in\mathbb{N}$ ) sequences $x^{n}:=(x_{1},\cdots,x_{n}),\tilde{x}^{n}:=(\tilde{x}_{1},\cdots,\tilde{x}_{n})\in\mathcal{X}^{n}$ (where $\left|\mathcal{X}\right|>0$) is said to be $k$-neighbors, represented by $x^{n}\overset{k}{\sim}\tilde{x}^{n}$ if,
-$k=\frac{1}{2}\sum_{a\in\mathcal{X}}\left|f(a|x^{n})-f(a|\tilde{x}^{n})\right|,$ (1)
-where, $f(a|x^{n})$ denotes the number of appearance of the alphabet $a\in\mathcal{X}$ in the sequence $x^{n}$.
-Further, for any sequence $x^{n}$ and its some permutation $\tilde{x}^{n}$, we denote $x^{n}\overset{0}{\sim}\tilde{x}^{n}$.
-Definition 4 (Minimun type-II error).
+Similarly, for any distribution $P\in\mathcal{P}(\mathcal{X})$, we denote $\mbox{supp}(P):=\{x\in\mathcal{X}\mid P(x)>0\}.$
+Definition 2 (Classical minimum type-II error).
 Given a pair of probability distributions $P_{1},P_{2}$ over a finite set $\mathcal{X}$, the minimum type-II error of a fixed order $\alpha\in[0,1]$ is defined as follows,
-$\beta(\alpha,P_{1},P_{2}):=\min_{\begin{subarray}{c}\phi:\mathcal{X}\to[0,1]\\ \sum_{x\in\mathcal{X}}P_{1}(x)\phi(x)\leq\alpha\end{subarray}}1-\sum_{x\in\mathcal{X}}P_{2}(x)\phi(x).$ (2)
+$\beta(\alpha|P_{1}\|P_{2}):=\min_{\begin{subarray}{c}\phi:\mathcal{X}\to[0,1]\\ \sum_{x\in\mathcal{X}}P_{1}(x)\phi(x)\leq\alpha\end{subarray}}1-\sum_{x\in\mathcal{X}}P_{2}(x)\phi(x).$ (2)
 Definition 3 (Classical hypothesis testing divergence).
@@
 where $t]_{+}:=\max\{t,0\}$ denotes the positive part of $t$.
-Definition 7 (Quantum Instrument [22]).
-A quantum instrument consists of a collection $\{\mathcal{E}_{j}\}$ of completely positive, trace non-increasing maps such that the sum map $\sum_{j}\mathcal{E}_{j}$ is trace preserving.
-Let $\{|j\rangle\}_{j}$ be an orthonormal basis for a Hilbert space $\mathcal{H}_{J}$.
-The action of a quantum instrument on a density operator $\rho\in\mathcal{D}(\mathcal{H})$ is the following quantum channel, which features a quantum and classical output:
-$\rho\mapsto\sum_{j}\mathcal{E}_{j}(\rho)\otimes|j\rangle\langle j|_{J}.$ (4)
-Definition 8 (Smooth Max-Relative Entropy).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\varepsilon\in[0,1)$, the smooth max-relative entropy of $\rho$ with respect to $\sigma$ is defined as
-$D_{\max}^{\varepsilon}(\rho\|\sigma):=\min_{\tilde{\rho}\in\mathcal{B}^{\varepsilon}(\rho)}D_{\max}(\tilde{\rho}\|\sigma),$ (5)
-where $D_{\max}(\tilde{\rho}\|\sigma):=\inf\{\lambda\in\mathbb{R}:\tilde{\rho}\leq 2^{\lambda}\sigma\}$ and $\mathcal{B}^{\varepsilon}(\rho):=\left\{\tilde{\rho}\in\mathcal{D}(\mathcal{H}_{A}):\frac{1}{2}\|\tilde{\rho}-\rho\|_{1}\leq\varepsilon\right\}$.
+Definition 5 (Classical alpha divergence).
+For any pair of probability distributions $P,Q$ over a finite set $\mathcal{X}$, and for any $\alpha\in[0,1]$, the classical $\alpha$ divergence of order $\alpha$ is defined as
+$D_{\alpha}(P\|Q):=\frac{1}{\alpha-1}\log\sum_{x\in\mathcal{X}}P(x)^{\alpha}Q(x)^{1-\alpha}.$ (4)
 Definition 6.
 A function $\mathbb{D}:\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})\to\mathbb{R}^{+}$ (where $\mathcal{X}$ is an arbitrary set of finite cardinality) is called a divergence if for any classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$, (where $\mathcal{Y}$ is another arbitrary set of finite cardinality), the following holds,
-$\mathbb{D}(P_{1}\|P_{2})\geq\mathbb{D}(P^{\mathcal{K}}_{1}\|P^{\mathcal{K}}_{2}),\forall P_{1},P_{2}\in\mathcal{P}(\mathcal{X}),$ (6)
-where for each $i=1,2$, $P^{\mathcal{K}}_{i}(y):=\mathbb{E}_{X\sim P_{i}}\left[\mathcal{K}(y\mid X)\right]$.
+$\mathbb{D}(P_{1}\|P_{2})\geq\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2})),\forall P_{1},P_{2}\in\mathcal{P}(\mathcal{X}),$ (5)
+where for each $i=1,2$, $\mathcal{K}(P_{i})(y):=\mathbb{E}_{X\sim P_{i}}\left[\mathcal{K}(y\mid X)\right]$.
+Definition 7 (Quantum minimum type-II error).
+For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the minimum type-II error of a fixed order $\alpha\in[0,1]$ is defined as follows,
+$\beta(\alpha|\rho\|\sigma):=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (6)
+Definition 8 (Quantum hypothesis testing divergence [32]).
+For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, the quantum hypothesis testing divergence $D^{\alpha}_{H}{(\rho\|\sigma)}$ of order $\alpha$ is defined as follows,
+$D^{\alpha}_{H}{(\rho\|\sigma)}:=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}-\ln\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (7)
+Definition 9 (Quantum hockey stick divergence [30]).
+For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we define the quantum hockey stick divergence $E_{\gamma}(\rho\|\sigma)$ of order $\gamma\geq 0$ as follows,
+$E_{\gamma}(\rho\|\sigma):=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}\end{subarray}}\mathrm{Tr}[\Lambda(\rho-\gamma\sigma)].$ (8)
+Definition 10 ([19]).
+Consider $\mathcal{F}$ be the set of functions $f:(0,\infty):\to\mathbb{R}$ that are convex and twice differentiable with $f(1)=0$.
+Then, for any $f\in\mathcal{F}$ and for any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we define the quantum $f$-divergence $D_{f}(\rho\|\sigma)$ as follows,
+$D_{f}(\rho\|\sigma):=\int_{1}^{\infty}\left(f^{\prime\prime}(\gamma)E_{\gamma}(\rho\|\sigma)+\gamma^{-3}f^{\prime\prime}(\gamma^{-1})E_{\gamma}(\sigma\|\rho)\right)d\gamma.$ (9)
+whenever, the integral is finite and $D_{f}(\rho\|\sigma):=+\infty$ otherwise.
+Definition 11.
+Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, then, we define the quantum Divergence between $\rho$ and $\sigma$ as follows
+$D(\rho\|\sigma):=\begin{cases}\mathrm{Tr}[\rho(\log\rho-\log\sigma)],&\text{ if }\rho\ll\sigma,\\ +\infty,&\text{ else}.\end{cases}$
+Definition 12 (Hockey Stick Rényi Divergence [19]).
+Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1)\cup(1,+\infty)$. Then, hockey stick Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
+$D_{\alpha}(\rho\|\sigma)$ $:=\frac{1}{\alpha-1}\log\big(1+(\alpha-1)H_{\alpha}(\rho\|\sigma)\big),$ (10)
+$H_{\alpha}(\rho\|\sigma)$ $:=\alpha\int_{1}^{\infty}\gamma^{\alpha-2}E_{\gamma}(\rho\|\sigma)+\gamma^{-\alpha-1}E_{\gamma}(\sigma\|\rho)\,\mathrm{d}\gamma,$ (11)
+when a parameter $\alpha\in[0,1)\cup(1,+\infty)$ satisfies the condition $\alpha<1\cap\rho\not\perp\sigma$ or the condition $\rho\ll\sigma$.
+Definition 13 (Measured Rényi Divergence [16, 19]).
+Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1)\cup(1,+\infty)$. Then, measured Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
+$\check{D}_{\alpha}(\rho\|\sigma):=\sup_{M:{\rm POVM}}D_{\alpha}(P^{M}_{\rho}\|P^{M}_{\sigma}),$
+where $P^{M}_{\rho}$ is the distribution given as $P^{M}_{\rho}(i):=\mathrm{Tr}\rho M_{i}$ for a POVM $M=\{M_{i}\}_{i}$.
 Definition 14.
@@
 […]
-Definition 11.
-Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, then, we define the quantum Divergence between $\rho$ and $\sigma$ as follows
-$D(\rho\|\sigma):=\begin{cases}\mathrm{Tr}[\rho(\log\rho-\log\sigma)],&\text{ if }\rho\ll\sigma,\\ +\infty,&\text{ else}.\end{cases}$
-Definition 12 (Petz Quantum Rényi Divergence [23]).
-Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and $\alpha\in[0,1]\cup(1,+\infty)$. Then, Petz Quantum Rényi Divergence of order $\alpha$ between $\rho$ and $\sigma$ is defined as follows,
-$D_{\alpha}(\rho\|\sigma):=\begin{cases}\frac{1}{\alpha-1}\log\mathrm{Tr}\left[\rho^{\alpha}\sigma^{1-\alpha}\right],&{\mbox{if }}(\alpha<1\cap\rho\not\perp\sigma)\cup(\rho\ll\sigma),\\ +\infty,&\mbox{else}.\end{cases}$
-Definition 13 (Quantum hockey stick divergence [24]).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we define the quantum hockey stick divergence $E_{\gamma}(\rho\|\sigma)$ of order $\gamma\geq 0$ as follows,
-$E_{\gamma}(\rho\|\sigma):=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}\end{subarray}}\mathrm{Tr}[\Lambda(\rho-\gamma\sigma)].$ (8)
 Fact 1.
-For any type ${\bf f}=(f_{1},\cdots,f_{d})$, the set $T_{\bf f}$ (see Definition 2) is a permutation invariant set, i.e., for every sequence $x^{n}:(x_{1},\cdots,x_{n})\in\mathcal{X}^{n}$, under any permutation $\pi\in S_{n},$ $\pi\left(x^{n}\right)\in T_{\bf f}$ where $\pi_{n}\left(x^{n}\right):=(x_{\pi^{-1}(1)},\cdots,x_{\pi^{-1}(n)}).$
-Fact 2.
-[25] $T^{n}_{d}$ (See Definition 1) satisfies the following upper-bound :
-$\left|\Lambda^{n}_{d}\right|\leq\left|T^{n}_{d}\right|\leq(n+1)^{d-1}.$ (9)
-Fact 3.
 For any two pair $(P,Q)$ and $(P^{\prime},Q^{\prime})$ probability distributions, the following holds for any $\gamma\geq 1$,
-$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\leq E_{\gamma}(P\|Q)+{\left\lVert P^{\prime}-P\right\rVert}_{1}+\gamma{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$ (10)
+$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\leq E_{\gamma}(P\|Q)+\frac{1}{2}{\left\lVert P^{\prime}-P\right\rVert}_{1}+\frac{\gamma}{2}{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$ (13)
 Proof.
 From Definition 4, we have
-$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $=\max_{\begin{subarray}{c}0\leq\Lambda(x)\leq 1,\\ x\in\mathcal{X}\end{subarray}}\sum_{x\in\mathcal{X}}\Lambda(x)(P^{\prime}(x)-\gamma Q^{\prime}(x))$
-$\leq\max_{\begin{subarray}{c}0\leq\Lambda(x)\leq 1,\\ x\in\mathcal{X}\end{subarray}}\sum_{x\in\mathcal{X}}\Lambda(x)(P(x)-\gamma Q(x))+\sum_{x\in\mathcal{X}}|P^{\prime}(x)-P(x)|+\gamma\sum_{x\in\mathcal{X}}|Q^{\prime}(x)-Q(x)|$
-$\leq E_{\gamma}(P\|Q)+{\left\lVert P^{\prime}-P\right\rVert}_{1}+\gamma{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$
-This proves Fact 3.
-Fact 4 (Hölder’s inequality).
+$E_{\gamma}(P^{\prime}\|Q^{\prime})$ $\overset{(a)}{=}P^{\prime}(A)-\gamma Q^{\prime}(A)$
+$\leq P(A)-\gamma Q(A)+P^{\prime}(A)-P(A)+\gamma(Q^{\prime}(A)-Q(A))$
+$\leq\max_{B\subseteq\mathcal{X}}(P(B)-\gamma Q(B))+\max_{C\subseteq\mathcal{X}}(P^{\prime}(C)-P(C))+\gamma\max_{D\subseteq\mathcal{X}}(Q^{\prime}(D)-Q(D))$
+$=E_{\gamma}(P\|Q)+\frac{1}{2}{\left\lVert P^{\prime}-P\right\rVert}_{1}+\frac{\gamma}{2}{\left\lVert Q^{\prime}-Q\right\rVert}_{1}.$
+where in $(a)$, let $A$ be the maximal set for $E_{\gamma}(P^{\prime}\|Q^{\prime})$ according to Definition 4. This proves Fact 1.
+Fact 2 (Hölder’s inequality).
 For any two vectors $x=(x_{1},\dots,x_{n})$ and $y=(y_{1},\dots,y_{n})$ in $\mathbb{R}^{n}$ and for any real numbers $p,q\in[1,\infty)$ such that $\frac{1}{p}+\frac{1}{q}=1$, we have,
 $\sum_{i=1}^{n}\left|x_{i}y_{i}\right|\leq\left(\sum_{i=1}^{n}\left|x_{i}\right|^{p}\right)^{1/p}\left(\sum_{i=1}^{n}\left|y_{i}\right|^{q}\right)^{1/q}.$ (14)
-Further, if $p=1$ and $q=\infty$, then the right-hand side of (11) can be interpreted as follows,
+Further, for the case when $p=1$ and $q=\infty$, the right-hand side of (14) can be interpreted as follows,
 $\left(\sum_{i=1}^{n}\left|x_{i}\right|^{p}\right)^{1/p}\left(\sum_{i=1}^{n}\left|y_{i}\right|^{q}\right)^{1/q}=\left(\sum_{i=1}^{n}\left|x_{i}\right|\right)\max_{i\in[n]}\left|y_{i}\right|.$ (15)
-Fact 5 ([18, Eq. $428$]).
+Fact 3 ([29]).
 Let $P$ and $Q$ be probability distributions on a finite (or measurable) space $\mathcal{X}$.
 Then, the classical relative entropy (Kullback-Leibler divergence) admits the following integral representation in terms of the classical hockey-stick divergence:
-$D(P\|Q)=\int_{0}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(P\|Q)+\frac{1}{\gamma^{2}}E_{\gamma}(Q\|P)\right)d\gamma,$ (13)
+$D(P\|Q)=\int_{1}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(P\|Q)+\frac{1}{\gamma^{2}}E_{\gamma}(Q\|P)\right)d\gamma,$ (16)
 where $E_{\gamma}(P\|Q)$ is the classical hockey-stick divergence as defined in Definition 4.
+Fact 4.
+For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, the quantum hypothesis testing divergence $D^{\alpha}_{H}{(\rho\|\sigma)}$ of order $\alpha$ satisfies the following,
+$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log\beta(\alpha|\rho\|\sigma),$ (17)
+where for any $\alpha\in[0,1]$, $\beta(\alpha|\rho\|\sigma)$ is defined in Definition 7.
+Fact 5 ([19]).
+For any $f\in\mathcal{F}$ (where $\mathcal{F}$ is defined in Definition 10) and for any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we have,
+$D_{f}(\rho\|\sigma):=\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\rho\|\sigma)d\gamma,$ (18)
+where $D_{f}(\rho\|\sigma)$ is defined in Definition 10.
 Fact 6 ([19]).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the quantum relative entropy $D(\rho\|\sigma)$ has the following integral representation in terms of quantum hockey stick divergence,
+For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, and setting $f=x\log x$ in Definition 10, we have the following integral representation of the quantum relative entropy $D(\rho\|\sigma)$,
 $D(\rho\|\sigma)=\int_{0}^{\infty}\left(\frac{1}{\gamma}E_{\gamma}(\rho\|\sigma)+\frac{1}{\gamma^{2}}E_{\gamma}(\sigma\|\rho)\right)d\gamma.$ (19)
-Fact 7.
+Fact 7 ([19]).
+For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we have the following relation between hockey-stick Rényi Divergence $D_{\alpha}(\rho\|\sigma)$ and measured Rényi Divergence $\check{D}_{\alpha}(\rho\|\sigma)$,
+$\check{D}_{\alpha}(\rho\|\sigma)\leq D_{\alpha}(\rho\|\sigma),$ (20)
+for a parameter $\alpha\in[0,1)\cup(1,+\infty)$.
+Fact 8.
+Let $\rho$ and $\sigma$ be two commuting quantum states on a Hilbert space $\mathcal{H}$ (i.e., $\rho,\sigma]=0$).
+For any order $\alpha\in(0,\infty)\setminus\{1\}$, the Hockey Stick Rényi divergence coincides with the Measured Rényi divergence:
+$D_{\alpha}(\rho\|\sigma)=\check{D}_{\alpha}(\rho\|\sigma)=\frac{1}{\alpha-1}\log\left(\sum_{i}p_{i}^{\alpha}q_{i}^{1-\alpha}\right),$ (21)
+where $\{p_{i}\}$ and $\{q_{i}\}$ are the eigenvalues of $\rho$ and $\sigma$ in their common eigen-basis.
+Fact 9.
 For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, we have the following equality,
-$\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]=\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (15)
+$\beta(\alpha|\rho\|\sigma)=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]=\min_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]=\alpha\end{subarray}}\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (22)
 Proof.
 […]
-$\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]$ $=\delta\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$
-$\overset{a}{<}\beta^{\prime},$
-where $a$ follows from the fact that $\delta<1$ and $\beta^{\prime}=\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$.
+$\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]=\delta\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]\overset{(a)}{<}\beta^{\prime},$
+where $(a)$ follows from the fact that $\delta<1$ and $\beta^{\prime}=\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$.
 Thus, for any POVM $\{\tilde{\Lambda},\mathbb{I}-\tilde{\Lambda}\}$, such that $\mathrm{Tr}[\tilde{\Lambda}\rho]<\alpha$, there exists a POVM $\{\hat{\Lambda},\mathbb{I}-\hat{\Lambda}\}$ such that $\mathrm{Tr}[\tilde{\Lambda}\rho]=\alpha$ and $\mathrm{Tr}[(\mathbb{I}-\hat{\Lambda})\sigma]<\mathrm{Tr}[(\mathbb{I}-\tilde{\Lambda})\sigma]$.
 Therefore, the optimum $\Lambda$ in the LHS of (22) satisfies $\mathrm{Tr}[{\Lambda}\rho]=\alpha$.
-This proves Fact 7.
-Fact 8 ([24]).
+This proves Fact 9.
+Fact 10 ([30]).
 For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, the quantum hockey stick divergence $E_{\gamma}(\rho\|\sigma)$ of order $\gamma\geq 0$ (see Definition 9) has the following equivalent form,
-$E_{\gamma}(\rho\|\sigma)=\mathrm{Tr}|\Lambda(\rho-\gamma\sigma)|_{+},$ (16)
+$E_{\gamma}(\rho\|\sigma)=\mathrm{Tr}|(\rho-\gamma\sigma)|_{+},$ (23)
 where $|\cdot|_{+}$ denotes the positive part of the operator, i.e., $|O|_{+}:=\sum_{i}\max\{0,\lambda_{i}\}|\psi_{i}\rangle\langle\psi_{i}|$, where $\lambda_{i}$ and $|\psi_{i}\rangle$ are the eigenvalues and eigenvectors of $O$, respectively.
-Fact 9.
+Fact 11 ([6, Lemma $3.1$]).
+Let $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathbb{C}^{2})$ be commuting qubit quantum states and let $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathbb{C}^{d})$. Then the following two conditions are equivalent:
+  1. 1.
+    There exists a CP-TP map $\mathcal{E}:\mathcal{S}(\mathbb{C}^{2})\to\mathcal{S}(\mathbb{C}^{d})$ such that $\mathcal{E}(\rho_{1})=\rho_{2}$ and $\mathcal{E}(\sigma_{1})=\sigma_{2}$;
+  2. 2.
+    $D_{\max}(\rho_{1}\|\sigma_{1})\geq D_{\max}(\rho_{2}\|\sigma_{2})$ and $D_{\max}(\sigma_{1}\|\rho_{1})\geq D_{\max}(\sigma_{2}\|\rho_{2})$.
+Fact 12.
 For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$ the following equality holds,
@@
 Proof.
-Proof of Fact 9 follows directly from the proof of Fact 7.
-Fact 10 (Reverse quantum data processing inequality).
-Consider $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$.
-Then, if $\forall\gamma\geq 0$, $E_{\gamma}(\rho_{1}\|\rho_{2})\geq E_{\gamma}(\sigma_{1},\sigma_{2})$, then there exists a completely positive map $\mathcal{T}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ such that $\sigma_{i}=\mathcal{T}(\rho_{i})$, for each $i\in\{1,2\}$.
-Proof.
-The proof of Fact 10 follows directly from Theorem 55 and Theorem $6$ of [27].
-Fact 11 (Data-processing inequality of quantum relative entropy).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, $D(\rho\|\sigma)$ the following holds,
-$D(\rho\|\sigma)\geq D(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (18)
-where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
-Fact 12 (Data-processing inequality for quantum hypothesis testing divergence [28]).
-Consider any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$ and any CP-TP map $\mathcal{N}:\mathcal{H}_{A}\to\mathcal{H}_{B}$. Then for any $\alpha\in[0,1]$ the following satisfies,
-$\mathcal{D}^{\alpha}_{H}(\rho\|\sigma)\geq\mathcal{D}^{\alpha}_{H}(\left(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\right).$ (19)
-Fact 13 (Data-processing inequality of Petz quantum Rényi divergence [23]).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ and $\forall\alpha\in[0,1)\cup(1,2]$, $D_{\alpha}(\rho\|\sigma)$ satisfies the following,
-$D_{\alpha}(\rho\|\sigma)\geq D_{\alpha}(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (20)
-where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
-Fact 14 (Data-processing inequality of quantum Hockey stick divergence [24, Lemma $4$]).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ and $\forall k\geq 0$, $E_{k}(\rho\|\sigma)$ satisfies the following,
-$E_{k}(\rho\|\sigma)\geq E_{k}(\mathcal{E}(\rho)\|\mathcal{E}(\sigma)),$ (21)
-where $\mathcal{E}$ is any completely positive and trace-preserving (CP-TP) map.
-Fact 15 (Monotonicity of SLD Fisher Information [29]).
-Let $\rho_{\theta}$ be a differentiable family of quantum states and let $\mathcal{E}$ be a completely positive trace-preserving (CP-TP) map.
-Then, the SLD Fisher information is monotonic under $\mathcal{E}$:
-$J_{\theta}(\rho_{\theta})\geq J_{\theta}(\mathcal{E}(\rho_{\theta})).$ (22)
-Fact 16 (Quantum Reversed Pinsker inequality[30, theorem 2]).
-For quantum states $\rho$ and $\sigma$, the following inequality holds:
-$D(\rho\,\|\,\sigma)\leq\frac{2}{\lambda_{\min}(\sigma)}E_{1}(\rho\,\|\,\sigma)^{2}.$ (23)
-where $\lambda_{\min}(\sigma)$ is the smallest non-zero eigenvalue of $\sigma$.
-Fact 17.
-Consider $\rho,\rho^{\prime},\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, the following inequality holds:
-$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$ (24)
-Proof.
-By the definition of the max-relative entropy, we have
-$\rho^{\prime}\leq e^{D_{\max}(\rho^{\prime}\|\sigma)}\,\sigma.$
-Equivalently,
-$\sigma\geq e^{-D_{\max}(\rho^{\prime}\|\sigma)}\,\rho^{\prime}.$
-Since the logarithm is operator monotone, this implies
-$\log\sigma\succeq\log\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
-Multiplying both sides by $\rho$ and taking the trace (which reverses the inequality), we obtain
-$\mathrm{Tr}[\rho\log\sigma]\leq-\mathrm{Tr}[\rho\log\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
-Adding $\mathrm{Tr}[\rho\log\rho]$ to both sides gives
-$\mathrm{Tr}[\rho(\log\rho-\log\sigma)]\leq\mathrm{Tr}[\rho(\log\rho-\log\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
-which can be written as
-$D(\rho\|\sigma)\leq D(\rho\|\rho^{\prime})+D_{\max}(\rho^{\prime}\|\sigma).$
-This completes the proof of Fact 17.
-Fact 18.
+Proof of Fact 12 follows directly from the proof of Fact 9.
+Fact 13.
 Consider $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$. Then, for any $\varepsilon\geq 0,\delta\in[0,1)$, the following statements are equivalent,
@@
 (2) $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$, where $E_{e^{\varepsilon}}(\cdot\|\cdot)$ is the quantum hockey stick divergence (see Definition 9) of order $e^{\varepsilon}$.
-Proof.
-We first prove that (1) implies (2). From (25), we have that for any $0\preceq\Lambda\preceq\mathbb{I}$,
-$\mathrm{Tr}[\Lambda\rho]-e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]\leq\delta,$
-which can be rewritten as
-$\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
-Taking the maximum over all $0\preceq\Lambda\preceq\mathbb{I}$ on the left-hand side, we get
-$\max_{0\preceq\Lambda\preceq\mathbb{I}}\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
-By the definition of the quantum hockey stick divergence (Definition 13), the left-hand side is exactly $E_{e^{\varepsilon}}(\rho\|\sigma)$. Thus, $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$.
-Next, we prove that (2) implies (1). Assume $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$. By definition, this means
-$\max_{0\preceq\Lambda^{\prime}\preceq\mathbb{I}}\mathrm{Tr}[\Lambda^{\prime}(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
-This implies that for any specific choice of $0\preceq\Lambda\preceq\mathbb{I}$, we must have
-$\mathrm{Tr}[\Lambda(\rho-e^{\varepsilon}\sigma)]\leq\max_{0\preceq\Lambda^{\prime}\preceq\mathbb{I}}\mathrm{Tr}[\Lambda^{\prime}(\rho-e^{\varepsilon}\sigma)]\leq\delta.$
-Rearranging the terms, we get
-$\mathrm{Tr}[\Lambda\rho]\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,$
-which is the condition in (25). This completes the proof.
-Fact 19 ([31, Lemma $6.9$]).
-If $E_{e^{\varepsilon}}(\rho\|\sigma)\leq\delta$, for a pair $(\rho,\sigma)\in\mathcal{D}(\mathcal{H}_{A})$, then,
-$D_{\max}(\tilde{\rho}\|\sigma)\leq\varepsilon-\log\left({1-\delta}\right),$
-where $\varepsilon\geq 0,\delta\in[0,1)$ and $\tilde{\rho}:=\frac{G\rho G^{\dagger}}{\mathrm{Tr}\left[G\rho G^{\dagger}\right]}$, has the property that,
-$\frac{1}{2}{\left\lVert\rho-\tilde{\rho}\right\rVert}_{1}\leq\sqrt{\delta(2-\delta)},$
-where $G:=\left(e^{\varepsilon}\sigma\right)^{\frac{1}{2}}\left(e^{\varepsilon}\sigma+\left|\rho-e^{\varepsilon}\sigma\right|_{+}\right)^{-\frac{1}{2}}.$ Further, if $\operatorname{supp}(\rho)=\operatorname{supp}(\sigma),$ then, $\operatorname{supp}(\tilde{\rho})=\operatorname{supp}(\rho)=\operatorname{supp}(\sigma).$
-III Blackwell’s Dominance of Quantum Mechanisms
-Blackwell [9] in classical statistics provided a framework to compare two statistical experiments in terms of their informativeness.
-An experiment is considered more informatively dominating if it allows for better decision-making based on the observed data.
-This idea of dominance can be viewed from the point of view of hypothesis testing.
-Consider $\mathcal{E}_{1}:=\left\{P_{1},P_{2}\right\}$ and $\mathcal{E}_{2}:=\left\{Q_{1},Q_{2}\right\}$ be two classical statistical experiments, where $P_{1},P_{2}$ and $Q_{1},Q_{2}$ are probability distributions over two finite sets $\mathcal{X}$ and $\mathcal{Y}$.
-Consider the two following Hypothesis tests:
-$H(\mathcal{E}_{1})$ $:=\begin{cases}H^{\mathcal{E}_{1}}_{0}:\mbox{A random variable }X\sim P_{1},\\ H^{\mathcal{E}_{1}}_{1}:\mbox{A random variable }X\sim P_{2}.\end{cases}$
-$H(\mathcal{E}_{2})$ $:=\begin{cases}H^{\mathcal{E}_{2}}_{0}:\mbox{A random variable }Y\sim Q_{1},\\ H^{\mathcal{E}_{2}}_{1}:\mbox{A random variable }Y\sim Q_{2}.\end{cases}$
-We now say that $\mathcal{E}_{2}$ is dominated by $\mathcal{E}_{1}$ (written as $\mathcal{E}_{2}\preceq_{B}\mathcal{E}_{1}$) if for any test (decision rule) $\phi_{\mathcal{E}_{2}}:\mathcal{Y}\to[0,1]$ for $H(\mathcal{E}_{2})$, there exists a test $\phi_{\mathcal{E}_{1}}:\mathcal{X}\to[0,1]$ for $H(\mathcal{E}_{1})$ such that $\phi_{\mathcal{E}_{1}}$ performs at least as well as $\phi_{\mathcal{E}_{2}}$.
-We formalize this notion in the following proposition.
-Proposition 1 (Blackwell’s Theorem of Informative Dominance[10, Theorems $12.4.1$ & $12.4.2$]).
-Given two pairs of probability distributions $\left(P_{1},P_{2}\right)$ and $\left(Q_{1},Q_{2}\right)$ over two finite sets $\mathcal{X}$ and $\mathcal{Y}$, respectively, the following statements are equivalent,
-  1. 1.
-    $\left\{Q_{1},Q_{2}\right\}\preceq_{B}\left\{P_{1},P_{2}\right\}$.
-  2. 2.
-    $\beta(\alpha,Q_{1},Q_{2})\geq\beta(\alpha,P_{1},P_{2})$ for all $\alpha\in[0,1]$, where $\beta(\alpha,P_{1},P_{2})$ and $\beta(\alpha,Q_{1},Q_{2})$ is the minimum type-II error of order $\alpha$ (see Definition 4) of the hypothesis tests $H(\left\{P_{1},P_{2}\right\})$ and $H(\left\{Q_{1},Q_{2}\right\})$ respectively.
-  3. 3.
-    $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}\leq D^{\alpha}_{H}{(P_{1}\|P_{2})}$ for all $\alpha\in[0,1]$, where $D^{\alpha}_{H}{(P_{1}\|P_{2})}$ and $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ are the classical hypothesis testing divergence of order $\alpha$ (see Definition 5) between the pair of distributions $\left\{P_{1},P_{2}\right\}$ and $\left\{Q_{1},Q_{2}\right\}$ respectively.
-  4. 4.
-    There exists a stochastic map $\mathcal{T}:\mathcal{X}\to\mathcal{Y}$ such that $Q_{i}=\mathcal{T}(P_{i})$ for each $i\in\{1,2\}$.
-Observe that Proposition 1 above provides a partial order $\preceq_{B}$ on experiments, i.e., we say an experiment $\mathcal{E}_{2}$ is dominated by (or less informative than) another experiment $\mathcal{E}_{1}$ if $\mathcal{E}_{2}\preceq_{B}\mathcal{E}_{1}$.
-Similarly, in binary asymmetric quantum hypothesis testing, given any pair of quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we perform a hypothesis test between $\rho$, then, the null hypothesis, and $\sigma$, the alternative hypothesis.
-The test is biased towards the null hypothesis, i.e., if the given state is originally $\rho$, then the test should reject the null hypothesis with a small probability, while accepting the null hypothesis with a small probability if the given state is $\sigma$.
-A quantum hypothesis test is typically performed using a binary POVM $\{\Lambda,\mathbb{I}-\Lambda\}$, where $\Lambda$ is the rejection operator and $\mathbb{I}-\Lambda$ is the acceptance operator.
-The hypothesis test can incur two types of errors, namely, the type-I error (false negative) and the type-II error (false positive).
-The type I error is the probability of rejecting the null hypothesis when it is true, while the type II error is the probability of accepting the null hypothesis when it is false.
-Under the binary POVM $\{\Lambda,\mathbb{I}-\Lambda\}$, the type-I error is given by $\mathrm{Tr}[\Lambda\rho]$ and the type-II error is given by $\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma]$.
-To characterize the trade-off between the type-I and type-II errors, in asymmetric quantum hypothesis testing, we fix the type-I error to a value $\alpha\in[0,1]$ and maximize the negative logarithm of type-II error over all possible POVMs $\{\Lambda,\mathbb{I}-\Lambda\}$, which gives us the following definition of quantum hypothesis testing divergence.
-Definition 14 (Quantum hypothesis testing divergence [28]).
-For any $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, for any $\alpha\in[0,1]$, the quantum hypothesis testing divergence $D^{\alpha}_{H}{(\rho\|\sigma)}$ of order $\alpha$ is defined as follows,
-$D^{\alpha}_{H}{(\rho\|\sigma)}:=\max_{\begin{subarray}{c}0\preceq\Lambda\preceq\mathbb{I}:\\ \mathrm{Tr}[\Lambda\rho]\leq\alpha\end{subarray}}-\ln\mathrm{Tr}[(\mathbb{I}-\Lambda)\sigma].$ (26)
-For any pair of quantum states $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A})$, we denote its fixed point as the level of type-I error $\alpha\in[0,1]$ for which the type-II error is equal to the type-I error, i.e., $e^{-D^{\alpha}_{H}{(\rho\|\sigma)}}=\alpha$.
-This is the point where both hypotheses are equally likely.
-Similar to classical Blackwell’s ordering, we define quantum Blackwell’s order as follows,
-Definition 15 (Quantum Blackwell’s order).
-Given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$ over two finite-dimensional Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, we say that $\left\{\sigma_{1},\sigma_{2}\right\}$ is dominated by (or less informative than) $\left\{\rho_{1},\rho_{2}\right\}$ (denoted as $\left\{\sigma_{1},\sigma_{2}\right\}\preceq_{B_{Q}}\left\{\rho_{1},\rho_{2}\right\}$) if for any $\alpha\in[0,1]$, the following holds,
-$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})}.$ (27)
-Lemma 1 (Quantum Blackwell’s Theorem).
-Given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$ over two finite-dimensional Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, the following statements are equivalent,
-  1. 1.
-    $\left\{\sigma_{1},\sigma_{2}\right\}\preceq_{B_{Q}}\left\{\rho_{1},\rho_{2}\right\}$.
-  2. 2.
-    There exists a completely positive map $\mathcal{T}:\mathcal{D}(\mathcal{H}_{A})\to\mathcal{D}(\mathcal{H}_{B})$, such that $\sigma_{1}=\mathcal{T}(\rho_{1})$ and $\sigma_{2}=\mathcal{T}(\rho_{2})$.
-Proof.
-Before proceeding with the proof of Lemma 1, we first prove the following lemma, which will be useful in proving Lemma 1.
-Lemma 2.
-Consider $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$. If
-$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})},\forall\alpha\in[0,1],$ (28)
-then,
-$E_{\gamma}(\rho_{1}\|\rho_{2})\geq E_{\gamma}(\sigma_{1}\|\sigma_{2}),\forall\gamma\geq 0,$
-where $E_{\gamma}(\cdot\|\cdot)$ is defined in Definition 13 for any $\gamma\geq 0$.
-Proof.
-See Appendix -A for the proof.
-It is easy to follow that $(2)\implies(1)$ as a consequence of Definition 15 and Fact 12.
-We now proceed to prove $(1)\implies(2)$.
-From Definition 15 $(1)$ implies (28).
-Further, from Lemma 2 and Fact 10, it follows that, if for the pairs $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$, (28) holds, then there exists a completely positive map $\mathcal{T}:\mathcal{H}_{A}\to\mathcal{H}_{B}$ such that $\sigma_{i}=\mathcal{T}(\rho_{i})$, for each $i\in\{1,2\}$.
-This completes the proof of Lemma 1.
-The concept of Blackwell dominance for pairs of quantum states, as presented in Definition 15, is related to the framework for the statistical comparison of quantum channels developed by Buscemi [11].
-In that work, an informational ordering is defined between two quantum channels, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, based on the guessing probability of any ensemble of states passed through them.
-Specifically, channel $\mathcal{N}_{1}$ is considered more informative than $\mathcal{N}_{2}$ if, for every ensemble of input states, the guessing probability after passing through $\mathcal{N}_{1}$ is at least as high as that after passing through $\mathcal{N}_{2}$.
-This ordering is shown to be equivalent to the existence of a completely positive map that transforms the output of $\mathcal{N}_{1}$ into that of $\mathcal{N}_{2}$, mirroring the condition in Lemma 1.
-However, it remains an open question whether this ordering can be fully characterized by the quantum hypothesis testing divergence, as in Definition 15.
-Addressing this question would deepen our understanding of the relationship between quantum channel comparison and hypothesis testing.
-IV Characteristic Region of $(\varepsilon,\delta)$-Differentially Private Quantum Mechanisms
-In this section, we introduce the concept of the characteristic region of a quantum mechanism that satisfies $(\varepsilon,\delta)$-quantum differential privacy (QDP) as defined in [4, Definition $2$].
+III Characteristic Region of $(\varepsilon,\delta)$-Differentially Private Quantum Mechanisms
+In this section, we introduce the concept of the characteristic region of a quantum mechanism that satisfies $(\varepsilon,\delta)$-quantum differential privacy (QDP) as defined in [36, Definition $2$].
 The characteristic region provides a geometric representation of the trade-offs between the privacy parameters $\varepsilon$ and $\delta$ for a given private quantum mechanism.
-To illustrate this concept, we consider the following definition of a pair of quantum states to be $(\varepsilon,\delta)$-DP.
-Definition 16.
-A pair $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ is defined to be $(\varepsilon,\delta)$-DP (differentially private) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, if every POVM measurement $0\preceq\Lambda\preceq\mathbb{I}$, the following holds,
-$\begin{split}\mathrm{Tr}[\Lambda\rho]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\sigma]+\delta,\\ \mathrm{Tr}[\Lambda\sigma]&\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\rho]+\delta.\end{split}$ (29)
-Further, we denote $\mathfrak{D}_{(\varepsilon,\delta)}$ to be the collection of all pairs of quantum states that satisfy (29).
-Moreover, for $\delta=0$, we denote $(\rho,\sigma)$ to be pure $\varepsilon$-DP or just $\varepsilon$-DP i.e.
-$(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$.
-Consider a hypothesis test between two quantum states $\rho$ and $\sigma$ over a finite dimensional Hilbert space $H$ as follows,
-$H$ $:=\begin{cases}H_{0}:\mbox{Given quantum state is}~\rho,\\ H_{1}:\mbox{Given quantum state is}~\sigma.\end{cases}$
+To illustrate this concept, consider a hypothesis test between two quantum states $\rho$ and $\sigma$ over a finite dimensional Hilbert space $H$ as follows,
+$H$ $:=\begin{cases}H_{0}:\mbox{Given quantum state is}~\rho,\\ H_{1}:\mbox{Given quantum state is}~\sigma.\end{cases}$ (26)
 Under the choice of a rejection POVM $0\preceq\Lambda\preceq\mathbb{I}$ for $H_{0}$, the type-I error is given by $\alpha_{\Lambda}:=\mathrm{Tr}\left[\Lambda\rho\right]$ and the type-II error is given by $\beta_{\Lambda}:=1-\mathrm{Tr}\left[\Lambda\sigma\right]$.
@@
 $\mathcal{R}(\rho,\sigma):=\left\{(\alpha_{\Lambda},\beta_{\Lambda}):0\preceq\Lambda\preceq\mathbb{I}\right\}.$ (27)
-Note that from Definition 15 it follows that for any two pairs of quantum states $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, we have $\{\sigma_{1},\sigma_{2}\}\preceq_{B_{Q}}\{\rho_{1},\rho_{2}\}$ if and only if $\mathcal{R}(\sigma_{1},\sigma_{2})\subseteq\mathcal{R}(\rho_{1},\rho_{2})$.
 Further, if $(\rho,\sigma)$ is $(\varepsilon,\delta)$-QDP, then for any rejection POVM $\Lambda$, from (1) the following constraints hold,
-$\beta_{\Lambda}$ $\geq e^{-\varepsilon}(1-\delta-\alpha_{S}),$ (32)
-$\beta_{\Lambda}$ $\geq 1-\delta-e^{\varepsilon}\alpha_{S},$ (33)
-$\beta_{\Lambda}$ $\leq 1-e^{-\varepsilon}(\alpha_{S}-\delta),$ (34)
-$\beta_{\Lambda}$ $\leq e^{\varepsilon}(1-\alpha_{S})+\delta.$ (35)
+$\beta_{\Lambda}$ $\geq e^{-\varepsilon}(1-\delta-\alpha_{\Lambda}),$ (28)
+$\beta_{\Lambda}$ $\geq 1-\delta-e^{\varepsilon}\alpha_{\Lambda},$ (29)
+$\beta_{\Lambda}$ $\leq 1-e^{-\varepsilon}(\alpha_{\Lambda}-\delta),$ (30)
+$\beta_{\Lambda}$ $\leq e^{\varepsilon}(1-\alpha_{\Lambda})+\delta.$ (31)
 $\alpha\longrightarrow$
@@
 Figure 1: Graphical representation of $\mathcal{R}(\varepsilon,\delta)$ : characteristic region of $(\varepsilon,\delta)$-QDP (in shaded region), where we define the boundaries (a), (b), (c) and (d) to be $\beta=1-\delta-e^{\varepsilon}\alpha$, $\beta=e^{-\varepsilon}(1-\delta-\alpha)$, $\beta=1-e^{-\varepsilon}(\alpha-\delta)$ and $\beta=e^{\varepsilon}(1-\alpha)+\delta$ respectively and $\mathcal{R}(\varepsilon,\delta)$ has two fixed points $(\frac{1-\delta}{1+e^{\varepsilon}},\frac{1-\delta}{1+e^{\varepsilon}})$ and $(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}})$ as its extremal points, where the former and the latter points are known to be the worst and the best fixed points respectively from the perspective of privacy.
-1n Figure 1, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ of $(\varepsilon,\delta)$-QDP.
-Therefore, under the constraints mentioned in eqs.˜32, 33, 34 and 35, we define the characteristic (or operating) region of quantum $(\varepsilon,\delta)$-QDP as follows,
+In Figure 1, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ of $(\varepsilon,\delta)$-QDP.
+Therefore, under the constraints mentioned in eqs.˜28, 29, 30 and 31, we define the characteristic (or operating) region of quantum $(\varepsilon,\delta)$-QDP as follows,
 Definition 16 (Characteristic region of $(\varepsilon,\delta)$-QDP).
@@
 in [18] plotted $\mathcal{R}(\varepsilon,\delta)$ for certain values of $\varepsilon,\delta$.
-From eqs.˜32, 33, 34 and 35, it is trivial to see that for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$), we have $\mathcal{R}(\rho,\sigma)\subseteq\mathcal{R}(\varepsilon,\delta)$.
+From eqs.˜28, 29, 30 and 31, it is trivial to see that for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$), we have $\mathcal{R}(\rho,\sigma)\subseteq\mathcal{R}(\varepsilon,\delta)$.
 Moreover, the characteristic region $\mathcal{R}(\varepsilon,\delta)$ is a convex set.
@@
 we show that for $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$, under certain choices of measurements, the pair of type-I ($\alpha$) and type-II errors ($\beta$) achieves the extremal/corner points of $\mathcal{R}(\varepsilon,\delta)$.
-Consider that $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ are two quantum states over a finite-dimensional Hilbert space $\mathcal{H}$ such that for an one-dimensional projector $|v\rangle\langle v|$ we have,
+Consider a pair $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ over a finite-dimensional Hilbert space $\mathcal{H}$ such that for a one-dimensional projector $|v\rangle\langle v|$,
 $\mathrm{Tr}[|v\rangle\langle v|\rho_{(\varepsilon,\delta)}]$ $\triangleq\delta,$ (33)
@@
 Similarly, for the one-dimensional projector $\mathbb{I}-|x\rangle\langle x|$, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the corner point $(1,\delta)$ of $\mathcal{R}(\varepsilon,\delta)$.
-Now, we notice that the cardinality of the support of the state pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ must be at least $3$ since we need two more corner points to be achieved.
+Now, we notice that the cardinality of the support of the state pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ must be at least $3$, since we need two more corner points to be achieved.
 Thus, we consider another one-dimensional projector $\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|$ which is perpendicular to both $|v\rangle\langle v|$ and $|x\rangle\langle x|$.
@@
 Thus, the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ achieves the best fixed point $(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}},\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}})$ of $\mathcal{R}(\varepsilon,\delta)$ with the help of the two-dimensional projector $|v\rangle\langle v|+|y\rangle\langle y|$.
-Now eqs.˜37, 38, 43 and 44 gives us,
+Now eqs.˜33, 34, 39 and 40 gives us,
 $\mathrm{Tr}[|y\rangle\langle y|\rho_{(\varepsilon,\delta)}]$ $=\frac{e^{\varepsilon}(1-\delta)}{1+e^{\varepsilon}},$ (41)
@@
 Finally, consider the one-dimensional projector $|z\rangle\langle z|=\mathbb{I}-|v\rangle\langle v|-|x\rangle\langle x|-|y\rangle\langle y|$ which is perpendicular to $|v\rangle\langle v|$, $|x\rangle\langle x|$ and $|y\rangle\langle y|$.
-Then, from eqs.˜37, 38, 39, 40, 45 and 46 we have,
+Then, from eqs.˜33, 34, 35, 36, 41 and 42 we have,
 $\mathrm{Tr}[|z\rangle\langle z|\rho_{(\varepsilon,\delta)}]$ $=\frac{1-\delta}{1+e^{\varepsilon}},$ (43)
@@
 $\sigma_{(\varepsilon,\delta)}$ $=(1-\delta)\left(\frac{1}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|10\rangle\langle 10|\right)+\delta|11\rangle\langle 11|.$ (48)
-Remark 1.
-  • It is important to note that in (43) and (44), if we had chosen the RHS to be $(\frac{1-\delta}{1+e^{\varepsilon}}$ and $\frac{1-\delta}{1+e^{\varepsilon}}$ we would get another pair of quantum states $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ which also achieves the extremal points of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ for any given $\varepsilon\geq 0$ and $\delta\in[0,1]$.
-This pair is given as follows,
-    $\rho^{\prime}_{(\varepsilon,\delta)}$ $=\delta|00\rangle\langle 00|+\frac{1-2\delta-e^{\varepsilon}\delta}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}|10\rangle\langle 10|,$ (53)
-    $\sigma^{\prime}_{(\varepsilon,\delta)}$ $=\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{1-2\delta-e^{\varepsilon}\delta}{1+e^{\varepsilon}}|10\rangle\langle 10|+\delta|11\rangle\langle 11|.$ (54)
-    However, the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ also achieves two points $\left(\frac{1-\delta}{1+e^{\varepsilon}}-\delta,\frac{1-\delta}{1+e^{\varepsilon}}-\delta\right)$ and $\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta,\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta\right)$, which are outside the characteristic region $\mathcal{R}(\varepsilon,\delta)$.
-From this, we can conclude that the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})\not\in\mathfrak{D}_{(\varepsilon,\delta)}$.
-  • Further, one can verify that the above pair satisfies $(\log(\frac{e^{\varepsilon}}{1-\delta(2+e^{\varepsilon})}),\delta)$-DP.
-In Figure 2 below, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ along with the extremal points achieved by the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$.
 $\alpha\longrightarrow$
@@
 Figure 2: Graphical representation of $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ (in the whole shaded region) which is strictly larger than $\mathcal{R}(\varepsilon,\delta)$ (in inner shaded region), where the two extremal points of $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ outside $\mathcal{R}(\varepsilon,\delta)$ are $\left(\frac{1-\delta}{1+e^{\varepsilon}}-\delta,\frac{1-\delta}{1+e^{\varepsilon}}-\delta\right)$ and $\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta,\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta\right)$.
-V Differential Privacy : A Quantum Hypothesis Testing Perspective
-V-A Main characterization
-In this section, we consider differential privacy from a quantum hypothesis testing perspective.
+Remark 1.
+  • It is important to note that in (39) and (40), if we had chosen the RHS to be $\frac{1-\delta}{1+e^{\varepsilon}}$ and $\frac{1-\delta}{1+e^{\varepsilon}}$ respectively, we would get another pair of quantum states $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ which also achieves the extremal points of the characteristic region $\mathcal{R}(\varepsilon,\delta)$ for any given $\varepsilon\geq 0$ and $\delta\in[0,1]$.
+This pair is given as follows,
+    $\rho^{\prime}_{(\varepsilon,\delta)}$ $=\delta|00\rangle\langle 00|+\frac{1-2\delta-e^{\varepsilon}\delta}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}|10\rangle\langle 10|,$ (49)
+    $\sigma^{\prime}_{(\varepsilon,\delta)}$ $=\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}|01\rangle\langle 01|+\frac{1-2\delta-e^{\varepsilon}\delta}{1+e^{\varepsilon}}|10\rangle\langle 10|+\delta|11\rangle\langle 11|.$ (50)
+    However, the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ also achieves two points $\left(\frac{1-\delta}{1+e^{\varepsilon}}-\delta,\frac{1-\delta}{1+e^{\varepsilon}}-\delta\right)$ and $\left(\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta,\frac{e^{\varepsilon}+\delta}{1+e^{\varepsilon}}+\delta\right)$, which are outside the characteristic region $\mathcal{R}(\varepsilon,\delta)$.
+From this, we can conclude that the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})\not\in\mathfrak{D}_{(\varepsilon,\delta)}$.
+  • Further, one can verify that the above pair satisfies $(\log(\frac{e^{\varepsilon}}{1-\delta(2+e^{\varepsilon})}),\delta)$-DP.
+In Figure 2 above, we illustrate a graphical representation of the characteristic region $\mathcal{R}(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$ along with the extremal points achieved by the pair $(\rho^{\prime}_{(\varepsilon,\delta)},\sigma^{\prime}_{(\varepsilon,\delta)})$.
+Remark 2.
+In [5], the authors obtained a figure similar to Figure 1 by plotting type-1 and type-2 errors for a pair of states $\{\rho_{1},\rho_{2}\}$.
+In contrast, Figure 1 is with respect to all pairs $\{\rho_{1},\rho_{2}\}$ which are $(\varepsilon,\delta)$-DP and therefore it helps to characterize the worst-case (most informative) $(\varepsilon,\delta)$-DP pair of quantum states.
+IV Quantum Information Ordering and Differential Privacy
+IV-A Quantum ordering of informativeness
+In [4, 3], Blackwell introduced a notion of an ordering of classical informativeness between two pairs of distributions $\{P_{1},P_{2}\}$ and $\{Q_{1},Q_{2}\}$.
+In particular, he proved the following proposition,
+Proposition 1 (Blackwell’s Theorem of Informativeness Ordering [3, Theorems $12.2.2$ & $12.4.2$]).
+Given two pairs of probability distributions $\left(P_{1},P_{2}\right)$ and $\left(Q_{1},Q_{2}\right)$ over two finite sets $\mathcal{X}$ and $\mathcal{Y}$, respectively, the following statements are equivalent,
+  1. 1.
+    $\left\{P_{1},P_{2}\right\}\succeq_{B}\left\{Q_{1},Q_{2}\right\}$.
+  2. 2.
+    $D^{\alpha}_{H}{(P_{1}\|P_{2})}\geq D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ for all $\alpha\in[0,1]$, where $D^{\alpha}_{H}{(P_{1}\|P_{2})}$ and $D^{\alpha}_{H}{(Q_{1}\|Q_{2})}$ are the classical hypothesis testing divergence of order $\alpha$ (see Definition 3) between the pair of distributions $\left\{P_{1},P_{2}\right\}$ and $\left\{Q_{1},Q_{2}\right\}$ respectively.
+  3. 3.
+    There exists a stochastic map $\mathcal{T}:\mathcal{X}\to\mathcal{Y}$ such that $Q_{i}=\mathcal{T}(P_{i})$ for each $i\in\{1,2\}$.
+Analogous to Blackwell’s order of classical informativeness (as mentioned in Proposition 1), we discuss a quantum version of it, which provides an ordering of quantum statistical experiments based on their distinguishability.
+Using the quantum hypothesis testing divergence as the underlying measure, we formalize when one pair of quantum states can be considered more informative than another.
+We further establish that this dominance relation implies a corresponding ordering across all quantum $f$-divergences, thereby linking hypothesis testing quantities to a broad class of information-theoretic measures.
+We now make the following definition.
+Definition 17 (Quantum information order).
+Consider two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$.
+We say that the pair $\left(\rho_{1},\rho_{2}\right)$ dominates (or, is more informative than) the pair $\left(\sigma_{1},\sigma_{2}\right)$ (denoted as $\left\{\rho_{1},\rho_{2}\right\}\succeq_{Q}\left\{\sigma_{1},\sigma_{2}\right\}$), if for all $\alpha\in[0,1]$, the following holds,
+$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})}.$ (51)
+Intuitively, the order $\left\{\rho_{1},\rho_{2}\right\}\succeq_{Q}\left\{\sigma_{1},\sigma_{2}\right\}$ indicates that the states $\rho_{1}$ and $\rho_{2}$ are more distinguishable as compared to the states $\sigma_{1}$ and $\sigma_{2}$.
+[…]
+However, we can circumvent this issue by directly showing that if for any two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, (51) holds, then, every quantum $f$-divergence (see Definition 10) between the former pair dominates that of the latter pair.
+We formalize this claim in the Theorem below.
+Theorem 1.
+Consider two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, for which the following holds,
+$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})},\forall\alpha\in[0,1].$ (52)
+Then, for any $f\in\mathcal{F}$ (see Definition 10) we have,
+$D_{f}(\rho_{1}\|\rho_{2})\geq D_{f}(\sigma_{1}\|\sigma_{2}),$
+whenever $D_{f}$ is well-defined (see Definition 10) for the pairs $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$.
+Before proceeding with the proof of Theorem 1, we first state the following lemma, which will be required in proving Theorem 1.
+Lemma 1.
+Consider $\rho_{1},\rho_{2}\in\mathcal{D}(\mathcal{H}_{A})$ and $\sigma_{1},\sigma_{2}\in\mathcal{D}(\mathcal{H}_{B})$. If
+$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})},\forall\alpha\in[0,1],$ (53)
+then,
+$E_{\gamma}(\rho_{1}\|\rho_{2})$ $\geq E_{\gamma}(\sigma_{1}\|\sigma_{2}),\forall\gamma\geq 0,$ (54)
+$D_{\tilde{\alpha}}(\rho_{1}\|\rho_{2})$ $\geq D_{\tilde{\alpha}}(\sigma_{1}\|\sigma_{2}),\forall\tilde{\alpha}\in[0,1)\cup(1,+\infty),$ (55)
+$D(\rho_{1}\|\rho_{2})$ $\geq D(\sigma_{1}\|\sigma_{2}),$ (56)
+[…]
+Proof.
+See Appendix -A for the proof.
+Using the above lemma, we now prove Theorem 1 using the following series of inequalities,
+$D_{f}(\rho_{1}\|\rho_{2})$ $\overset{(a)}{=}\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\rho_{1}\|\rho_{2})d\gamma$
+$\overset{(b)}{\geq}\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\sigma_{1}\|\sigma_{2})d\gamma$
+$\overset{(c)}{=}D_{f}(\sigma_{1}\|\sigma_{2}),$
+where $(a)$ and $(c)$ follows from Fact 5 and $(b)$ follows from (54) of Lemma 1 and the fact that the double derivative of $f$ exists and $f$ is a convex function.
+This completes the proof of Theorem 1.
+Remark 3.
+We note here that (54) of Lemma 1 was also proven in [5, Theorem 22].
+Remark 4.
+There have been various notions of ordering of informativeness in the quantum case in the earlier works, see for example, [28, 8, 21, 24, 9, 10, 7, 33].
+However, it is not very clear whether these orderings are useful for characterizing differentially private mechanisms, which is the main aim of this manuscript.
+IV-B Differential privacy: A quantum hypothesis testing perspective
+In this subsection, we consider differential privacy from a quantum hypothesis testing perspective.
 The main goal of quantum differential privacy is to ensure that a quantum mechanism produces output states that are nearly indistinguishable when applied to neighboring quantum inputs.
@@
 More formally, with high probability, no hypothesis test—regardless of the investigator’s strategy—can reliably infer the respondent’s individual contribution from the output.
-Further, it is important to observe that given two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, if for any $\alpha\in[0,1]$ the hypothesis testing divergence satisfies if $D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})}$, then $(\sigma_{1},\sigma_{2})$ is harder to distinguish than $(\rho_{1},\rho_{2})$ for all possible values of the Type-I error parameter $\alpha$.
-Intuitively, this indicates that the states $\rho_{1}$ and $\rho_{2}$ are more distinguishable—or equivalently, more well-separated—than the states $\sigma_{1}$ and $\sigma_{2}$.
-This observation naturally leads to the following result.
-Theorem 1.
-Consider two pairs of quantum states $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$, for which the following holds,
-$D^{\alpha}_{H}{(\rho_{1}\|\rho_{2})}\geq D^{\alpha}_{H}{(\sigma_{1}\|\sigma_{2})},\forall\alpha\in[0,1].$ (55)
-[…]
-[…]
-Proof.
-[…]
-Motivated by the close relation between privacy and hypothesis testing in the lemma below, we give an equivalent condition for a pair of quantum states to be $(\varepsilon,\delta)$-DP.
+Motivated by the close relation between privacy and hypothesis testing, in the lemma below, we give an equivalent condition for a pair of quantum states to be $(\varepsilon,\delta)$-DP (see Definition 1).
 Lemma 2.
-A pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 16) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1)$ if and only if the following holds,
-$D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1].$ (56)
-where $f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1]$.
+Consider any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 1) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1)$. Then, the following holds,
+$D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1],$ (57)
+or, equivalently,
+$\beta(\alpha|\rho\|\sigma)\geq f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1],$ (58)
+where,
+$f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\right\},$ (59)
+for any $\alpha\in[0,1]$.
 Proof.
 See Appendix -B for the proof.
-Note that the above version of quantum $(\varepsilon,\delta)$-DP implies that if for each $\alpha\in[0,1]$ $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}(\alpha)$ for some pair of quantum states $\rho^{\prime},\sigma^{\prime}$ over any arbitrary finite dimension Hilbert space, then for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 16) is at least as hard as distinguishing between the pair $(\rho^{\prime},\sigma^{\prime})$.
+Note that the above version of quantum $(\varepsilon,\delta)$-DP implies that if for each $\alpha\in[0,1]$ $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}(\alpha)$ for some pair of quantum states $\rho^{\prime},\sigma^{\prime}$ over any arbitrary finite dimensional Hilbert space, then for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 1) is at least as hard as distinguishing between the pair $(\rho^{\prime},\sigma^{\prime})$.
 The above intuition provides a notion of the weakest (most informative) quantum $(\varepsilon,\delta)$-DP pairs of quantum states, which are the least “hard to differentiate”.
@@
 For $\varepsilon\geq 0$ and $\delta\in[0,1),$ a pair of quantum states $(\rho,\sigma)$ is defined to be the weakest (most informative) quantum $(\varepsilon,\delta)$-DP if
-$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1].$ (57)
-where $f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1]$.
-[…]
-Theorem 2.
-[…]
-[…]
-where $(\rho^{\prime},\sigma^{\prime})$ is some pair of quantum states which is the weakest (most informative) $(\varepsilon,\delta)$-DP (see Definition 19).
-Proof.
-If a pair $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then, from Lemma 3 and the fact that $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}$ for every $\alpha\in[0,1]$, the following holds,
-$D^{\alpha}_{H}(\rho\|\sigma)\leq D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime}),\forall\alpha\in[0,1].$ (59)
-[…]
-In the lemma below, we show the existence of a pair of quantum states that is the weakest (most informative) quantum $(\varepsilon,\delta)$-DP (see Definition 19) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$ in the Blackwell sense.
+$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1],$ (60)
+In the lemma below, we show the existence of a pair of quantum states that is the weakest (most informative) quantum $(\varepsilon,\delta)$-DP (see Definition 18) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$.
 Lemma 3.
 Consider the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in (47) and (48), for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$.
-Then, we have $D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=-\log\max\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\}$ for any $\alpha\in[0,1]$.
+Then, we have $D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=-\log f_{\varepsilon,\delta}(\alpha)$ for any $\alpha\in[0,1]$.
 Proof.
 See Appendix -C for the proof.
-V-B Blackwell Dominance under Quantum Differential Privacy
-From Lemmas 3 and 4 and Definition 15, we note that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, gets dominated by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
+Therefore, for a pair $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ with some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, Lemmas 2 and 3 guarantee the relation,
+$D^{\alpha}_{H}(\rho\|\sigma)$ $\leq-\log f_{\varepsilon,\delta}(\alpha)=D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (61)
+for every $\alpha\in[0,1]$.
+Further, from Lemmas 2 and 3 and Definition 17, we note that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, gets dominated by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
 Formally, it implies the following.
@@
 If any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then the following holds,
-$\left\{(\rho,\sigma)\right\}\preceq_{B_{Q}}\left\{(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\right\}.$ (60)
+$\left\{\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)}\right\}\succeq_{Q}\left\{\rho,\sigma\right\}.$ (62)
 Remark 5.
-  • Note that the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in Lemma 4 satisfies weakest (most informative) $(\varepsilon,\delta)$-DP condition for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$ and therefore, they have a fixed-point at $\alpha=\frac{1-\delta}{e^{\varepsilon}+1}$.
-  • For $\delta=0$ and any $\varepsilon\geq 0$, we define two quantum states $\rho_{\varepsilon}:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|$ and $\sigma_{\varepsilon}:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|$.
-Observe that, it can be verified that $\rho_{\varepsilon}$ and $\sigma_{\varepsilon}$ are weakest (most informative) $(\varepsilon,0)$-DP states.
+  • Note that the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in Lemma 3, which satisfies weakest (most informative) $(\varepsilon,\delta)$-DP condition for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, has a fixed-point at $\alpha=\frac{1-\delta}{e^{\varepsilon}+1}$.
+  • For $\delta=0$ and any $\varepsilon\geq 0$, we define two quantum states,
+    $\rho_{(\varepsilon)}$ $:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|,$ (63)
+    $\sigma_{(\varepsilon)}$ $:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|,$ (64)
+    Observe that it can be verified that $\rho_{(\varepsilon)}$ and $\sigma_{(\varepsilon)}$ are the weakest (most informative) $\varepsilon$-DP pair of quantum states.
+We now state the main theorem of this section, which provides a bound on $D_{f}(\rho\|\sigma)$ (see Definition 10) for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$.
+Theorem 2.
+If a pair of quantum states $(\rho,\sigma)$ belongs to $\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then for any $f\in\mathcal{F}$, we have,
+$D_{f}(\rho\|\sigma)\leq D_{f}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (65)
+whenever $D_{f}$ is well-defined for the pairs $(\rho_{1},\rho_{2})$ and $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
+Also, for any $\gamma\geq 0$ and $\alpha\in[0,1)\cup(1,+\infty)$, we have,
+$E_{\gamma}(\rho\|\sigma)$ $\leq E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (66)
+$D_{\alpha}(\rho\|\sigma)$ $\leq D_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (67)
+$D(\rho\|\sigma)$ $\leq D(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (68)
+$\check{D}_{\alpha}(\rho\|\sigma)$ $\leq\check{D}_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}).$ (69)
+Proof.
+Suppose a pair $(\rho,\sigma)$ belongs to $\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$.
+Then, the combination of (61) and Theorem 1 implies the relation $D_{f}(\rho\|\sigma)\leq D_{f}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$.
+The combination of (61) and Lemma 1 also implies (66), (67), and (68).
+Finally, using Fact 7, we have,
+$\check{D}_{\alpha}(\rho\|\sigma)\leq D_{\alpha}(\rho\|\sigma)\leq D_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=\check{D}_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (70)
+where the final equality follows from Fact 8, since $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ are commutative with each other. This proves Theorem 2.
+Further, for $\delta=0$, the theorem below gives an upper-bound on $D_{f}$ (for any $f\in\mathcal{F}$) between any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$.
+Corollary 2.
+If any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$ for some fixed $\varepsilon\geq 0$, then for any $f\in\mathcal{F}$, $\gamma\geq 0$ and $\alpha\in[0,1)\cup(1,+\infty)$, we have,
+$D_{f}(\rho\|\sigma)$ $\leq D_{f}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$ (71)
+$E_{\gamma}(\rho\|\sigma)$ $\leq E_{\gamma}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$ (72)
+$D_{\alpha}(\rho\|\sigma)$ $\leq D_{\alpha}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$ (73)
+$\check{D}_{\alpha}(\rho\|\sigma)$ $\leq\check{D}_{\alpha}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}).$ (74)
+whenever $D_{f}$ is well-defined for the pairs $(\rho,\sigma)$ and $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ (defined in (63) and (64) respectively).
+Remark 6.
+Note that for some fixed $\varepsilon\geq 0$, the pair $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ is commutative and $D_{\max}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)})=D_{\max}(\sigma_{(\varepsilon)}\|\rho_{(\varepsilon)})$ $\varepsilon$.
+Further, Definition 1 directly guarantees that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$ satisfies
+$D_{\max}(\rho\|\sigma)$ $\leq\varepsilon=D_{\max}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$
+$D_{\max}(\sigma\|\rho)$ $\leq\varepsilon=D_{\max}(\sigma_{(\varepsilon)}\|\rho_{(\varepsilon)}).$
+Therefore, from Fact 11, there exists a CP-TP map $\mathcal{E}$, such that $\rho=\mathcal{E}(\rho_{(\varepsilon)})$ and $\sigma=\mathcal{E}(\sigma_{(\varepsilon)})$. Hence, Definition 14 implies
 […]
-Corollary 2.
-[…]
-[…]
-The reference [19, Theorem 18] showed Corollary 2 in the classical case.
-Corollary 3.
-[…]
-[…]
-where $\rho_{(\varepsilon)}:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|$ and $\sigma_{(\varepsilon)}:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|$.
-Now using Corollary 2, for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, we can derive upper-bounds in terms of quantum hockey-stick divergence.
-However, it is important to note that for $\delta>0$, $\operatorname{supp}(\rho_{(\varepsilon,\delta)})\not\subseteq\operatorname{supp}(\sigma_{(\varepsilon,\delta)})$.
+It is important to note that for $\delta>0$, $\operatorname{supp}(\rho_{(\varepsilon,\delta)})\not\subseteq\operatorname{supp}(\sigma_{(\varepsilon,\delta)})$.
 Therefore, the divergence that requires support inclusion is not well-defined for the pair of quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$.
-Although for $\delta=0$, using Corollary 3, since $\operatorname{supp}(\rho_{\varepsilon})=\operatorname{supp}(\sigma_{\varepsilon})$, for any $\varepsilon,0$-DP quantum state pairs, we can derive upper-bounds on quantum relative entropy, quantum Rényi divergence in the corollary below.
+Although for the case when $\delta=0$, since $\operatorname{supp}(\rho_{\varepsilon})=\operatorname{supp}(\sigma_{\varepsilon})$, for any $\varepsilon$-DP quantum state pairs, using Corollary 2, we derive upper-bounds on quantum relative entropy and $L_{1}$ distance as mentioned in the corollary below.
 Corollary 3.
@@
   2. (ii)
-    $D_{\alpha}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\left[\log\left(e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}\right)-\log\left(1+e^{\varepsilon}\right)\right]$,  for all $\alpha\in[0,1]\cup(1,2)$ ,
-  3. (iii)
-    ${\left\lVert\rho-\sigma\right\rVert}_{1}\leq\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)$.
+    ${\left\lVert\rho-\sigma\right\rVert}_{1}\leq 2\tanh\left(\frac{\varepsilon}{2}\right)$.
 Proof.
-(i) From Fact 11 it follows that $D(\cdot\|\cdot)$ is a valid divergence and thus, we can write the following,
-$D(\rho^{\prime}\|\sigma^{\prime})$ $=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\varepsilon-\frac{1}{1+e^{\varepsilon}}\varepsilon$
-$=\varepsilon\left(\frac{e^{\frac{\varepsilon}{2}}-e^{-\frac{\varepsilon}{2}}}{e^{\frac{\varepsilon}{2}}+e^{-\frac{\varepsilon}{2}}}\right)$
-$=\varepsilon\tanh\left(\frac{\varepsilon}{2}\right).$ (63)
-Therefore, (61) of Corollary 2 completes the proof of 1) of Corollary 4.
-(ii) From Fact 13 it follows that for $\alpha\in[0,1]\cup(1,2]$, $D_{\alpha}(\cdot\|\cdot)$ is a valid divergence and thus, we can write the following,
-$D_{\alpha}(\rho^{\prime}\|\sigma^{\prime})$ $=\frac{1}{\alpha-1}\log\left(\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\right)^{\alpha}\left(\frac{1}{1+e^{\varepsilon}}\right)^{1-\alpha}+\left(\frac{1}{1+e^{\varepsilon}}\right)^{\alpha}\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\right)^{1-\alpha}\right)$
-$=\frac{1}{\alpha-1}\log\left(\frac{e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}}{1+e^{\varepsilon}}\right)$
-$=\frac{1}{\alpha-1}\left[\log\left(e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}\right)-\log\left(1+e^{\varepsilon}\right)\right].$ (64)
-Therefore, (61) of Corollary 2 completes the proof of 2) of Corollary 4.
-(iii) As ${\left\lVert\mathcal{N}(\rho)-\mathcal{N}(\sigma)\right\rVert}_{1}=2E_{1}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))$, from Fact 14 it follows that ${\left\lVert\cdot-\cdot\right\rVert}_{1}$ is a valid divergence and thus, we can write the following,
-${\left\lVert\rho^{\prime}-\sigma^{\prime}\right\rVert}_{1}$ $=\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|+\left|\frac{1-e^{\varepsilon}}{1+e^{\varepsilon}}\right|$
-$\overset{a}{=}2\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|$
-$=2\tanh\left(\frac{\varepsilon}{2}\right).$ (65)
-Therefore, (61) of Corollary 2 completes the proof of 3) of Corollary 4.
-VI Quantum Privatized Parameter Inference
-In the context of randomized response, a fundamental challenge is balancing the privacy of respondents with the statistical utility of the information gathered by an investigator.
-Let’s consider a scenario where binary information, represented by $\{0,1\}$, is encoded into two quantum states, $\rho$ and $\sigma$.
-The privacy of this encoding is quantified by the $(\varepsilon,\delta)$-DP constraint on the pair $(\rho,\sigma)$.
-We assume that the behavior of the respondents is considered as random sampling with replacement from the investigator’s perspective, and the investigator is interested only in the ratio between 0 and $1$ so that the underlying statistical information $\theta$ follows a binomial distribution $p_{\theta}$, where $p_{\theta}(0)=\theta$ and $p_{\theta}(1)=1-\theta$.
-That is, the key task is the inference of the parameter $\theta$ by the investigator.
-With the quantum encoding, this translates to a parametrized state $\rho_{\theta}:=\theta\rho+(1-\theta)\sigma$, where $\theta\in[0,1]$.
-If the investigator collects $n$ responses, the composite state is $\rho_{\theta}^{\otimes n}$.
-More generally, in quantum parameter estimation, the goal is to estimate an unknown parameter $\theta$ encoded in a family of quantum states $\{\rho_{\theta}\}$, where the encoding is constrained by privacy requirements.
-In the privatized setting, the states $\rho$ and $\sigma$ must satisfy the $(\varepsilon,\delta)$-DP condition, which restricts the distinguishability of the states and thus limits the amount of information that can be extracted about $\theta$.
-The estimation performance is fundamentally limited by the quantum Fisher information, which quantifies the sensitivity of the state $\rho_{\theta}$ to changes in $\theta$.
-In the quantum setting, the relevant quantity is the Symmetric Logarithmic Derivative (SLD) Fisher information, defined as
-$J_{\theta}:=\mathrm{Tr}\left[L_{\theta}^{2}\rho_{\theta}\right],$
-where $L_{\theta}$ is the SLD operator satisfying
-$\frac{1}{2}(L_{\theta}\rho_{\theta}+\rho_{\theta}L_{\theta})=\frac{d\rho_{\theta}}{d\theta}.$
-The quantum Cramér-Rao bound states that the mean squared error (MSE) of any unbiased estimator $\hat{\theta}$ is lower bounded by $1/J_{\theta}$.
-In the privatized scenario, the investigator is restricted to the pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$, and thus the achievable Fisher information is maximized over all such admissible pairs.
-This leads to a constrained optimization problem,
-$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}J_{\theta},$
-where the maximum is taken over all pairs of quantum states satisfying the privacy constraint.
-The optimal value quantifies the fundamental tradeoff between privacy and estimation accuracy in quantum settings.
-Furthermore, the structure of the optimal $(\varepsilon,\delta)$-DP pair, as characterized in Lemma 4, allows for explicit computation of the Fisher information and the corresponding Cramér-Rao bound.
-This provides a precise benchmark for the best possible estimation performance under quantum differential privacy constraints, and generalizes classical results to the quantum regime.
+(i) Fact 6 yields that $D(\cdot\|\cdot)=D_{f}(\cdot\|\cdot)$, where $f(x)=x\log x$. Further, the pair $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ satisfies the following,
+$D(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)})$ $=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\varepsilon-\frac{1}{1+e^{\varepsilon}}\varepsilon$
+$=\varepsilon\left(\frac{e^{\frac{\varepsilon}{2}}-e^{-\frac{\varepsilon}{2}}}{e^{\frac{\varepsilon}{2}}+e^{-\frac{\varepsilon}{2}}}\right)=\varepsilon\tanh\left(\frac{\varepsilon}{2}\right).$ (75)
+Therefore, (71) of Corollary 2 completes the proof of (i) of Corollary 3.
+(ii) The pair $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ satisfies the following,
+${\left\lVert\rho_{(\varepsilon)}-\sigma_{(\varepsilon)}\right\rVert}_{1}$ $=\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|+\left|\frac{1-e^{\varepsilon}}{1+e^{\varepsilon}}\right|$
+$\overset{(a)}{=}2\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|=2\tanh\left(\frac{\varepsilon}{2}\right).$ (76)
+where $(a)$ follows from the fact that $e^{\varepsilon}\geq 1$ as $\varepsilon\geq 0$.
+Note that we have ${\left\lVert\rho-\sigma\right\rVert}_{1}=2\left|\rho-\sigma\right|_{+}=2E_{1}(\rho\|\sigma)$, where the last equality follows from Fact 10.
+Therefore, the relation (72) of Corollary 2 derives that ${\left\lVert\rho-\sigma\right\rVert}_{1}=2E_{1}(\rho\|\sigma)\leq 2E_{1}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)})={\left\lVert\rho_{(\varepsilon)}-\sigma_{(\varepsilon)}\right\rVert}_{1}$.
+This completes the proof of (ii) of Corollary 3.
+V Quantum Privatized Parameter Inference
+V-A Problem setting
+We consider a statistical inference problem with $n$ respondents with the privacy setting.
+Respondent $i\in\{1,\ldots,n\}$ holds a private bit $X_{i}\in\{0,1\}$, and the investigator aims to infer a population parameter $\theta\in[0,1]$ that governs the relative frequency of the two outcomes.
+Concretely, since (from the investigator’s perspective) the respondents’ behavior is modeled as random sampling with replacement, the $n$ private bits can be treated as i.i.d.
+draws from a Bernoulli law $p_{\theta}$ on $\{0,1\}$ with $p_{\theta}(0)=\theta$ and $p_{\theta}(1)=1-\theta$.
+Equivalently, the investigator treats the collected answers as random sampling with replacement from a population with composition parameterized by $\theta$, so that $(X_{1},\ldots,X_{n})$ is (approximately) i.i.d.
+under $p_{\theta}$.
+The key task is to estimate $\theta$ from the $n$ (privatized) responses.
+To gather information while protecting privacy, each respondent releases not the raw bit $X_{i}$ but a quantum system whose state depends on $X_{i}$.
+Specifically, we fix two quantum states $\rho$ and $\sigma$, and respondent $i$ releases state $\rho$ if $X_{i}=0$ and state $\sigma$ if $X_{i}=1$.
+Thus, each response is produced by an encoding map $x\in\{0,1\}\mapsto\rho_{x}$, where $\rho_{0}:=\rho$ and $\rho_{1}:=\sigma$.
+The privacy requirement is imposed at the level of each respondent’s released system.
+Namely, we require that the pair of possible outputs of an individual respondent, $(\rho,\sigma)$, satisfies $(\varepsilon,\delta)$-DP in the sense of Definition 1, i.e., $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$.
+Parametrized states induced by the population model.
+Under the Bernoulli model with parameter $\theta$, a single privatized response is described by the mixture $\rho_{\theta}:=\theta\,\rho+(1-\theta)\,\sigma$ with $\theta\in[0,1]$.
+Since the investigator collects $n$ responses, the resulting joint state is modeled as the product state $\rho_{\theta}^{\otimes n}$.
+This is the quantum analogue of observing $n$ i.i.d.
+privatized samples drawn according to $p_{\theta}$, where the unknown parameter $\theta$ controls the mixture proportions between $\rho$ and $\sigma$.
+We also define $\rho_{(\varepsilon,\delta),\theta}$ analogously when the respondents encode their information using the weakest/most informative $(\varepsilon,\delta)$-DP pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ constructed earlier in the paper.
+That is, $\rho_{(\varepsilon,\delta),\theta}:=\theta\,\rho_{(\varepsilon,\delta)}+(1-\theta)\,\sigma_{(\varepsilon,\delta)}$.
+This reference pair plays a canonical role in our analysis: since it is the most informative among all $(\varepsilon,\delta)$-DP pairs, it provides the fundamental benchmark for the maximum inferential power (e.g., hypothesis testing performance and Fisher information) achievable by any privatized mechanism under the same privacy parameters.
+More generally, quantum parameter inference concerns estimating an unknown parameter $\theta$ encoded in a family of states $\{\rho_{\theta}\}$, where the encoding is constrained by privacy requirements.
+In the present privatized setting, the requirement $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ restricts the distinguishability of the two single-shot outputs under any measurement and hence limits how much information about $\theta$ can be extracted from $\rho_{\theta}^{\otimes n}$.
+In the following subsections, we quantify this limitation through (i) privatized hypothesis testing and (ii) bounds on the (SLD) Fisher information under $(\varepsilon,\delta)$-QDP.
+V-B Hypothesis testing with one respondent
+First, for the theoretical and fundamental analysis, we consider the scenario where the investigator aims to distinguish between two original information sources, $\rho_{\theta_{0}}$ and $\rho_{\theta_{1}}$ with $\theta_{0}<\theta_{1}$, only from one observation; however, this scenario is impractical as it addresses only a single respondent.
+The performance in this hypothesis testing task is characterized by quantities like the hypothesis testing divergence $D_{H}^{\alpha}(\rho_{\theta_{0}}\|\rho_{\theta_{1}})$.
+Theorem 3.
+The maximum hypothesis testing divergence under the $(\varepsilon,\delta)$-DP constraint is given by:
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}D_{H}^{\alpha}(\rho_{\theta_{1}}\|\rho_{\theta_{2}})=D_{H}^{\alpha}(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}).$ (77)
+This maximum is uniformly attained by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
+Proof.
+It is sufficient to show
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})=\beta^{c}(\alpha|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}).$ (78)
+where $\beta^{c}(\alpha|\sigma_{1}\|\sigma_{2}):=1-\beta(\alpha|\sigma_{1}\|\sigma_{2})$. The relation (61) implies
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}\beta^{c}(\alpha|\rho\|\sigma)=\beta^{c}(\alpha|\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}).$ (79)
+Hence, the following discussion will show (78) from (79).
+Assume that the states $\rho$ and $\sigma$ satisfy the $(\varepsilon,\delta)$-DP condition, which implies
+$\beta^{c}(\alpha|\rho\|\sigma)\leq\beta^{c}(\alpha|\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}).$ (80)
+The quantity $\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ is written as
+$\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})=$ $\max_{\begin{subarray}{c}\alpha^{\prime}\in[0,1]:\\ \theta_{1}\alpha^{\prime}+(1-\theta_{1})\beta^{c}(\alpha^{\prime}|\rho\|\sigma)\leq\alpha\end{subarray}}\theta_{0}\alpha^{\prime}+(1-\theta_{0})\beta^{c}(\alpha^{\prime}|\rho\|\sigma).$ (81)
+We choose a real number $\alpha_{0}$ satisfying $\theta_{1}\alpha_{0}+(1-\theta_{1})\beta^{c}(\alpha_{0}|\rho\|\sigma)=\alpha$.
+Then, $\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})=\theta_{0}\alpha_{0}+(1-\theta_{0})\beta^{c}(\alpha_{0}|\rho\|\sigma)$ because $\theta_{0}\alpha^{\prime}+(1-\theta_{0})\beta^{c}(\alpha^{\prime}|\rho\|\sigma)$ and $\theta_{1}\alpha^{\prime}+(1-\theta_{1})\beta^{c}(\alpha^{\prime}|\rho\|\sigma)$ are strictly monotonically increasing for $\alpha^{\prime}$.
+Also, we choose a real number $\alpha_{1}$ satisfying $\theta_{1}\alpha_{1}+(1-\theta_{1})\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=\alpha$.
+Hence, the relation (80) implies that $\alpha_{1}\leq\alpha_{0}$.
+Hence,
+$\beta^{c}(\alpha|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-\beta^{c}(\alpha|\rho_{\theta_{1}}\|\rho_{\theta_{0}})$
+$=$ $\theta_{0}\alpha_{1}+(1-\theta_{0})\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-(\theta_{0}\alpha_{0}+(1-\theta_{0})\beta^{c}(\alpha_{0}|\rho\|\sigma))$
+$=$ $(\theta_{1}-\theta_{0})(\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-\alpha_{1})-(\theta_{1}-\theta_{0})(\beta^{c}(\alpha_{0}|\rho\|\sigma)-\alpha_{0})$
+$=$ $(\theta_{1}-\theta_{0})(\beta^{c}(\alpha_{1}|\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}})-\beta^{c}(\alpha_{0}|\rho\|\sigma)+\alpha_{0}-\alpha_{1})\geq 0,$
+which implies (78). This completes the proof of Theorem 3.
+The combination of Theorems 1 and 3 derives the following corollary.
+Corollary 4.
+For any $\alpha\in[0,1)\cup(1,+\infty)$, we have
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}D_{\alpha}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=D_{\alpha}(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}),$ (82)
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}D(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=D(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}),$ (83)
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}\check{D}_{\alpha}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=\check{D}_{\alpha}(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}),$ (84)
+This maximum is uniformly attained by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
+V-C Hypothesis testing with $n$ respondents
+Next, we discuss the hypothesis testing when the investigator receives $n$ responses.
+This scenario focuses on the hypothesis testing divergence $D_{H}^{\alpha}(\rho_{\theta_{1}}^{\otimes n}\|\rho_{\theta_{0}}^{\otimes n})$.
+We study the following quantity
+$B_{\epsilon}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $:=\lim_{n\to\infty}\frac{1}{n}D_{H}^{\epsilon}(\rho_{\theta_{1}}^{\otimes n}\|\rho_{\theta_{0}}^{\otimes n}),$ (85)
+for $\epsilon\in(0,1)$.
+Since it is known that the above values are given by $D(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ as Stein’s lemma [17, 27], [16, Chapter 3], thus, (83) of Corollary 4 yields the following relation.
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}B_{\epsilon}(\rho_{\theta_{1}}\|\rho_{\theta_{0}})$ $=D(\rho_{(\varepsilon,\delta),\theta_{1}}\|\rho_{(\varepsilon,\delta),\theta_{0}}).$ (86)
+The above relations show that the asymptotic behavior of the hypothesis testing problem is optimized when respondents encode their information by using $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$.
+V-D Parameter estimation
+Next, we discuss the estimation of the parameter $\theta$ when the investigator has $n$ data.
 The investigator’s goal is to estimate the parameter $\theta$.
@@
 $J_{\theta}:=\mathrm{Tr}(L_{\theta}^{2}\rho_{\theta}).$ (88)
-The MSE of any unbiased estimator for $\theta$ is lower-bounded by $1/J_{\theta}$.
-This bound is asymptotically achievable, for instance, via a two-step estimation process.
+The MSE of any unbiased estimator for $\theta$ is lower-bounded by $1/J_{\theta}$. This bound is asymptotically achievable, for instance, via a two-step estimation process.
+In the privatized scenario, the investigator is restricted to the pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$.
 Consequently, to optimize the estimation, it is natural to maximize the SLD Fisher information $J_{\theta}$ subject to the $(\varepsilon,\delta)$-DP privacy constraint on the states $\rho$ and $\sigma$.
+This leads to a constrained optimization problem,
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}J_{\theta},$
+where the maximum is taken over all pairs of quantum states satisfying the privacy constraint.
+The optimal value quantifies the fundamental tradeoff between privacy and estimation accuracy in quantum settings.
+Furthermore, the structure of the optimal $(\varepsilon,\delta)$-DP pair, as characterized in Lemma 3, allows for explicit computation of the Fisher information and the corresponding Cramér-Rao bound.
+This provides a precise benchmark for the best possible estimation performance under quantum differential privacy constraints, and generalizes classical results to the quantum regime.
 Theorem 4.
 The maximum SLD Fisher information achievable under the $(\varepsilon,\delta)$-DP constraint is given by:
-$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}J_{\theta}=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$ (68)
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}J_{\theta}=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$ (89)
 This maximum is uniformly attained by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
@@
 Let $\rho_{(\varepsilon,\delta),\theta}:=\theta\rho_{(\varepsilon,\delta)}+(1-\theta)\sigma_{(\varepsilon,\delta)}$.
-We first compute the SLD Fisher information for this specific state, which we denote by $J_{(\varepsilon,\delta),\theta}$.
+First, we denote the SLD Fisher information for this specific states $\rho_{(\varepsilon,\delta),\theta}$ by $J_{(\varepsilon,\delta),\theta}$, and we will show the relation
+$\max_{(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}}J_{\theta}=J_{(\varepsilon,\delta),\theta}.$ (90)
+It is known that [16, Theorem 6.3]
+$J_{\theta}=\lim_{\epsilon\to 0}\frac{8(1-\mathrm{Tr}[|\sqrt{\rho_{\theta}}\sqrt{\rho_{\theta+\epsilon}}|])}{\epsilon^{2}}.$ (91)
+Since $\check{D}_{\frac{1}{2}}(\rho\|\sigma)=-\frac{1}{2}\log\mathrm{Tr}[|\sqrt{\rho}\sqrt{\sigma}|]$, the above relation is rewritten as
+$J_{\theta}=\lim_{\epsilon\to 0}\frac{4\check{D}_{\frac{1}{2}}(\rho_{\theta}\|\rho_{\theta+\epsilon})}{\epsilon^{2}}.$ (92)
+Therefore, the relation (84) with the limit like (92) implies (90).
+Next, we compute the SLD Fisher information $J_{(\varepsilon,\delta),\theta}$.
 The state $\rho_{(\varepsilon,\delta),\theta}$ is diagonal in the standard basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$.
@@
 $=\frac{\delta}{\theta(1-\theta)}+(1-\delta)(e^{\varepsilon}-1)^{2}\left(\frac{1}{e^{\varepsilon}(\theta^{2}+(1-\theta)^{2})+\theta(1-\theta)(1+(e^{\varepsilon})^{2})}\right)$
-$=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$
-Now, consider any pair of states $(\rho,\sigma)$ that satisfies the $(\varepsilon,\delta)$-DP condition.
-From Corollary 1, there exists a CP-TP map $\Gamma$ such that $\Gamma(\rho_{(\varepsilon,\delta)})=\rho$ and $\Gamma(\sigma_{(\varepsilon,\delta)})=\sigma$.
-From Fact 15, the SLD Fisher information preserves monotonicity under CP-TP maps.
-Therefore, since $J_{\theta}$ is the SLD Fisher information for the state $\rho_{\theta}=\theta\rho+(1-\theta)\sigma$, Corollary 2 implies
-$J_{\theta}\leq J_{(\varepsilon,\delta),\theta}.$ (70)
-Since the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ itself satisfies the $(\varepsilon,\delta)$-DP condition, it is a valid choice for the maximization.
-Thus, the equality in (70) is achievable, and the maximum value is $J_{(\varepsilon,\delta),\theta}$.
-As an alternative application, consider the scenario where the investigator aims to distinguish between two original information sources, $\rho_{\theta_{0}}$ and $\rho_{\theta_{1}}$.
-The performance in this hypothesis testing task is characterized by quantities like the hypothesis testing divergence $D_{H}^{\alpha}(\rho_{\theta_{0}}^{\otimes n}\|\rho_{\theta_{1}}^{\otimes n})$.
-Also, Corollary 2 implies the following theorem.
-Theorem 4.
-The maximum hypothesis testing divergence under the $(\varepsilon,\delta)$-DP constraint is given by:
-$\max_{(\rho,\sigma):(\varepsilon,\delta)\text{-DP}}D_{H}^{\alpha}(\rho_{\theta_{0}}^{\otimes n}\|\rho_{\theta_{1}}^{\otimes n})=D_{H}^{\alpha}(\rho_{(\varepsilon,\delta),\theta_{0}}^{\otimes n}\|\rho_{(\varepsilon,\delta),\theta_{1}}^{\otimes n}),$ (71)
-This maximum is uniformly attained by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$.
-The asymptotic behavior of the hypothesis testing problem for the optimal states is characterized by the quantum Rényi divergences between $\rho_{(\varepsilon,\delta),\theta_{0}}$ and $\rho_{(\varepsilon,\delta),\theta_{1}}$.
-The specific expressions for $D(\rho_{(\varepsilon,\delta),\theta_{0}}\|\rho_{(\varepsilon,\delta),\theta_{1}})$ and $D_{\alpha}(\rho_{(\varepsilon,\delta),\theta_{0}}\|\rho_{(\varepsilon,\delta),\theta_{1}})$ can be computed from their definitions.
-VII Stability of Learning Algorithms
-In the previous setting, the investigator receives information directly from the respondents.
-However, when there exists a trusted third party, the data processor, who applies the learning algorithm to the training data sent by the respondents between the respondents and the investigator, it is sufficient to discuss the output of the learning algorithm for the privacy.
-Under this assumption, we can relax our condition for privacy.
-To discuss this case, we use the framework developed earlier to study a fundamental problem in learning theory: the privacy of training data.
-Specifically, we ask how much information the output of a learning algorithm reveals about its training data.
-Clearly, if the output were completely independent of the training data, it would reveal no information.
-However, such an algorithm would be useless for learning because it could not leverage the data to make predictions.
-Thus, there is an inherent trade-off between preserving privacy and maintaining the utility of the learning algorithm.
-To address this trade-off, a private learning algorithm should ensure that its output does not change significantly when a single training example is modified.
-This property is known as stability, which was first introduced in [1].
-Formally, stability refers to the insensitivity of the algorithm’s output to small changes in the training dataset, such as replacing one data point with a neighboring one.
-The authors in [1] studied this in terms of $L_{1}$ distance.
-In the classical setting, this notion is well established, and differentially private learning algorithms are known to provide strong stability guarantees [33].
-In [6], the authors studied the stability of $\varepsilon$-DP learning algorithms and obtained an upper-bound on the mutual information between the training data at the respondent’s end and the algorithm’s output at the investigator’s end.
-The main aim of the subsequent subsections is to generalize the [6, Proposition 2] in the quantum setting for $\delta\neq 0$.
-Towards this, we first discuss a differentially private quantum learning framework in the subsection below.
-VII-A Framework for $1$-neighbor $(\varepsilon,\delta)$-DP quantum learning algorithms
-Motivated by the learning frameworks of [14] and [15], we will assume that the input to the quantum learning algorithm is a classical $n$-length ($n\in\mathbb{N}$) data $s:=(z_{1},\cdots,z_{n})\in\mathcal{S}:=\mathcal{Z}^{n}$ (where $\left|\mathcal{Z}\right|>0$) and a $\rho_{s}\in\mathcal{D}((\mathbb{C}^{d})^{\otimes n})$.
-For each $i\in[n]$, $z_{i}:=(x_{i},y_{i})$, where $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ for some non-empty sets $\mathcal{X}$ and $\mathcal{Y}$ and $y_{i}=f(x_{i})$ for some concept function $f:\mathcal{X}\to\mathcal{Y}$, which is unknown to the the learning algorithm.
-In a typical case, each data $z_{i}$ for $i\in[n]$ has the form $z_{i}:=(x_{i},y_{i})$, where $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ for some non-empty sets $\mathcal{X}$ and $\mathcal{Y}$ and $y_{i}=f(x_{i})$ for some concept function $f:\mathcal{X}\to\mathcal{Y}$, which is unknown to the the data processor.
-In this quantum learning framework, the algorithm’s objective is to learn a concept function $f$ from a classical sequence $s$ and a quantum state $\rho_{s}$.
-Thus, the manuscripts [14] and [15] formulated the input to a learning algorithm to be a classical-quantum state $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$.
-The quantum mechanics formalism restricts the algorithm to only use measurements and CP-TP maps.
-Therefore, the joint state between the training data $S$ and the output of any learning algorithm in this framework can be modeled by a classical-quantum state,
-$\sigma^{SB}:=\sum_{s\in\mathcal{S}}P_{S}(s)|s\rangle\langle s|\otimes\sigma^{B}_{s},$ (72)
-where,
-$\sigma^{B}_{s}:=\sum_{w\in\mathcal{W}}(\mathcal{N}_{s,w}(\rho_{s}))^{B^{\prime}}\otimes|w\rangle\langle w|^{W},$ (73)
-and $w$ is a hypothesis in the hypothesis class $\mathcal{W}$.
-Further, the map $\mathcal{N}_{s}:\rho_{s}\to\sum_{w\in\mathcal{W}}\mathcal{N}_{s,w}(\rho_{s})\otimes|w\rangle\langle w|$ is a quantum instrument, i.e., for every $(s,w)\in\mathcal{S}\times\mathcal{W},$ $\mathcal{N}_{s,w}$ is a completely positive trace non-increasing map.
-In the discussions below for every $s\in\mathcal{S},$ we will use following the notation
-$\mathcal{N}_{s}(\rho_{s}):=\sum_{w\in\mathcal{W}}\mathcal{N}_{s,w}(\rho_{s})\otimes|w\rangle\langle w|.$
-From the learning point of view, it is quite natural to expect that the order in which the training data is fed to the algorithm should not change the algorithm’s output.
-Therefore, we will assume that if $s$ and $s^{\prime}$ have the same type as $T_{s},$ then both $\rho_{s}$ and $\rho_{s^{\prime}}$ get mapped to the same state at the output of the learning algorithm.
-Formally, upon getting an input $\{s,\rho_{s}\}$, the algorithm maps, $\rho_{s}\to V_{\pi_{s}}\rho_{s}V^{\dagger}_{\pi_{s}}$, where $\pi_{s}:s\to T_{s}$ permutes $s$ to $T_{s}$ which is the type representative of $s$.
-Thus, for every $s$ in a type $T_{s},$ its associated quantum output is $\sum_{w\in\mathcal{W}}\mathcal{N}_{T_{s},w}(\rho_{T_{s}})\otimes|w\rangle\langle w|.$
-This property is also desirable from the viewpoint of the privacy because this property disables the investigator to identify which respondent has the respective data $z$.
-Therefore, we consider that this property is a part of the conditions for the privacy as follows.
-Definition 20.
-An algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$ is said to be a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm if and only if the following properties are true.
-  1. 1.
-    For every $s\in\mathcal{S}$ the quantum instrument $\mathcal{N}_{s}$ should be 11-neighbor $(\varepsilon,\delta)$-DP, i.e., for every $s\overset{1}{\sim}s^{\prime}$ (see Definition 3), the following holds,
-    $\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]+\delta,$ (74)
-    $\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]$ $\leq e^{\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]+\delta,$ (75)
-    for every $0\preceq\Lambda\preceq\mathbb{I}.$ The property desired in (74) and (75) is to ensure privacy during the learning process.
-  2. 2.
-    It should be support consistent. Formally, for each $s\overset{1}{\sim}s^{\prime}\in\mathcal{S}$, the following holds,
-    $\operatorname{supp}(\mathcal{N}_{s}(\rho_{s}))=\operatorname{supp}(\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})).$ (76)
-    The property mentioned in (76) is restrictive and may not be true for many learning algorithms.
-However, many of the information-theoretic quantities which can be used to analyze the stability require this condition.
-Therefore, in light of this requirement, we make the above assumption.
-Further, for $\delta=0$, this condition is inherently there.
-Otherwise, for no finite $\varepsilon\geq 0$ the pair of quantum states will satisfy the $\varepsilon$-DP criterion.
-A property which follows from the Definition 20 is that if a learning algorithm $\mathcal{A}$ is $1$-neighbor $(\varepsilon,\delta)$-DP, then it is also a differentially private algorithm with respect to data points which are $k$-neighbors.
-However, the privacy parameters will degrade with $k.$ We formalize this in the corollary below.
-Corollary 5.
-If a learning algorithm $\mathcal{A}$ satisfies Definition 20, then, any pair of $k$-neighbor (where $k>1$) inputs $s\overset{k}{\sim}s^{\prime}$ and for every $0\preceq\Lambda\preceq\mathbb{I}$, the following holds,
-$\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]+g_{k}(\delta),$ (77)
-$\mathrm{Tr}[\Lambda\mathcal{N}_{s^{\prime}}(\rho_{s^{\prime}})]\leq e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}_{s}(\rho_{s})]+g_{k}(\delta),$ (78)
-where,
-$g_{k}(\varepsilon,\delta):=\frac{e^{k\varepsilon}-1}{e^{\varepsilon}-1}\delta,$ (79)
-and is assumed to be less than $1.$
-Proof.
-See Appendix -F for the proof.
-Remark 3.
-It is important to note that in this section, our focus is not on the training process of the learning algorithm itself.
-Instead, we are primarily concerned with the privacy preservation between the input training data and the output of the quantum learning algorithm.
-We analyze how the $(\varepsilon,\delta)$-DP property of the algorithm limits the information an investigator can infer about the training data from the algorithm’s output.
-VII-B Stability of a $1$-neighbor $(\varepsilon,\delta)$-DP support consistent quantum learning algorithm
-From the point of view of privacy-preserving learning, the output of a learning algorithm should reveal minimal information about the training dataset it was trained on.
-In this scenario, the trusted party, the data processor, has access to the learning algorithm and the input state $\sum_{s}P_{S}(s)|s\rangle\langle s|\otimes\rho_{s}$, and the investigator, who has access to the output subsystem $B$ and may be interested in learning about $S$ from $B$.
-The investigator attempts to infer as much information as possible about the respondent’s original training data $S$ from the algorithm’s output $B$.
-In contrast, the data processor’s goal is to ensure that the learning process discloses little to no information about the input dataset.
-Figure 3 below illustrates this framework.
-[…]
-Figure 3: Privacy based learning framework.
-In the context of the above adversarial setting between the respondents and the investigator, motivated from [34], we focus on the mutual information $I\left[S;B\right]$, which is calculated with respect to $\sum_{s\in\mathcal{S}}P_{S}(s)|s\rangle\langle s|\otimes{\mathcal{N}_{s}(\rho_{s})}^{B}$.
-This quantity satisfies the following chain rule.
-$I[S;B]=I[S;B^{\prime}W]=I[S;B^{\prime}\mid W]+I[S;W].$ (80)
-The respondents impose the requirement that the output should remain essentially unchanged under the modification of a single training example.
-This property, called stability, is formalized in the following definition.
-Definition 21.
-(Stability) A quantum learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s}$ is $\gamma$-stable, if
-$\max_{Q_{S}}I\left[S;B\right]\leq\gamma.$
-The above definition provides a quantitative upper-bound on the maximum classical information that an adversary or investigator can extract from $B$ about $S$.
-Consequently, a small upper-bound implies that the algorithm’s output is not strongly dependent on any single training data point, indicating that the algorithm is information-theoretically stable.
-Therefore, the respondent will always aim to minimize $\gamma$ by using algorithms that are not sensitive to small changes in the input training sequence $s$.
-Towards this, we have the following definition.
-Now, using the framework presented in Section VII-A, we analyze the stability of the quantum private learning algorithm by deriving an upper-bound on the Holevo information in the theorem below, under the assumption that
-$g_{n(\left|\mathcal{Z}\right|-1)}(\varepsilon,\delta)=\frac{e^{n(\left|\mathcal{Z}\right|-1)\varepsilon}-1}{e^{\varepsilon}-1}\delta<1.$ (81)
-We now have the following theorem.
-Theorem 5.
-For $\varepsilon\in\left[\frac{1}{n},1)\right$., consider a learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 20 and (81).
-Then, the following holds,
-$I[S;B]_{\sigma}\leq(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (82)
-where, $n$ is the length of the training data and for some constant $m\in(0,1],$ $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\log\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ and has a property that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$.
-Proof.
-See Appendix -D for the proof.
-[…]
-Corollary 6.
-For $\varepsilon\in\left.[0,\frac{1}{n}\right)$, consider a learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 20 and (81).
-Then, the following holds,
-$I[S;B]_{\sigma}\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$ (83)
-Corollary 7.
-[…]
-$I[S;B]_{\sigma}\leq(|\mathcal{Z}|-1)\log\left(n+1\right).$ (84)
-The proof of Theorem 5 above relies on the lemma below, which can be thought of as a quantum analog of [6, Lemma $2$].
-Lemma 5.
-Consider $\rho$ and $\sigma$ be two quantum states over Hilbert space $\mathcal{H}_{A}$ such that $\rho\ll\sigma$ and $\sigma$ is a finite mixture of probability distributions such that $\sigma=\sum_{b=1}^{m}P(b)\sigma_{b}$, where $\sum_{b=1}^{m}P(b)=1$, and $\rho\ll\sigma_{b}$ for all $b\in[m]$.
-Then, the following holds,
-$D(\rho\|\sigma)\leq\min_{b\in[m]}\left\{D\left(\rho\|\sigma_{b}\right)-\log P(b)\right\}.$ (85)
-Further, a tighter bound in comparison to (85) is as follows,
-$D(\rho\|\sigma)\leq-\log\left(\sum_{b=1}^{m}P(b)\exp\left(-D\left(\rho\|\sigma_{b}\right)\right)\right).$
-Proof.
-See Appendix -E for the proof.
-Theorem 5 establishes a crucial quantitative link between differential privacy and stability by providing an explicit upper-bound on the Holevo information $I[S;B]_{\sigma}$ between the training dataset $S$ and the algorithm output $B$.
-Furthermore, the upper-bound given in Theorem 5 is uniform and explicitly depends on dataset size $n$, alphabet size $|\mathcal{Z}|$, and privacy parameters $(\varepsilon,\delta)$.
-This makes clear how stability scales with these variables.
-In particular, the theorem translates the $(\varepsilon,\delta)$-privacy guarantees into a provable stability bound, thereby linking differential privacy directly to stability-based generalization controls measured by mutual information.
-This establishes a precise and quantitative connection between privacy and algorithmic stability.
-Theorem 5 allows us to obtain a classical version which can be considered as $(\varepsilon,\delta)$-DP generalization of Proposition 2.
-Before discussing this result, we first briefly discuss a classical learning framework, which is $1$-neighbor $(\varepsilon,\delta)$-DP.
-Remark 4.
-The upper-bound obtained in Theorem 5 is independent of $P_{S}$ and thus, Theorem 5 implies that if a quantum learning algorithm $\mathcal{A}=\left\{\mathcal{N}_{s}\right\}_{s\in\mathcal{S}}$ satisfies Definition 20, then $\mathcal{A}$ is $\left((|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+\right$.
-$\left.h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right)$-stable (see Definition 21).
-A similar observation also follows for Corollaries 6 and 7.
-VII-C $1$-neighbor $(\varepsilon,\delta)$-DP support consistent classical learning algorithms and its stability bound
-In the classical setting a learning algorithm takes input $s\in\mathcal{S}$ and produces a $w\in\mathcal{W}$ according to some condition distribution $p_{W\mid s}.$ For more details on $s$ and $w$ see subsection VII-A Further, in the context of learning, it is reasonable to assume that the condition distribution is independent of the order in which the training samples are fed to it (i.e., it depends only on the type of $s$).
-We say that a learning algorithm is $1$-neighbor $(\varepsilon,\delta)$-DP support consistent learning algorithm if for every $s\overset{1}{\sim}s^{\prime}$ (see Definition 3), we have
-$p_{W\mid s}(\mathcal{J})$ $\leq e^{\varepsilon}p_{W\mid s^{\prime}}(\mathcal{J})+\delta,$
-$p_{W\mid s^{\prime}}(\mathcal{J})$ $\leq e^{\varepsilon}p_{W\mid s}(\mathcal{J})+\delta,$ (86)
-for $\forall\mathcal{J}\subseteq\mathcal{W}.$
-When the learning algorithm $\mathcal{A}=\left\{p_{W\mid s}\right\}_{s\in\mathcal{S}}$ satisfies the above condition for every $s\overset{1}{\sim}s^{\prime},$ $\operatorname{supp}(p_{W\mid s})=\operatorname{supp}(p_{W\mid s^{\prime}})$, Theorem 5 implies the following,
-Corollary 8.
-Any classical learning algorithm $\mathcal{A}=\left\{p_{W\mid s}\right\}_{s\in\mathcal{S}}$ which satisfies the properties mentioned in (86), satisfies the following upper-bounds,
-  1. i)
-    for $\varepsilon>1$,
-    $I[S;W]\leq(|\mathcal{Z}|-1)\log\left(n+1\right).$
-  2. ii)
-    for $\frac{1}{n}\leq\varepsilon\leq 1$,
-    $I[S;W]\leq(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
-  3. iii)
-    for $\varepsilon<\frac{1}{n}$,
-    $I[S;W]\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
-Further, for $\delta=0$, Corollary 8 above, recovers [6, Proposition 22] mentioned below.
-Proposition 2 ([6, Proposition 22]).
-For $\varepsilon<1$ and $\delta=0$ any classical learning algorithm $\mathcal{A}=\left\{p_{W\mid s}\right\}_{s\in\mathcal{S}}$ which satisfies the properties mentioned in (86), satisfies the following upper-bound,
-$I[S;W]\leq(|\mathcal{Z}|-1)\log\big(1+\varepsilon en\big),$
-where $n$ is the length of the training data.
-VII-D Comparison between Theorem 5 and [20, Proposition 10]
-In [20, Proposition 10], the authors claimed to have derived a stability upper-bound on Holevo information for quantum $(\varepsilon,\delta)$-LDP quantum channels.
-However, it is doubtful and misleading for two reasons.
-The first reason is that the applicability of the upper-bound in [20, Proposition 10] to $(\varepsilon,\delta)$-LDP quantum channels is questionable.
-Although the proposition claims that the bound holds for general $(\varepsilon,\delta)$-LDP channels, the derived inequality is valid only in the special case $\delta=0$.
-That is, it is evident that the bound can be violated when $\delta$ becomes sufficiently large.
-The second and more important reason is that the stability of a learning algorithm is closely related to how sensitive the output of a learning algorithm is with respect to minor changes in the input training data.
-Even though the authors in [20, Proposition 10] claim to study the stability, the results obtained by them nowhere capture the relation between the stability and the sensitivity.
-Formally, they have not considered the algorithms which are $(\varepsilon,\delta)$ differentially private quantum channel only for neighboring (defined appropriately) input quantum states.
-In contrast, Theorem 5 takes care of all the issues discussed above.
-VII-E Comparison between Theorem 5 and [14, Appendix C.7]
-In [14, Appendix C.7], the authors obtain an upper-bound on the Holevo information assuming their learning framework under pure $\varepsilon$-LDP constraints.
-However, they don’t study this from the point of view of the stability of a quantum learning algorithm.
-The bound on the Holevo information that they obtain (under pure $\varepsilon$-LDP constraints) is conceptually wrong. This is because of the reasons mentioned below.
-  1. 1.
-    In their context of the learning algorithm, the $\varepsilon$-LDP channels (or measurements) act only on the training part of the data available.
-In particular, to obtain [14, Eq.
-(C.7.1)], the authors assume that $\Lambda^{\mathcal{A}}_{s,w}$, for any $0\preceq M\preceq\mathbb{I}^{\mbox{hyp}}$ and $\rho^{\mbox{train}}_{1},\rho^{\mbox{train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mbox{train}})$, $\Lambda^{\mathcal{A}}_{s,w}$ satisifies the following,
-    $\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{1})\right]\leq e^{\varepsilon}\mathrm{Tr}\left[M\Lambda^{\mathcal{A}}_{s,w}(\rho^{\mbox{train}}_{2})\right].$
-    The authors further claim that even this local action would maintain the $\varepsilon$-LDP globally. However, for [14, Eq. (C.7.1)] to be true, the following must hold,
-    $\mathrm{Tr}\left[O\left(\mathbb{I}^{\mbox{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\right)(\rho^{\mbox{test;train}}_{1})\right]\leq e^{\varepsilon}\mathrm{Tr}\left[O\left(\mathbb{I}^{\mbox{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}\right)(\rho^{\mbox{test;train}}_{2})\right],$ (90)
-    for every $0\preceq O\preceq\mathbb{I}^{\mbox{test;hyp}}$.
-From [4, Theorem $4$], (90) does not hold because $\mathbb{I}^{\mbox{test}}$ is not a differentially private channel for some fixed $\varepsilon\geq 0$.
-    Therefore, the validity of [14, Eq. (C.7.1)] is questionable under the given assumptions and may require further scrutiny.
-  2. 2.
-    Further, in their framework, they consider the training part of their learning algorithm to consist of certain POVM measurements (which depend on the classical training data) that act on the quantum part of the training data.
-However, while studying the Holevo information under $\varepsilon$-LDP constraints, they assume that the POVMs are independent of the training data.
-In particular, on the page $59$ of [14] the authors mention the following,
-    “Next, we turn our attention to the classical MI term in our generalization bounds.
-Here, we assume that the data processor $\mathcal{A}$ uses an overall $\varepsilon$-LDP POVM.
-As the POVM $\left\{|s\rangle\langle s|\otimes E_{s}^{\mathcal{A}}(w)\right\}_{s,w}$ is not LDP even if every $\left\{E_{s}^{\mathcal{A}}(w)\right\}_{w}$ is, we make the simplifying assumption that the data processor uses an $s$-independent $\varepsilon$-LDP POVM $\left\{E^{\mathcal{A}}(w)\right\}_{w}$.”
-    The above over-simplified assumption is as good as saying that the learning algorithm doesn’t depend on the training data.
-In that case, the algorithm is trivially private.
-However, it will incur severe errors.
-Theorem 5 does not have any such issues.
-VIII Upper-Bounds on Relative Entropy between the outputs of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
-VIII-A Formulation of $(\varepsilon,\delta)$-LDP classical channels
-[…]
+$=\frac{\delta}{\theta(1-\theta)}+\frac{(1-\delta)(1-e^{\varepsilon})^{2}}{e^{\varepsilon}+(1-e^{\varepsilon})^{2}\theta(1-\theta)}.$ (94)
+The combination of (90) and (94) yields the desired relation (89).
+VI Upper-Bounds on Relative Entropy between the outputs of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
+VI-A Formulation of $(\varepsilon,\delta)$-LDP classical channels
+In this section, we show some other applications of the dominance of quantum informativeness as discussed in Section IV.
+A primary question which is often studied in the context of $(\varepsilon,\delta)$-DP mechanisms is to determine upper-bounds on the divergence with respect to the output induced by the mechanisms.
+To do this, we can easily invoke Theorem 2 and easily get an upper-bound on the divergence between the states induced at the output of the mechanism in terms of $D_{f}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$, where $D_{f}$ is some $f$-divergence as defined in Definition 10.
+However, note that $\operatorname{supp}(\rho_{(\varepsilon,\delta)})\nsubseteq\operatorname{supp}(\sigma_{(\varepsilon,\delta)})$ and therefore there will be many $f$-divergences for which $D_{f}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$ is not well defined even for the case when the divergence $D_{f}(\cdot\|\cdot)$ between the output distributions induced by the $(\varepsilon,\delta)$-DP mechanism is well defined.
+Hence, naively using Theorem 2 will not lead to any meaningful upper-bound on $D_{f}(\cdot\|\cdot)$ between the output distributions induced by an $(\varepsilon,\delta)$-DP mechanism.
 In this section, we will obtain a meaningful bound on the relative entropy $D(\cdot\|\cdot)$ between the output distributions induced by an $(\varepsilon,\delta)$-LDP channel (defined below) for the case when at least the support of one of the output distributions is a subset of the support of the other distribution.
-We will accomplish this with the help of the integral representation of the relative entropy mentioned in Fact 5.
+We will accomplish this with the help of the integral representation of the relative entropy mentioned in Fact 3.
 This integral representation uses the contraction coefficient of $(\varepsilon,\delta)$-LDP channels for the hockey stick divergence.
@@
 Lemma 4.
-A classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP, if and only if any pair of distributions $P_{X},Q_{X}\in\mathcal{P}(\mathcal{X})$ satisfies the relation
+A classical channel $\mathcal{K}$ is $(\varepsilon,\delta)$-LDP, if and only if any pair of distributions $P,Q\in\mathcal{P}(\mathcal{X})$ satisfies the relation
 $\mathcal{K}(P)(S)\leq e^{\varepsilon}\mathcal{K}(Q)(S)+\delta,$ (96)
@@
 In the subsections below, we define the contraction coefficient of classical and quantum $(\varepsilon,\delta)$-LDP channels with respect to general divergences and obtain almost matching upper and lower bounds on the contraction coefficient of the hockey stick divergence for $(\varepsilon,\delta)$-LDP mechanism.
-VIII-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
-For any classical divergence $\mathbb{D}$, from Definition 9, we know that for any pair of probability distributions $P,Q\in\mathcal{P}(\mathcal{X})$ (where $\mathcal{X}$ is some finite set) and a classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ (where $\mathcal{Y}$ is another finite set), $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\mathbb{D}(P_{1}\|P_{2})$, where for each $i=1,2$, $\mathcal{K}(P_{i})$ is the marginal output distribution of $\mathcal{K}$ corresponding to $P_{i}$.
+VI-B Contraction Coefficient for Divergences under Private Classical/Quantum Learning Algorithms
+For any classical divergence $\mathbb{D}$, from Definition 6, we know that for any pair of probability distributions $P_{1},P_{2}\in\mathcal{P}(\mathcal{X})$ (where $\mathcal{X}$ is some finite set) and a classical channel $\mathcal{K}:\mathcal{X}\to\mathcal{Y}$ (where $\mathcal{Y}$ is another finite set), $\mathbb{D}(\mathcal{K}(P_{1})\|\mathcal{K}(P_{2}))\leq\mathbb{D}(P_{1}\|P_{2})$, where for each $i=1,2$, $\mathcal{K}(P_{i})$ is the marginal output distribution of $\mathcal{K}$ corresponding to $P_{i}$.
 However, this inequality is not always strict.
@@
 $\eta^{\mathcal{N}}_{E_{\gamma}}:=\sup_{\begin{subarray}{c}\rho,\sigma\in\mathcal{D}(\mathcal{H}_{A}):\\ E_{\gamma}(\rho\|\sigma)\neq 0\end{subarray}}\frac{E_{\gamma}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))}{E_{\gamma}(\rho,\sigma)}.$ (99)
-Hirche et al. in [32] have studied the following bound on $\eta^{\mathcal{N}}_{E_{\gamma}}$ for any CP-TP some fixed $\gamma\geq 1$.
-Proposition 3 ([32, Lemma II.4]).
+Hirche et al. in [18] studied the following bound on $\eta^{\mathcal{N}}_{E_{\gamma}}$ for any CP-TP some fixed $\gamma\geq 1$.
+Proposition 2 ([18, Lemma II.4]).
 For any $\gamma\geq 1$ and a CP-TP map $\mathcal{N}:\mathcal{D}(\mathcal{H}_{A})\to\mathcal{D}(\mathcal{H}_{B})$, the contraction coefficient $\eta^{\mathcal{N}}_{E_{\gamma}}$ with respect to the quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ (see Definition 9) satisfies the following,
@@
 This is given in the proposition below.
-Proposition 4 ([20, Theorem 5]).
+Proposition 3 ([26, Theorem 5]).
 For any $\varepsilon\geq 0$ and $\delta\in[0,1)$, the contraction coefficient $\eta^{\varepsilon,\delta}_{E_{1}}$ with respect to the trace distance satisfies
 $\eta^{\varepsilon,\delta}_{E_{1}}=\frac{(e^{\varepsilon}-1+2\delta)}{e^{\varepsilon}+1}.$ (102)
-In the following lemma, we now generalize the above result for the contraction coefficient of quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ for any $\gamma\geq 1$ under $(\varepsilon,\delta)$-LDP quantum mechanisms for any $\varepsilon\geq 0$ and $\delta\in[0,1]$.
+In the following lemma, we now generalize the above result for the contraction coefficient of quantum hockey stick divergence $E_{\gamma}(\cdot\|\cdot)$ for any $\gamma\geq 1$ under $(\varepsilon,\delta)$-LDP CP-TP maps for any $\varepsilon\geq 0$ and $\delta\in[0,1]$.
 Lemma 5.
@@
 (1) Lower-bound: Consider the following series of inequalities
-$\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{a}{\geq}1-\gamma(1-\eta^{\varepsilon,\delta}_{E_{1}})$
-$\overset{b}{=}1-2\gamma\frac{(1-\delta)}{e^{\varepsilon}+1}$
+$\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{(a)}{\geq}1-\gamma(1-\eta^{\varepsilon,\delta}_{E_{1}})$
+$\overset{(b)}{=}1-2\gamma\frac{(1-\delta)}{e^{\varepsilon}+1}$
 $=\frac{e^{\varepsilon}+1-2\gamma+2\gamma\delta}{e^{\varepsilon}+1}$
-$\overset{c}{\geq}\frac{e^{\varepsilon}+\delta-\gamma(1-\delta)}{e^{\varepsilon}+1}+\frac{(1-\gamma)}{e^{\varepsilon}+1}$
+$\overset{(c)}{\geq}\frac{e^{\varepsilon}+\delta-\gamma(1-\delta)}{e^{\varepsilon}+1}+\frac{(1-\gamma)}{e^{\varepsilon}+1}$
 $=\eta^{\varepsilon,\delta}_{E_{1}}+\frac{(1-\gamma)(2-\delta)}{e^{\varepsilon}+1},$
-where $a$ follows from Proposition 3, $b$ follows from Proposition 4, $c$ follows because $\gamma\geq 1.$
+where $(a)$ follows from Proposition 2, $(b)$ follows from Proposition 3, $(c)$ follows because $\gamma\geq 1.$
 (2) Upper-bound:
@@
     For the case when $\gamma\in[1,e^{\varepsilon}]$. Consider the following series of inequalities,
-    $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{a}{=}\sup_{\begin{subarray}{c}\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta},\\ |\psi\rangle\perp|\phi\rangle\end{subarray}}E_{\gamma}(\mathcal{N}(|\psi\rangle\langle\psi|)\|\mathcal{N}(|\phi\rangle\langle\phi|))$
-    $\overset{b}{\leq}E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$
-    $\overset{c}{=}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},$ (100)
+    $\eta_{E_{\gamma}}^{\varepsilon,\delta}$ $\overset{(a)}{=}\sup_{\begin{subarray}{c}\mathcal{N}\in\mathfrak{P}_{\varepsilon,\delta},\\ |\psi\rangle\perp|\phi\rangle\end{subarray}}E_{\gamma}(\mathcal{N}(|\psi\rangle\langle\psi|)\|\mathcal{N}(|\phi\rangle\langle\phi|))$
+    $\overset{(b)}{\leq}E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$
+    $\overset{(c)}{=}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},$ (104)
     $=\frac{e^{\varepsilon}-1+2\delta}{e^{\varepsilon}+1}+\frac{(1-\delta)(1-\gamma)}{e^{\varepsilon}+1}$
     $=\eta^{\varepsilon,\delta}_{E_{1}}+\frac{(1-\delta)(1-\gamma)}{e^{\varepsilon}+1},$ (105)
-    where in $a$ $|\psi\rangle,|\phi\rangle\in\mathcal{H}_{A}$ are two orthogonal quantum states and the inequality follows from [32, Theorem II.2], $b$ follows because $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜51 and 53 and $c$ follows from Claim 1 below.
+    where in $(a)$, $|\psi\rangle,|\phi\rangle\in\mathcal{H}_{A}$ are two orthogonal quantum states and the inequality follows from [18, Theorem II.2], $(b)$ follows from Theorem 2 and $(c)$ follows from Claim 1 below.
     Claim 1.
@@
     $E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=\begin{cases}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},&~\mbox{if}~\gamma\in[1,e^{\varepsilon}],\\ \delta,&~\mbox{if}~\gamma>e^{\varepsilon}.\end{cases}$ (106)
-    where $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜51 and 53.
+    where $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ is the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states as mentioned in eqs.˜47 and 49.
     Proof.
-    See Appendix -G for the proof.
+    See Appendix -D for the proof.
   2. (ii)
@@
 $\eta_{E_{\gamma}}^{(c),\varepsilon,\delta}\leq\begin{cases}\frac{e^{\varepsilon}-\gamma+\delta(\gamma+1)}{e^{\varepsilon}+1},&~\mbox{if}~\gamma\in[1,e^{\varepsilon}],\\ \delta,&~\mbox{if}~\gamma>e^{\varepsilon}.\end{cases}$ (107)
-In the quantum setting, Nuradha et al in [20, Proposition 3] obtain an upper-bound on quantum relative entropy under $\varepsilon$-LDP CP-TP maps via an integral representation of quantum relative entropy (Fact 6) in terms of quantum hockey stick divergence.
-Moreover, the upper-bound obtained in [20, Proposition 3] is tighter than the upper-bound obtained in $(i)$ of Corollary 4.
+In the quantum setting, Nuradha et al in [26, Proposition 3] obtained an upper-bound on the quantum relative entropy under $\varepsilon$-LDP CP-TP maps via an integral representation of quantum relative entropy (Fact 6) in terms of quantum hockey stick divergence.
+Moreover, the upper-bound obtained in [26, Proposition 3] is tighter than the upper-bound obtained in $(i)$ of Corollary 3.
 However, for $\delta>0$, it is not clear if one can obtain a similar upper-bound on quantum relative entropy under $(\varepsilon,\delta)$-LDP CP-TP maps via an integral representation of quantum relative entropy in terms of quantum hockey stick divergence.
@@
 This makes the integration mentioned in Fact 6 un-integrable.
-To resolve this issue, in the subsection below, we come up with a technique which we call truncation. This technique allows us to distill pure DP from non-pure DP.
-VIII-C upper-bound on the relative entropy of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
+In the subsection below, we resolve this issue in classical setting, by coming up with a technique which we call truncation. This technique allows us to distill pure DP from non-pure DP.
+VI-C Upper-bound on the relative entropy of $(\varepsilon,\delta)$-LDP classical channels via the integral representation
 In this subsection, we obtain a tight upper-bound on the relative entropy of the output distributions of any $(\varepsilon,\delta)$-LDP classical channel (Markov kernel) via the integral representation of relative entropy in terms of hockey-stick divergence.
@@
 […]
-  1. i)
+  1. (i)
     $\|\tilde{P}-P\|_{1}\leq 2\delta$ and $\|\tilde{Q}-Q\|_{1}\leq 2\delta.$
-  2. ii)
+  2. (ii)
 […]
-  3. iii)
+  3. (iii)
     if $\operatorname{supp}(P)\subseteq\operatorname{supp}(Q)$, then $\left|D(P\|Q)-D(\tilde{P}\|\tilde{Q})\right|\leq 2\delta\left(\varepsilon+\log\frac{1}{1-\delta}+\frac{2}{m}\right)$.
@@
 Proof.
-See Appendix -H for the proof.
+See Appendix -E for the proof.
 Using Lemma 6 and the integral representation of relative entropy in terms of hockey-stick divergence (Fact 6), we now obtain an upper-bound on the relative entropy of the output distributions of any $(\varepsilon,\delta)$-LDP classical channel (Markov kernel) in Theorem 5 below.
@@
 Proof.
-See Appendix -I for the proof.
-Alternatively, we can obtain a different upper-bound on $D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ by directly applying the continuity bound for relative entropy from (ii) of Lemma 8 and substituting $\varepsilon\leftarrow\left(\varepsilon+\log\frac{1}{1-\delta}\right)$ in Corollary 4.
+See Appendix -F for the proof.
+Alternatively, in the theorem below, we obtain a different upper-bound on $D(\mathcal{K}(P_{X})\|\mathcal{K}(Q_{X}))$ by directly applying the continuity bound for relative entropy from $(ii)$ of Lemma 6 and substituting $\varepsilon\leftarrow\left(\varepsilon+\log\frac{1}{1-\delta}\right)$ in Corollary 3.
 Theorem 6.
@@
 It is important to note that the bound obtained in Theorem 5 is tighter than the bound in Theorem 6.
-This is because the bound in Theorem 7 is obtained by first approximating the $(\varepsilon,\delta)$-DP pair with an $(\varepsilon^{\prime},0)$-DP pair, where $\varepsilon^{\prime}=\varepsilon+\log\frac{1}{1-\delta}$, and then applying the known bounds for pure DP.
+This is because the bound in Theorem 6 is obtained by first approximating the $(\varepsilon,\delta)$-DP pair with an $\varepsilon^{\prime}$-DP pair, where $\varepsilon^{\prime}=\varepsilon+\log\frac{1}{1-\delta}$, and then applying the known bounds for pure DP.
 This approximation introduces looseness, particularly in the leading term, which becomes $\mathcal{O}((\varepsilon+\delta)\tanh(\varepsilon+\delta))$.
@@
 For small $\varepsilon$, the bound in Theorem 5 is therefore significantly tighter.
-IX Conclusion
-In this work, inspired by [7], we develop a framework for studying quantum differential privacy from the perspective of hypothesis testing and Blackwell’s ordering [9].
-We provided a characterization of quantum $(\varepsilon,\delta)$-differential privacy in terms of quantum hypothesis testing divergences, and identified the most informative (in the Blackwell sense) pair of quantum states.
-We use this framework for studying the stability of differentially private quantum learning algorithms.
-Our stability result also generalizes (in the sense that $\delta>0$) the existing bound in the classical settings.
-We also study the problem of quantum privatized parameter estimation, where the trade-off between privacy and statistical utility is characterized via the quantum Fisher information.
+VII Conclusion
+In this work, motivated by Blackwell’s order of informativeness [4, 3] between two statistical experiments, we define an order of informativeness in the quantum setting, which is based on the hypothesis testing divergence dominance.
+In the classical case, Blackwell’s order implies that there always exists a Markov kernel that maps the more informative statistical experiment to the lesser informative statistical experiment.
+However, in the quantum case, a CP-TP map may not always exist [23] from a more informative pair of quantum states to the lesser informative pair of quantum states.
+Despite the lack of existence of such CP-TP maps, we still upper-bound the $f$-divergence of lesser informative pair of quantum states by that of a more informative pair of quantum states.
+This approach allows us to fully characterize $(\varepsilon,\delta)$-QDP mechanisms.
+We use this ordering to study the problem of quantum privatized hypothesis testing and quantum privatized parameter estimation, where the trade-off between privacy and statistical utility is characterized via the quantum hypothesis testing divergence and the quantum Fisher information, respectively.
 We derived explicit expressions for the maximal Fisher information achievable under quantum $(\varepsilon,\delta)$-DP constraints, thereby quantifying the fundamental limits of parameter estimation in the presence of quantum privacy mechanisms.
@@
 Acknowledgments
-The work of NAW was supported in part by MTR/2022/000814.
-The work of MH was supported in part by the National Natural Science Foundation of China under Grant 62171212 and in part by the General Research and Development Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund under Grant GRDP2025-022.
+We are grateful to Henrik Wilming for pointing out an error in the earlier version of this manuscript.
+His comments have greatly helped us to make the technical contributions of this manuscript more rigorous.
+We thank Francesco Buscemi for pointing out several useful references and also thank Mark Wilde for his comments on the earlier version.
+The work of NAW was supported by MTR/2022/000814.
+The work of AD was supported by TCS Research Scholar Program (Cycle 19).
+The work of MH was supported in part by the General R&D Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund (Grant No.
+GRDP2025-022), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No.
+GDZX2505003), the Shenzhen International Quantum Academy (Grant No.
+SIQA2025KFKT07), and the National Natural Science Foundation of China under Grant 62171212.